A Complete Bibliography of the Journal of the Royal Statistical Society, Series A family: 1930–1939

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.02

Title word cross-reference

t [280].

. [552].

1 [559]. 100 [685]. 15th [307, 854]. 16th [739]. 1780 [242]. 1790 [355]. 1806 [1089]. 1817 [163]. 1852 [295]. 1856 [944]. 1860 [723, 919]. 1870 [501]. 18th [611, 607]. 1907 [188]. 1909/10 [176]. 1913 [972, 688, 537]. 1913/14 [176]. 1914 [835]. 1926 [110]. 1926/7 [176]. 1927 [138, 137]. 1928 [246, 59]. 1929 [139, 326, 268, 62, 652]. 1930 [108, 78, 946, 61]. 1930/1 [176]. 1931 [245, 477, 320, 321]. 1932 [464]. 1933 [375, 502]. 1935 [810]. 1936 [674, 1003, 1034, 893, 1058]. 1937 [979, 1012]. 1938 [1064]. 19th [499]. 1st [649].

2 [76, 178, 283, 401, 498, 610, 738, 854, 968, 1077]. 20th [402, 1078, 1077]. 21 [34]. 21st [284, 969, 968, 674]. 23 [212]. 26 [138]. 27 [1004]. 28 [17].

31 [378, 234]. 31st [854, 968, 1077]. 32 [409]. 33 [527]. 34 [617, 624]. 35 [811]. 36 [897].

4 [479].

50 [475].

8 [235].

A. [1018, 281, 282]. A.A.A [753]. Åkerman [244]. Abbey [659]. ABC [808]. Abolish [807]. Abram [216]. Abroad [1064, 497, 185]. Abundance [244]. Accident [407]. Accompanying [963]. According [369]. Account [236, 414, 1082, 593, 108]. Accounting [1063]. Accounts [537]. Act [1012, 851]. Activity [449, 395, 825, 768]. Acts [188]. Actuarial [126]. Actuarie [207]. Acworth [329]. Adam [567, 868]. Adams [295]. Adarkar [715]. Additions [94, 27, 47, 66, 93, 196, 119, 143, 168, 195, 305, 222, 251, 275, 304, 335, 361, 390, 422, 447, 468, 492, 513, 548, 578, 603, 635, 668, 700, 730, 761, 794, 822, 848, 872, 905, 933, 959, 988, 1020, 1048, 1072, 1096]. Address [734]. Addressed [705]. Addresses [247]. Adelphi [76, 178, 283, 401, 498, 610, 738, 854, 968, 1077]. Adequacy [281]. Adjustment [751, 1025]. Administration [163, 722, 1012, 463, 590]. Adult [673, 675]. Adults [189]. Advancement [203, 543]. Advances [913, 1052, 175, 255, 428, 519, 736, 912]. Advisory [830]. Affaires [151]. Affairs [616]. Africa [940, 1087]. Aftalion [456]. After [135, 71, 441, 55]. Aftermath [259]. Age [233, 597]. Ages [673, 675, 769]. Agrawal [648]. Agricole [235]. Agricultural [741, 909, 1030, 663, 811, 523, 429, 234, 409, 441, 527, 624, 661, 722, 897, 852, 800, 996, 383, 295, 112, 353, 573, 722]. Agriculture [351, 441, 967, 486, 214, 296, 721]. Aitken [369]. al [150]. Alanzo [662]. Alaouites [42]. Albert [456, 955]. Albitreccia [694]. Alec [534, 787]. Alexander [383]. Alfred [130, 244, 728, 242, 461, 1017, 1039, 286, 979]. Allan [657]. Allen [943, 137]. Allied [885]. allo [941]. Allocation [704]. Almannac [885]. America [722]. American [723, 620, 240, 922, 21, 149, 356, 355, 159]. Ammoun [42]. among [189]. Analysis [225, 122, 37, 913, 972, 125, 123, 805, 1064, 1028, 943, 778, 127, 152, 891]. Analytically [658]. Anders [262]. Anderson [645, 453]. Andreas [302]. Angas [655, 88, 809]. Angell [748, 35]. Angewandten [779]. Anglaise [42]. Anglo [620]. Anglo-American [620]. Animal [354]. Années [475]. Annual [94, 196, 305, 61, 299, 77, 76, 179, 178, 284, 283, 402, 401, 499, 498, 611, 610, 739, 738, 855, 854, 969, 968, 1078, 1077]. Annuities [207]. Annuities-Certain [207]. Anormali [564]. Ansiaux [538]. Anwar [722]. Apercu [1058]. Application [225, 707, 342, 520, 372]. Applications [53, 151, 647, 529, 740, 316, 500]. Applied [182, 539]. Appreciation [1054, 809]. Approach [30, 12]. Approchees [205]. Approximation [281, 282, 369]. Arable [606]. Arbitration [191]. Archbishop [976]. Area [777, 319, 784, 837]. Argentine [1040]. Arithmetic [1032]. Armitage [623]. Armitage-Smith [623]. Arno [87, 22]. Arnold [894, 479]. Art [345]. Arthur [723, 752, 131, 501, 260, 615, 834, 925, 709]. Articles [26, 46, 65, 92, 118, 142, 167, 194, 221, 250, 274, 303, 334, 360, 389, 421, 446, 467, 491, 512, 547, 577, 602, 634, 667, 699, 729, 760, 793, 821, 847, 871, 904, 932, 958, 987, 1019, 1047, 1071, 1095]. Arts [77, 76, 179, 178, 284, 283, 402, 401, 499, 498, 611, 610, 739, 738, 855, 854, 969, 968, 1078, 1077]. Arvid [649]. Ashton [476, 1089]. Aspects [57, 586, 86, 1054, 188, 340, 495, 327, 865, 676, 1075]. Assembly [797]. Assessment [999]. Assessments [875]. Assignats [85]. Association [203, 543, 299]. Associations [1008]. Assurance [894]. Assured [588]. Astor [661]. Auguste [1009]. Ausgehenden [842]. Australia [893, 279, 82, 440, 464]. Australian [40, 812, 20]. Auswirkung [951]. Authorities [703]. Authority [1084]. Autriche [213]. aux [151].

B [111, 139, 152, 776, 211, 161, 354, 655, 355, 657, 715, 88, 809, 661, 20, 21, 1094, 588, 1093]. Back [67, 95, 145, 169, 197, 363, 391, 424, 550, 580, 670, 796, 824, 907, 935, 960, 1098, 329]. Bagehot [1088]. Balance [927, 654, 201, 454]. Baldwin [504]. Ballore [206]. Bank [786, 481, 377, 1000, 835]. Banking [58, 135, 619, 688, 18, 414, 438, 134, 713, 1002, 1081, 345, 689, 324, 542]. Banks [135, 157, 714, 84, 618, 656]. Baptiste [816]. Barley [487]. Barnard [414, 565]. Barnes [217]. Barrett [135, 380]. Barth [951]. Bartlett [297]. Based [78, 1025]. Basées [235]. Basic [1061]. Basil [618, 632, 347]. Baster [84, 656, 922]. Bates [596]. Batson [62]. Bayliss [710]. be [442]. Beach [688]. Bearing [83]. Becontree [504]. Beet [185]. Beeton [511]. Before [71, 355, 135]. Beginn [842]. Behaviour [972, 748]. Being [22, 722]. Bell [166]. Bellman [659]. Belshaw [214]. Benham [1062]. Benjamin [1008]. Benn [108]. Benoy [786]. Bernard [34, 329, 332]. Bernoulli [641]. Bertil [621]. Berwick [236]. Bessie [442]. Betting [99]. Between [532, 706, 81, 827, 769, 737]. Beveridge [243]. Bevolkerungsproblem [951, 1059]. Bibliography [912, 62]. Biljoen [1087]. Bill [606, 852]. Binomial [770]. Biographical [724]. Biography [558, 386]. Biological [1030]. Biologie [500]. Biology [776]. Biometricians [317]. Biometry [128]. Birmingham [137]. Birth [132, 826, 133, 232, 132]. Birth-Rate [826, 232]. Births [769, 454]. bis [842]. Bituminous [860]. Black [137, 216]. Blackett [632, 618, 347]. Blacklock [1063]. Blé [528]. Blodgett [781]. Board [522]. Body [777]. Bonar [131, 154]. Bond [944]. Bonn [240]. Book [34, 57, 108, 136, 162, 209, 230, 265, 294, 329, 348, 780, 1086, 40, 58, 59, 111, 139, 191, 218, 243, 245, 381, 406, 418, 485, 530, 570, 588, 589, 590, 591, 614, 628, 647, 693, 711, 743, 810, 813, 816, 817, 840, 883, 884, 895, 1060, 453, 741, 82, 477, 12, 15, 19, 37, 54, 86, 106, 109, 127, 129, 152, 180, 184, 263, 264, 370, 371, 646, 691, 692, 708, 709, 719, 723, 740, 745, 752, 753, 776, 777, 779, 781, 786, 973, 350, 866, 976, 78, 84, 85, 126, 130, 131, 132, 134, 135, 138, 157, 160, 207, 210, 211, 212, 213, 244, 271, 291, 344, 346, 375, 411, 433, 434, 435, 457, 480, 481, 538, 567, 618, 619, 620, 687, 713, 714]. Book [744, 928, 948, 1002, 1054, 598, 649, 684, 804, 863, 946, 972, 977, 1061, 798, 11, 13, 14, 38, 55, 89, 150, 153, 161, 183, 185, 186, 235, 236, 285, 297, 349, 354, 378, 380, 383, 384, 385, 416, 437, 488, 501, 502, 536, 558, 592, 623, 626, 627, 645, 648, 650, 686, 710, 721, 746, 747, 748, 782, 785, 803, 815, 835, 837, 838, 839, 889, 891, 894, 920, 921, 927, 950, 951, 981, 1003, 1009, 1010, 1011, 1034, 1042, 1058, 1059, 1062, 1063, 1065, 1066, 1082, 1083, 1085, 1088, 17, 79, 103, 104, 105, 128, 133, 154, 155, 181, 182, 188, 189, 190, 217, 231, 232, 233, 262, 270, 289, 290, 319, 326, 356, 377, 407, 417, 430, 442, 456, 478, 504, 505, 506]. Book [525, 595, 613, 654, 655, 685, 688, 742, 750, 806, 833, 858, 1001, 1036, 1037, 1081, 8, 53, 151, 203, 204, 205, 206, 229, 257, 314, 315, 316, 317, 543, 559, 707, 749, 831, 856, 898, 258, 292, 293, 379, 541, 859, 860, 266, 500, 842, 885, 886, 916, 917, 970, 974, 975, 998, 1030, 1031, 1032, 1055, 1056, 287, 323, 408, 482, 537, 653, 660, 694, 751, 754, 890, 949, 1008, 158, 18, 35, 61, 87, 107, 156, 163, 208, 237, 238, 239, 240, 241, 242, 247, 259, 260, 261, 268, 298, 299, 324, 345, 347, 352, 355, 382, 386, 410, 412, 413, 414, 432, 438, 440, 459, 460, 461, 475, 476, 479, 483, 503, 529, 531, 532, 534, 539, 562, 563, 565, 566, 568]. Book [571, 593, 594, 596, 612, 615, 616, 617, 622, 625, 652, 656, 657, 659, 663, 689, 690, 712, 715, 716, 717, 718, 724, 787, 788, 805, 807, 808, 811, 832, 834, 867, 868, 888, 893, 918, 919, 925, 926, 947, 1000, 1004, 1006, 1039, 1040, 1041, 1064, 1079, 1087, 1089, 941, 942, 971, 1028, 214, 295, 296, 16, 215, 267, 288, 439, 286, 318, 404, 405, 431, 436, 535, 542, 561, 569, 887, 1035, 1057, 9, 10, 36, 83, 88, 113, 320, 533, 560, 651, 683, 774, 775, 802, 809, 814, 836, 857, 861, 862, 864, 882, 892, 922, 923, 943, 944, 945, 978, 979, 980, 982, 999, 1005, 1038, 1084, 896, 22, 41, 42, 60, 62, 112, 234, 322, 351, 353, 376, 409, 441, 486]. Book [487, 527, 528, 572, 573, 597, 624, 658, 661, 662, 720, 722, 778, 783, 812, 841, 843, 897, 983, 1012, 1013, 1067, 20, 21, 39, 56, 110, 137, 159, 216, 246, 269, 327, 328, 462, 784, 80, 81, 102, 321, 325, 369, 372, 373, 374, 403, 415, 454, 455, 458, 463, 464, 484, 507, 508, 526, 540, 564, 621, 830, 865, 924, 1007, 1080, 187, 915, 1029, 1033]. Booms [373]. Booth [231]. Borel [500]. Boverat [981]. Bowley [917, 919]. Boyazoglu [383]. Boyd [777]. Bradford [856, 774]. Braithwaite [350]. Brigden [20]. Brinley [804, 536]. Brissenden [37]. Britain [326, 412, 867, 1064, 572, 553, 172, 639, 967, 570, 1083, 888, 158, 652, 296, 783, 462]. Britain-and [158]. British [895, 866, 481, 620, 744, 354, 1003, 1034, 751, 949, 652, 834, 868, 1040, 1064, 16, 441, 56, 269, 797, 365, 543, 4, 1023, 34, 157, 714, 378, 889, 688, 203, 543]. Broadley [1065]. Brookings [41]. Broomhall [957]. Broster [914]. Brown [717]. Bruck [566]. Brundage [189]. Brunt [315]. Buckingham [723]. Buckinghamshire [1013]. Budgetary [374]. Budgets [34, 374, 536]. Building [122, 890, 659, 485, 814]. Bunle [981]. Bunzel [442]. Burden [101]. Burkhardt [971]. Burns [501]. Burrows [867]. Burton [838]. Business [371, 723, 449, 502, 710, 200, 356, 414, 256, 768, 36, 507, 865, 1008, 264, 349, 838, 325, 372, 374]. By-Product [254]. By-Products [254].

C [34, 693, 15, 776, 207, 375, 713, 928, 782, 921, 1066, 181, 356, 806, 8, 204, 229, 316, 292, 541, 247, 352, 594, 690, 1064, 436, 561, 1035, 560, 683, 774, 775, 882, 573, 20, 39, 56, 137, 369, 455, 924]. Cahill [503]. Calculated [769]. Calculating [1031]. Calculation [771]. Calendar [557]. California [322]. Campion [1083, 862]. Canada [596, 918, 132, 751]. Canadian [299]. Cannan [410]. Cannon [261]. Cantillon [218]. Capital [592, 1074, 1037, 832, 888, 926, 945, 97, 658, 924, 862, 244]. Capitali [564]. Capitalism [566, 867, 922, 457, 921]. Caradog [650, 837, 920]. Cargoes [876]. Carl [502]. Caroline [133]. Carpet [60]. Carr [920, 1064, 858]. Carr-Saunders [920, 1064, 858]. Carrothers [16]. Carrying [1023]. Cartels [484]. Carter [859]. Case [243, 377]. Cassel [135, 457, 562, 712]. Causes [104, 742, 395, 79]. Cecil [377]. Census [75, 584, 866, 149, 772, 649, 643, 769, 355, 918, 228, 880, 147, 399, 477]. Cent [685, 801]. Centenary [472, 704, 476]. Central [134, 618, 713, 345, 1081]. Centre [745]. Century [530, 326, 1000, 1089]. Certain [86, 271, 236, 188, 207]. Ch [535]. Chair [607, 674]. Chamber [866]. Chamberlain [269]. Chambers [407]. Change [155, 233]. Changes [879, 396, 523, 800, 719, 184]. Changing [459]. Characteristics [406]. Charles [1086, 614, 435, 710, 231, 10, 923, 843, 45, 1094]. Charlier [8, 204, 316]. Chegwidden [570]. Children [555]. Chimistes [205]. China [22, 22, 60]. Christopher [711]. Churchill [570]. Chydenius [262]. Cinema [639]. Cinematograph [497]. Clare [808]. Clark [927, 241, 320, 560, 861]. Clash [657]. Class [555]. Classes [53, 101]. Classifications [288]. Clement [838, 1093]. Clement-Leon [1093]. Clinical [875]. Clinics [133]. Cloth [160]. Club [1027]. Co [162, 1064, 573, 1064, 297]. Co-operation [1064, 573, 1064, 297]. Co-operative [1064]. Co-Ordination [162]. Coal [329, 7, 860, 4, 172, 39, 40]. Coal-Mining [172]. Coates [113]. Coatmain [812]. Coefficient [770, 174, 280, 975]. Coefficients [10, 737]. Cohen [269, 485, 572, 720]. Cole [596, 723, 592, 156, 298, 460, 532]. Colin [320, 861]. Collective [1041]. Colonial [59, 797, 293]. Colonies [1034]. Colson [1093]. Combination [315, 86]. Combines [461, 1039, 246]. Coming [186]. Commerce [218, 866, 159]. Commercial [995, 622]. Commerece [816]. Commission [136, 783, 1012]. Commissions [826]. Committee [557, 866, 830]. Commodities [33, 124, 227, 338, 879]. Commodity [781, 415, 676]. Common [979]. Common-Stock [979]. Commonwealth [481]. Communications [265, 488]. Community [101]. Company [79]. Comparative [983]. Comparison [769, 282]. Comparisons [416, 51, 381]. Compete [160]. Competition [31, 834, 506]. Composition [53]. Compound [207, 1025]. Comptabilité [235]. Computation [403]. Concerns [484]. Concise [838]. Conditions [810, 319, 70, 939, 503, 438]. Condliffe [111, 139]. Conference [294, 375]. Conferences [973]. Connor [318]. Consequences [719, 595]. Considerations [663]. Considered [658]. Constant [293]. Construction [780, 405, 768, 978]. Consumable [350]. Consumer [1064]. Consumption [437, 766]. Contentment [244]. Continuous [588, 588]. Contribution [416]. Contributions [285]. Control [133, 257, 834, 310, 692, 211, 707, 374]. Controlling [653]. Convention [966]. Copland [20, 211]. Corn [352, 487]. Corporate [19]. Corporation [889]. Correct [597]. Corrections [939]. Correlation [586, 127, 770, 975, 521, 102, 1055, 915, 81].

Correlations [706]. Corrigenda [30, 673]. Corrigendum [3, 394]. Cost [678, 705, 384]. Costs [397, 416, 177]. Cotta [573]. Cotton [691, 496, 31, 87, 242, 22, 21, 753]. Cotton-Spinning [496]. Cough [340]. Council [76, 178, 283, 401, 498, 610, 738, 705, 854, 968, 1077]. Count [940]. Counties [236]. Countries [494, 582, 471, 694]. Country [896, 137]. Course [710, 970, 608]. Court [191]. Cowden [287]. Cowles [979]. Craig [333]. Craigie [25]. Cram [244]. Cramér [78, 53, 883, 53]. Crash [55]. Crawford [1065]. Credibility [80]. Credit [135, 433, 687, 156, 214, 722, 538, 383, 211, 565]. Crick [689]. Criminalite [42, 983]. Criminelle [42]. Crises [804, 538]. Crisi [564]. Crisis [244, 375, 209, 186]. Critical [891, 894]. Critique [437]. Crocker [186]. Croome [344]. Crop [994, 13]. Crops [138, 81]. Crowd [612]. Crowther [625]. Crum [1035, 19]. Crump [808]. Culpin [189]. Culture [21]. Currency [786, 211, 261]. Current [24, 44, 64, 91, 115, 141, 165, 193, 220, 249, 273, 301, 331, 358, 388, 420, 444, 466, 490, 510, 546, 576, 601, 631, 666, 697, 727, 757, 791, 820, 846, 901, 931, 954, 986, 1016, 1045, 1070, 1092, 594]. Curtis [1001]. Curve [1053]. Curves [853, 452, 914, 342, 520, 74, 915, 405]. Customs [541]. Cycle [938, 587, 743, 344, 721, 534, 807, 832, 865]. Cycles [502, 543, 1038, 371]. Cyclical [83, 781]. Czuber [971].

D [813, 840, 752, 211, 244, 346, 481, 592, 648, 650, 837, 920, 1063, 181, 750, 537, 18, 156, 241, 298, 324, 355, 460, 532, 596, 615, 925, 533, 943, 60, 20, 216, 246, 1094, 1093]. Dagenham [504]. Dahlgren [946]. Dairy [297, 487]. Dalton [536]. Daltroff [568]. Dan [747]. Dana [159]. Daniel [641]. Daniels [862, 711, 903]. dans [267, 978, 42]. d’Apres [213]. d’Apres-Guerre [213]. Darmois [647, 884]. Data [225, 797, 13, 73, 74, 369, 1025]. d’Aujourd’hui [291]. Davenport [776]. David [618, 315, 975]. Davis [708, 559]. Dawson [403]. day [162, 1074, 113]. Deal [922]. Dean [189]. Death [734, 79]. Death-Rates [734]. Deaths [769, 454]. December [649, 76, 178, 283, 401, 498, 556, 610, 738, 854, 968, 1077]. Decimals [1031]. Declining [232]. Decompositions [543]. Deductions [53]. Deficits [747]. Definitions [288]. Degree [80]. Delatour [955]. Delinquant [983]. Delinquency [1051]. Demand [614, 853, 914, 405]. Democracy [41]. Demographie [1058, 500]. Demography [797]. Dennis [247]. Dent [692]. Deolalkar [292]. Department [58, 79]. Dependencies [1034]. Depreciation [658]. Depressed [837]. Depression [394, 210, 244, 480, 437, 366, 239, 652, 400, 295, 752, 505, 412, 432, 440, 864]. Depressions [747, 653, 373]. Deprez [1031]. Derivation [520]. Deriving [853]. Description [1064]. d’Estimation [884]. Determination [80]. Deterrent [1076]. Deuren [646, 740]. Deutsche [482]. Deutschen [951]. Deutschland [842]. Development [200, 660, 552, 56, 137, 830, 111, 566]. Developments [722, 803]. Diagnosis [807]. Diary [479]. Dickson [543]. Dictator [724]. Died [556]. Dienste [370]. Difference [827]. Different [494, 582, 495, 471]. Differential [649, 737]. Dilemma [661]. Ding [60]. Dinner [1027]. d’Intensité [10]. Discount [690]. Discussion [494, 226]. Diseases [612]. Distribution [350, 866, 149, 891, 52, 1050, 456, 827, 994, 174, 280, 862, 880, 772]. Distributions [883]. Distributive [177]. Divers [1058]. Dobb [113]. Dobbs [350]. Dodwell [618]. Données [235]. Doodsoorzaak [153]. Doreen [1067]. Dorothea [350]. Dorp [924]. Douglas [417, 653, 569, 110]. Downfall [712]. Drawn [6]. Drink [327]. Dual [89]. Dubey [648]. Dublin [79, 442]. Duckham [354]. Dudley [287]. Dulles [17]. Duodecimal [1032]. Durand [159]. Durbin [687, 437]. During [258, 441, 972]. Duties [851, 830]. Dyason [20]. Dynamic [614].

each [6]. Early [18, 438]. Earners [643]. Earning [19, 189]. Earnings [678, 705, 37]. East [488]. Eastern [319]. Economic [26, 46, 65, 92, 118, 142, 167, 194, 225, 221, 250, 274, 303, 334, 360, 389, 421, 446, 467, 491, 512, 547, 577, 602, 634, 667, 699, 729, 760, 793, 821, 847, 871, 904, 932, 958, 987, 1019, 1047, 1071, 1095, 111, 591, 86, 138, 375, 1061, 1042, 1063, 583, 937, 595, 1036, 1037, 257, 859, 751, 158, 107, 240, 532, 539, 596, 615, 805, 867, 868, 925, 947, 825, 1035, 113, 922, 923, 982, 62, 441, 1013, 20, 56, 327, 321, 458, 526, 1080, 57, 754, 247, 298, 410, 476, 503, 531, 867, 893, 561, 41]. Economiche [540]. Economics [30, 40, 58, 614, 12, 15, 106, 184, 1022, 506, 806, 483, 441, 39, 265, 626, 460, 539, 1062, 459, 594, 718, 1067]. economique [843, 267]. Économiques [435]. Economist [449, 704]. Economists [943]. Economy [816, 609, 566, 787, 597, 1041]. Edgar [445, 1007]. Edge [1034]. Edgeworth [654, 654]. Edie [346]. Edin [649]. Edinburgh [784]. Edith [777]. Education [516, 1050]. Edward [649, 116, 233, 159, 463]. Edwards [688, 481]. Edwin [261, 410]. Effect [679, 879]. Effects [963, 60]. Efficiencies [770]. Efficiency [992]. Efficient [520]. Effort [596, 260]. Eighteenth [530, 1089]. Eighteenth-Century [530]. Eighth [284, 283]. Eilif [80]. Einar [946]. Einarsen [1038]. Einaudi [817]. Einführung [230, 645]. Einzig [833]. Eisler [375]. Ekas [776]. Elbourne [463]. Elderton [915]. Eldon [598]. Element [1076]. Elementary [53, 648, 455]. Elements [129, 917, 802]. Elizabeth [530]. Ellinger [414, 565]. Ellis [487]. Elsas [842, 482]. Elston [406]. Emanuel [971]. Emigration [16]. Emil [370]. Emile [500]. Empire [744, 354, 56, 271]. Empiriques [205]. Employment [570, 684, 832, 552, 711, 863, 980, 693]. Ended [76, 178, 283, 401, 610, 738, 968, 1077, 498, 854]. Ending [76, 178, 283, 401, 610, 738, 968, 1077, 498, 854]. Enfield [214]. England [481, 835, 920, 295, 673, 638, 176, 675, 994, 852, 800, 530, 18, 352]. English [237, 1084, 734]. Enquiry [683, 20]. Enterprise [587, 533]. Enterprises [995]. Entscheidenden [942]. Entstehen [942]. Enumerated [769]. Enumeration [644]. Environment [963]. Epidemics [612]. Epidemiology [774]. Epstein [560, 109, 61]. Equalisation [1082]. Equalization [593]. Erich [1042, 563]. Erik [946]. Ernest [108, 1086, 135, 457]. Errata [338]. Erratum [765]. Error [9]. Errors [53]. Especially [426]. Essai [42, 218]. Essais [435]. Essay [787, 458]. Essays [247, 1006, 158, 863, 386, 1080]. Estate [504]. Estates [1013]. Estimate [682, 313]. Estimated [174]. Estimating [494, 471]. Estimation [368]. Etat [42]. Ethel [359]. Etude [983]. Europe [810]. European [438]. Evans [570, 889]. Evarts [355]. Eve [112]. Evelyn [157]. Evidences [177]. Evolution [981, 663]. Examination [270, 1064, 683]. Examinations [683, 683]. Examined [243]. Examiners [775]. Exchange [570, 138, 1082, 450, 474, 593, 676, 655, 833, 212]. Exchanges [808]. Existence [706]. Exodus [486]. Expectations [1005]. Expenditure [394, 378, 703, 366, 967]. Experience [79, 1006, 804]. Experiment [225]. Experimental [282, 774]. Experimentalists [882]. Experiments [741, 834]. Export [148]. Exports [287]. Extended [280]. Extension [585, 681]. Eyskens [267]. Ezekiel [127].

F [209, 348, 40, 37, 129, 126, 138, 620, 687, 384, 437, 501, 626, 782, 981, 1066, 356, 407, 970, 975, 1030, 408, 754, 890, 412, 529, 593, 622, 689, 832, 1041, 1087, 971, 215, 431, 9, 36, 836, 658, 20]. F.I.A. [728]. F.R.C.P. [758]. F.R.S. [607, 698, 674]. Factors [21]. Factory [37]. Falbe [302]. Falbe-Hansen [302]. Fall [183]. Families [555, 73]. Family [256]. Fanno [564]. Far [488]. Farm [295]. Farmer [407]. Farming [236, 1067]. Farms [606, 236, 1013]. Fascist [573]. Fatti [941]. Fay [352, 1064, 573]. Feavearyear [237]. Federal [241, 355]. Felice [150, 941]. Felix [184]. Fellner [488]. Fenelon [265, 162, 265]. Ferdynand [718]. Ferenczi [215, 1033]. Fertility [649, 454]. Fields [543]. Fifth [1078, 1077, 543]. Figures [964]. Films [497]. Finance [157, 948, 889, 1003, 485, 746, 785, 1084]. Financial [76, 178, 283, 401, 498, 610, 738, 854, 968, 1077, 745, 375, 620, 236, 377, 163, 259, 414, 36]. Findlay [785]. First [307, 611, 610, 53, 970, 912]. Fisher [1030, 54, 55, 103, 685, 657, 373, 1029, 1006]. Fitted [452, 342]. Fitting [520, 74]. Five [113]. Five-Year [113]. Fletcher [834]. Flogging [1076]. Florence [1064, 439, 560, 12]. Fluctuations [781, 496, 395, 173, 83, 1080, 723]. Folketellingen [649]. following [827]. Fong [60, 60]. Food [138, 71, 161, 322, 572, 138, 1065]. Food-Stuffs [138]. Footnotes [964]. Ford [1085, 526]. Fordham [296]. Forecasting [13, 200, 372, 36, 325, 371]. Foreign [160, 948, 655, 808]. Form [8]. Formation [945]. Formules [205]. Försäkringsaktiebolaget [78]. Forschungsmethoden [971]. Forthcoming [147]. Forty [155, 694]. Forward [833]. Fouad [42]. Foundation [568]. Foundations [441]. Fourth [969, 968, 299]. Fowler [658]. Foxwell [765, 759]. Fragments [798]. Framework [717]. Franc [291, 17]. France [981, 843, 503, 947]. Francis [413]. Frank [212, 782, 431]. Franklin [60]. Frankreich [1059]. Franz [942]. Fraser [412]. Frechet [205]. Frederic [1062, 1031]. Frederick [614, 180, 244, 488, 561, 923, 944]. Free [772, 619]. Freights [876]. French [17, 293, 537, 622]. Freppel [573]. Frequency [52, 8, 827, 520, 915]. Friedrich [344, 475, 566]. Frisbee [264]. Frisch [404, 405, 102]. Fritz [860]. Front [2, 29, 49, 69, 96, 121, 146, 171, 199, 306, 337, 364, 393, 515, 551, 581, 605, 637, 672, 702, 732, 764, 874, 908, 936, 962, 991]. Function [516, 8]. Functions [708, 559]. Fundamental [803, 1037]. Fundamentals [686, 463]. Funds [840]. Future [329, 734, 949, 470, 616, 830, 676, 238, 562, 415].

G [57, 162, 265, 25, 570, 711, 180, 692, 745, 244, 480, 153, 592, 623, 782, 785, 839, 950, 1066, 181, 407, 898, 860, 751, 890, 156, 239, 298, 345, 460, 532, 596, 657, 832, 1000, 1040, 1079, 542, 814, 862, 943, 1005, 376, 597, 137, 415]. G. [1046]. Gaetan [843]. Gain [262]. Gandhi [160]. Gaston [267]. Gayer [752, 615, 925]. General [77, 76, 179, 178, 284, 283, 402, 401, 499, 498, 611, 610, 739, 738, 855, 854, 969, 968, 1078, 1077, 684, 1062, 218]. Genetical [103]. Geneva [831]. Geoffrey [625]. Geographical [994]. Geography [21]. Georg [262]. George [903, 889, 1032, 107, 808, 947, 431, 112, 34, 957]. Georges [647, 884, 291]. Georgievsky [956]. Gerhard [743]. German [135, 502]. Germany [135, 212, 566, 246, 367, 590, 35, 486]. Geschichte [842]. Ghosh [535]. Gibbon [896]. Giblin [20]. Gilbert [266]. Gilboy [530]. Giorgio [540]. Gjermoe [80]. Glass [750]. Glenday [595]. Glückstadt [433, 433]. Gold [209, 244, 879, 289, 688, 208, 238, 239, 412, 413, 562, 617, 712, 1079, 376, 346, 457, 481, 156]. Goodrich [859]. Goods [350]. Gordon [244]. Gosset [902]. Gottfried [346, 864]. Government [1003]. Grace [709]. Graduation [369]. Graham [212]. Grand [291]. Grant [888]. Granville [1034]. Graunt [311]. Gray [691]. Great [570, 628, 1083, 505, 867, 1064, 441, 783, 462, 553, 172, 639, 967, 412]. Greater [417]. Greene [355]. Greenwood [607, 674, 774, 612, 724]. Gregory [457, 562, 290, 238]. Greidanus [411]. Groups [236]. Grover [927]. Growing [609]. Growth [777, 504, 858, 868, 1041, 552, 749, 292]. Grundfragen [779]. Grundriss [314]. Grundzüge [204, 286]. Guerre [213]. Guide [808]. Gustav [712]. Gwilym [896].

H [418, 589, 883, 1046, 82, 477, 129, 692, 708, 781, 344, 346, 592, 626, 710, 721, 838, 950, 981, 1059, 1065, 1083, 326, 377, 417, 1081, 229, 559, 474, 653, 641, 156, 247, 298, 324, 412, 438, 459, 460, 532, 596, 717, 1004, 214, 569, 533, 802, 862, 882, 60, 572, 20, 110, 328, 325]. H. [765, 759]. Haberler [346, 864]. Hadley [133]. Haensel [113]. Haight [622]. Halbwachs [104]. half [912]. Hall [77, 76, 179, 178, 284, 283, 402, 401, 499, 498, 611, 610, 739, 738, 855, 854, 969, 968, 1078, 1077, 593]. Hamer [758]. Handbook [831]. Haney [325]. Hans [433, 268, 865]. Hansen [302]. Hanson [1040]. Harald [78, 285, 53, 792]. Hardy [815]. Harley [326]. Harold [816, 659, 62]. Harper [445, 129]. Harrington [355]. Harris [181, 85, 744, 216]. Harrison [723, 782, 1066]. Harrod [832]. Harry [627, 217, 537, 483]. Hartog [683, 775]. Hasard [886]. Hawtrey [239, 57, 480, 623, 239, 345, 832, 1000, 1079, 376]. Hayek [344]. Hays [857]. Head [397]. Health [271, 326]. Heaton [758]. Heberton [889]. Heilperin [213]. Held [77, 76, 179, 178, 284, 283, 402, 401, 499, 611, 610, 739, 969, 968, 1077, 498, 738, 855, 854, 1078]. Helen [11]. Henry [1094, 218, 753, 619, 333, 686, 891, 356, 511]. Herbert [116]. Heredity [598]. Herman [1028]. Hermann [894, 716]. Hersch [983]. Hevesy [528]. Hickernell [36]. Hicks [1003, 1037, 382]. Higgs [218]. Higher [708, 559, 733]. Hilda [187]. Hill [774, 856]. Him [705]. Hirst [413]. Historical [815, 894, 61]. History [516, 11, 285, 1042, 1082, 18, 237, 476, 1057, 441, 720, 59, 163, 690]. Hoare [1017]. Hobson [597]. Hodson [459, 1004]. Hoffman [415]. Hogg [418]. Holmes [838]. Holmstrom [660]. Home [185, 782, 1066, 1064]. Honor [1006]. Hooper [117, 633]. Hoover [1007]. Horder [1011]. Horobin [596]. Horst [978]. Hours [966]. Housing [638, 504, 70, 964]. Houston [686]. Hsien [60]. Hubbard [746]. Huber [981]. Hugh [536, 1001, 166, 328]. Hughes [536]. Human [777, 598, 21]. Humphrey [432]. Hundred [611, 610, 739, 738, 855, 854, 969, 968, 1078, 1077, 689]. Hundredth [499, 498]. Huntington [950]. Hutchinson [649]. Hutchison [1061]. Hutton [477]. Hyper [212]. Hyper-Inflation [212].

Ichihashi [385]. Ideals [543]. II [182, 1064, 215, 552, 680, 708]. III [319, 543]. Illiteracy [190]. Illness [189]. Illustrated [920]. im [370, 268]. Immigrant [751]. Imperfect [506]. Imperial [84]. Imperialism [927]. Import [622, 830, 851]. Importance [256, 148]. Imports [572]. Improvement [452, 585, 342, 13]. Imre [1033]. Inaccurate [1025]. Ince [543]. Incidence [418, 189]. Including [270]. Income [494, 1060, 109, 398, 946, 379, 919, 320, 861, 945, 471, 950, 1005]. Incomes [73, 733, 1085]. Independent [827, 256]. Index [449, 138, 5, 909, 522, 343, 642, 680, 768, 278, 996, 1025, 780, 887]. Index-Number [343]. Index-Numbers [5, 780]. Indexes [979]. India [477, 161, 541, 438, 470, 825, 308, 292, 379, 87, 542]. Indian [648, 253, 429, 786]. Indices [875]. Individualism [619]. Inductive [518]. Industrial [3, 30, 58, 191, 12, 86, 995, 138, 210, 609, 79, 1, 707, 395, 470, 242, 568, 716, 439, 552, 278, 137, 784, 463, 894, 541, 834, 560, 462]. Industrialisation [60]. Industrialist [1089]. Industrialization [138]. Industrie [978]. Industries [417, 652, 254, 1007]. Industry [50, 86, 496, 398, 397, 185, 354, 292, 860, 87, 910, 172, 639, 552, 22, 159, 60, 691, 157, 627, 890, 461, 1039, 1038, 39, 830]. Inédits [817]. Infant [368]. Infantile [279]. Inference [518, 916]. Inflation [212]. Influence [200, 552, 733]. Information [68, 170, 198, 276, 392, 425, 493, 604, 671, 731, 763, 849, 961, 990, 1073]. Ingenieurs [205]. Inheritance [106]. Inouye [138]. Inquiry [1037]. Institut [475]. Institute [100, 202, 473, 773, 997, 207, 616, 683, 322, 486]. Institutional [30, 12]. Institutions [414, 58]. Instrument [622]. Insurance [590, 79, 269, 328]. Interest [207, 684, 944, 508, 54, 742]. Interests [233, 589]. Interference [1087]. Internal [450, 474]. International [245, 866, 375, 416, 1010, 688, 831, 537, 208, 475, 565, 616, 656, 867, 966, 215, 51, 351, 486, 621, 865, 100, 202, 473, 773, 997, 381, 380, 461, 1039, 215, 683, 527, 662]. Interpolation [885, 9]. Interpretation [427, 797, 922]. Interpretations [215]. Interregional [621]. Introduction [693, 898, 1056, 625, 805, 882, 403, 187, 128, 1035, 980]. Introduzione [941, 150]. Invariants [6]. Inverse [640, 828]. Investigation [3, 588, 643, 200, 1]. Investigations [476]. Investment [1005, 592, 836, 88, 809]. Involved [125]. Iqbal [722]. Ira [264]. Ireland [163]. Irish [772, 880]. Iron [50, 397, 830]. Irvine [1088]. Irving [54, 55, 685, 1006, 373]. Irwin [13].

Isaac [558]. Isbell [109]. Ishii [982]. Isles [567, 16]. Issues [594]. Italy [573]. IV [319, 543].

J [111, 139, 406, 485, 589, 816, 741, 477, 691, 973, 84, 126, 620, 714, 977, 13, 89, 354, 383, 536, 838, 79, 181, 182, 232, 326, 407, 1037, 474, 266, 842, 287, 482, 660, 949, 163, 240, 382, 483, 656, 689, 805, 868, 1087, 887, 774, 922, 441, 812, 20, 507, 508]. J. [641]. J.M [182]. Jack [244]. Jackman [754]. Jackson [181]. Jacob [346]. Jahrhunderts [842]. James [957, 152, 131, 748, 154, 35]. Jameson [326]. January [674, 224]. Japan [138, 22, 138, 982]. Japanese [138, 186, 31, 385]. Jean [816]. Jerome [627]. Jeux [886]. Jevons [704]. Jewish [682, 313]. Jewkes [691]. Jews [367]. Jn [1007]. Joan [863, 506, 980]. Johan [244, 1038]. John [264, 138, 346, 684, 183, 721, 241, 386, 597, 76, 117, 178, 283, 401, 498, 610, 738, 854, 968, 1077]. Johnson [868]. Johnston [539]. Joint [689, 135]. Joint-Stock [135]. Jones [650, 837, 920, 474, 483]. Jordan [369, 10]. Jorgen [972]. Joseph [349, 378, 539, 1084]. Josiah [106, 244, 803, 289, 258, 949, 259, 926]. Joslyn [356]. Journey [22]. Juif [983]. Juive [983]. Jules [780]. Julia [242]. Julius [645]. July [1012]. June [77, 76, 179, 178, 284, 283, 402, 401, 499, 498, 611, 607, 610, 739, 738, 855, 854, 969, 968, 1078, 1077]. Junnosuke [138]. Jürgen [810]. Juvenile [1051].

K.C.B [728, 632]. K.C.S.I. [632]. Kaffee [268]. Kalder [344]. Kalecki [1080]. Kantorowitsch [1055]. Kapitalbildung [977]. Karin [946]. Karl [698, 649, 951, 317, 1054]. Kelley [974, 974]. Kendall [898]. Key [244]. Keynes [182, 788, 684, 182, 386]. Kiernan [163]. Kindersley [714]. King [109, 1018, 690].

Kingdom [919, 339, 277, 228, 608, 881, 1026]. Kirk [721, 407]. Kirsh [1008]. Kisch [377]. Kleeck [571]. Knight [807]. Knowles [56, 56]. Knut [742]. Koch [946]. Kock [1081]. Kohlweiler [370]. Kohn [384]. Korösy [10]. Kotschnig [892]. Kriege [482]. Kuczynski [749, 810, 132, 749, 454]. Kumar [786]. Kuznets [83, 945].

L [485, 82, 19, 109, 244, 928, 38, 648, 746, 838, 17, 505, 655, 8, 204, 229, 316, 543, 917, 534, 787, 788, 919, 439, 318, 1035, 88, 802, 809, 882, 1005, 60, 572, 720, 20, 56]. Labor [58]. Labour [245, 679, 691, 606, 1010, 155, 319, 831, 788, 172, 486, 852, 705, 231, 810, 430, 478, 525, 613]. Lacy [242]. Lagging [706]. Lambert [641]. Lancashire [691, 242, 253]. Laski [816]. l’Astronomie [316]. Law [829, 609, 1008]. Lawley [1041]. Lawrence [209, 348]. Laws [352]. Layton [625]. Leaders [356]. League [831]. Leaning [536]. Learned [892]. Least [9, 369]. Leather [1007]. Lectures [539, 594, 973]. Lederer [977]. Lee [356, 974]. Leisure [244]. Lejeune [780]. L’Emploi [884]. Lenin [628]. L’Enseignement [843]. Leo [841]. Leon [1093]. Leonard [693, 1082, 295]. Lessons [244, 1006, 321]. L’Étalon [434]. L’Étalon-Or [434]. l’etranger [843]. Letters [816]. Level [395, 733, 654, 567, 289, 324]. Levy [802, 894, 716]. Lewis [325]. Liban [42]. Library [94, 27, 47, 66, 93, 196, 119, 143, 168, 195, 305, 222, 251, 275, 304, 335, 361, 390, 422, 447, 468, 492, 513, 548, 578, 603, 635, 668, 700, 730, 761, 794, 822, 848, 872, 905, 933, 959, 988, 1020, 1048, 1072, 1096, 423, 514, 636, 762, 873, 989, 1097]. Liebmann [983]. Liefmann [484]. Life [589, 679, 1054, 79, 155, 231, 319, 430, 478, 525, 613, 704, 982, 240]. Light [803]. Likelihood [342, 520]. Limits [552].

Lindahl [946]. Lindquist [970, 970]. Line [679, 771]. L’Inflation [538]. Lionel [346, 867, 62, 458]. Lippert [487]. List [94, 27, 47, 66, 93, 196, 119, 143, 168, 195, 305, 447, 468, 492, 513, 548, 578, 603, 635, 668, 700, 730, 761, 794, 822, 848, 872, 905, 933, 959, 988, 1020, 1048, 1072, 1096]. Live [254]. Live-Stock [254]. Liverpool [73, 70]. Lives [588]. Living [678, 705, 1010, 416, 950]. Lloyd [751, 1046]. Local [703, 1084]. Location [1007]. Lochhead [263]. Logarithms [1031]. Logic [518, 439]. Logistic [1053]. Lohne [842]. lois [205]. London [714, 155, 231, 319, 430, 478, 525, 613, 690, 555, 679, 682, 313, 417, 266]. Long [937]. Long-Period [937]. L’Or [456]. Lord [976, 1011]. Loss [924]. Louis [558, 79, 442]. Loveday [158]. Lucas [834]. Lucien [151]. Luck [950]. Luigi [817]. Lyons [349].

M [607, 674, 741, 371, 691, 160, 344, 687, 11, 153, 437, 839, 920, 950, 981, 1011, 181, 182, 858, 1081, 205, 898, 541, 842, 1055, 482, 61, 240, 568, 596, 1064, 560, 892, 56, 1007, 1080, 187]. M.D [758]. Macaulay [180, 944]. Macfie [534, 787]. MacGregor [533, 587]. Machine [1031, 597]. Machines [813]. MacLaurin [893]. Macmillan [976]. MacNalty [270]. Made [976]. Madrid [202]. Mainly [598, 882]. Maintaining [210]. Major [612, 724, 774, 173, 25]. Making [111]. Malestroit [817]. Mallet [34, 332]. Malthus [816, 556]. Man [181]. Management [439, 1063, 571, 692]. Manchester [476, 585]. Manes [286, 130, 286]. Manifestation [1038]. Mann [242]. Manual [838, 970]. Manuale [941]. Manufactured [257, 310]. Maps [319]. Marbe [779]. March [151]. Marco [564]. Margaret [418, 745]. Marginal [404]. Marine [323]. Maritime [831]. Market [253, 55, 1074, 888, 714, 14, 782, 1066, 690, 1040]. Marketing [297, 812]. Markets [415]. Marks [775]. Marriages [649]. Marschak [977]. Marsh [693]. Martin [210]. Mary [191, 571, 617]. Material [123]. Materials [138]. Mathematical [225, 152, 708, 973, 913, 1052, 53, 543, 559, 125, 175, 255, 428, 519, 736, 912, 1055, 887, 741, 943]. Mathematique [886]. Mathematische [645, 230]. Mathematischen [204]. Matter [67, 95, 2, 29, 49, 69, 145, 169, 197, 96, 121, 146, 171, 199, 363, 391, 424, 306, 337, 364, 393, 550, 580, 515, 551, 581, 605, 670, 637, 672, 702, 732, 796, 824, 764, 907, 935, 960, 874, 908, 936, 962, 1098, 991]. Mauldon [40]. Maurice [538, 104, 113]. Max [778]. Maximum [342, 520]. May [189, 359]. Maynard [684, 386]. McGill [693]. McGregor [244, 597]. McKay [282]. McKay’s [281]. Meade [805, 508]. Mean [770, 9]. Meaning [38, 328]. Means [1085]. Measurement [396, 181, 749]. Measures [825, 287]. Measuring [404, 454]. Meat [1040]. Mechanism [433, 1082]. Mechanization [627]. Medical [128, 856, 1030, 724, 187, 1034]. Medicine [776]. Meeting [72, 98, 224, 307, 309, 517, 607, 735, 674, 767, 965, 299, 77, 76, 179, 178, 284, 283, 402, 401, 499, 498, 611, 610, 739, 738, 855, 854, 969, 968, 1078, 1077]. Meetings [993, 644]. Member [479]. Memorandum [584, 585]. Memorial [678, 705]. Men [813, 976]. Mendershausen [978]. Mercantile [323]. Mergers [86]. Merle [776]. Merriman [928]. Merseyside [123, 256, 650]. Meteorological [13]. Method [30, 853, 12, 452, 342, 520, 495, 310, 681, 11, 838]. Méthode [151]. Methodes [780, 884]. Methodischen [942]. Methods [225, 494, 776, 623, 229, 707, 288, 404, 882, 471, 454, 1029, 1056, 999, 455, 127]. Metodo [150]. Metropolitan [79]. Meulen [619]. Meyendorff [384]. Michel [213]. Migration [339, 1053, 400, 288, 859].

Migrations [215]. Milford [432]. Milk [720, 783]. Millais [189]. Millar [812]. Millions [659]. Mills [561]. Milnes [784]. Minerals [138]. Miners [571]. Mines [813]. Minimum [543]. Mining [172]. Minister [705]. Minor [173]. Mises [998]. Misuse [583]. Mitchell [782, 1066]. Mittelalter [842]. Moderate [950]. Modern [803, 292, 1056, 461, 1039, 62, 526, 1002, 1001, 261]. Modifying [679]. Moments [452, 342, 6]. Monde [1058]. Mondiale [528, 456]. Monetaire [213]. Monétaires [435]. Monetary [588, 346, 433, 481, 616, 715, 1006, 344, 744, 804, 432, 615, 617, 925]. Money [348, 375, 714, 801, 746, 200, 182, 5, 414, 182, 411, 684, 686, 748, 182, 685, 742, 1001, 237, 347, 563, 413, 507]. Monnoyes [817]. Montague [296]. Montessus [206]. Montgomery [887]. Moore [598]. Mordecai [127]. morrow [439, 113]. Mortality [673, 406, 588, 78, 279, 79, 340, 675, 323, 641, 368, 963, 426]. Mortara [540]. Mouvement [978]. Movement [58, 659, 408, 788, 1064]. Movements [688, 749, 83, 944]. Mr. [914, 788]. Muddle [348]. Mukund [81]. Munshi [541]. Murray [572]. Musikstatistik [841]. Myers [745]. Myra [1001]. Myrddin [570]. Myrddin-Evans [570]. Myth [240, 938].

N [264, 344, 354, 536, 645, 292, 541, 975, 593, 6]. N. [923]. nach [482]. Nancy [464]. National [494, 109, 398, 866, 262, 654, 926, 320, 862, 97, 471, 946, 861, 945]. Nationalism [539, 596]. Nations [831, 481]. Native [271]. Natural [103, 1057]. Nature [218, 1009, 83, 458]. Naturelles [151]. Nautical [885]. Neisser [865]. Nemesis [539]. Nervous [189]. Net [995, 398]. neuen [951]. Neunzehnten [842]. Neurdenburg [153]. Neville [269]. New-York [267]. Newbold [359]. Newton [558]. Neyman [973]. Ninety [77, 76, 179, 178, 284, 283, 402, 401]. Ninety-Eighth [284, 283]. Ninety-Ninth [402, 401]. Ninety-Seventh [179, 178]. Ninety-Sixth [77, 76]. Ninth [402, 401]. Non [643, 983]. Non-Earners [643]. Non-Juive [983]. Nora [784]. Norge [649]. Normal [52, 80]. Norman [808]. Norwegian [1038]. Note [820, 797, 801, 798, 377, 474, 1053, 706, 828, 367]. Note-Book [798]. Notes [24, 44, 64, 91, 115, 141, 165, 193, 220, 249, 273, 301, 331, 358, 388, 420, 444, 466, 490, 510, 546, 576, 601, 631, 545, 575, 600, 630, 666, 697, 727, 757, 665, 696, 726, 756, 791, 846, 790, 819, 845, 870, 901, 931, 954, 986, 900, 930, 953, 985, 1016, 1045, 1070, 1092, 1015, 1044, 1069, 1091, 940]. nous [291]. November [307]. Number [555, 909, 522, 343, 278, 367]. Numbers [5, 877, 642, 1025, 878, 680, 780]. Numerical [281, 152]. Nutrition [875].

O [928, 972, 13]. Oats [487]. Obituary [556]. Observations [884, 315]. Obtained [236]. Occupation [79, 963]. Occupational [79, 426]. Occupations [3, 1]. Office [831, 486, 245, 1010]. Official [585, 226]. Ogilvie [408]. Ohlin [621]. One [499, 498, 611, 610, 739, 738, 855, 854, 969, 968, 1078, 1077]. One-Hundredth [499]. operation [1064, 573, 1064, 297]. Operations [835]. operative [1064]. Opinions [617]. Opportunity [859]. Optimum [1033]. Orchard [138]. Order [867]. Ordinary [607]. Ordination [162]. Organization [439]. Organized [415]. Origine [1009]. Orphan [840]. Orr [354]. Orthogonal [369]. Oscar [590]. Oskar [645]. Oswald [34]. Other [23, 43, 63, 90, 114, 140, 164, 192, 219, 248, 272, 300, 330, 357, 387, 419, 443, 465, 489, 509, 544, 574, 599, 629, 664, 695, 725, 755, 789, 818, 844, 869, 899, 929, 952, 984, 1014, 1043, 1068, 1090, 879, 158, 539, 724, 926, 741]. Otto [230]. Our [71]. Outlay [861].

Outline [481, 619, 857]. Output [606, 398, 397, 172, 523, 800]. Overcrowded [73]. Overcrowding [123]. Overseas [277, 56, 660]. Oxford [1036].

P [25, 628, 12, 646, 740, 776, 350, 78, 135, 160, 207, 210, 244, 618, 649, 380, 1034, 1085, 292, 242, 347, 715, 1064, 295, 439, 535, 113, 999, 353, 507, 526]. Pacific [139]. Pagan [940]. Palestine [42, 399]. Palin [915]. Panandikar [542]. Paper [53, 474, 413]. Papers [271, 289, 1036, 299]. Parabolic [74]. Paradoxes [817]. Parallelism [737]. Pareto [829]. Paris [745]. Parker [346]. Parliament [479]. Part [1013]. Partial [521]. Particularly [188]. Parts [271]. Passenger [266]. Past [734, 858, 470, 278]. Paterson [181]. Patton [1035]. Paul [37, 833, 1056, 653, 694, 569, 113, 528, 110, 956]. Pavlovsky [112]. Pay [37]. Pay-Roll [37]. Pays [1058]. Pearl [128, 1057]. Pearse [87, 22]. Pearson [698, 1054, 1054, 317, 707, 890, 431]. Peasant [1067]. Peddie [89]. Pedersen [972]. Peel [107, 947]. Peers [1064]. Pennsylvania [860]. Penrose [138, 529]. Pension [840]. People [1065]. Percy [45, 269]. Performance [709]. Period [972, 937]. Periodical [28, 48, 120, 144, 223, 252, 336, 362, 448, 469, 549, 579, 669, 701, 795, 823, 821, 906, 934, 1021, 1049]. Periodicals [26, 46, 65, 92, 118, 142, 167, 194, 221, 250, 274, 303, 423, 334, 360, 389, 421, 514, 446, 467, 491, 512, 636, 547, 577, 602, 634, 762, 667, 699, 729, 760, 873, 793, 847, 871, 989, 904, 932, 958, 987, 1097, 1019, 1047, 1071, 1095]. Perplexity [377]. Personal [1085]. Persons [371]. Perspectives [981]. Peter [628, 1089]. Petersen [972]. Pethick [209, 348]. Pethick-Lawrence [209, 348]. Petty [311]. Phelps [717]. Philip [683, 775]. Phthisis [673, 675]. Physical [138, 875]. Physiciens [205]. Pick [782]. Piercy [928].

Pigon [594]. Pigou [247, 15, 921, 806, 436]. Pilgrim [976]. Pioneer [660]. Pirou [843]. Pitfalls [405]. Placement [590]. Places [1031]. Plan [113]. Planned [566, 596, 347]. Planning [692, 867, 893, 572, 662, 596, 654, 596, 923, 896]. Pleasures [596]. Plummer [461, 1039]. Plymouth [651]. Point [709, 6]. Points [426]. Poisson [827]. Policy [567, 619, 620, 804, 688, 541, 324, 432, 615, 617, 925, 947, 113, 662, 687, 241, 616, 622, 715, 805]. Political [30, 816, 12, 299, 532]. Pologne [213]. Polonaise [983]. Polynomials [369, 369]. Ponsonby [266]. Population [82, 131, 186, 981, 643, 858, 644, 749, 682, 769, 355, 552, 6, 983, 313, 78, 1011, 750, 1057, 1033, 161, 529, 982]. Populations [271, 769]. Port [267, 526]. Porteous [840]. Portuguese [377]. Posadowsky [1059]. Posadowsky-Wehner [1059]. Position [138, 481, 4, 830]. Possel [886]. Possessions [797]. Possibilities [290]. Possible [681]. Post [719, 888]. Post-War [719, 888]. Postulates [1061]. Potato [176]. Pound [620, 237]. Poverty [679, 771, 123, 1011]. Power [210, 437, 329, 19, 109]. Powers [543]. Pox [188]. Practical [129, 433, 808, 372]. Practice [594, 867, 239, 565, 1079, 318, 814, 376, 948]. Pratique [434]. Pre [496]. Pre-War [496]. Precision [642, 680]. Predecessors [868]. Preference [889]. Preise [842]. Preliminary [651, 945]. Present [162, 734, 188, 858, 70, 4, 470, 412, 438, 663, 830]. Present-day [162]. Presented [76, 178, 283, 401, 610, 738, 854, 968, 1077, 498]. Président [817, 607, 705, 674]. Pressure [982]. Prevention [188, 538]. Price [719, 567, 879, 972, 396, 289, 654, 395, 450, 474, 324, 617, 887, 733]. Price-Level [395]. Prices [33, 124, 227, 338, 341, 451, 554, 677, 799, 911, 1024, 397, 212, 879, 14, 176, 742, 909, 890, 522, 83, 944, 996, 614, 183, 625, 720, 507, 743, 431]. Pricing [717]. Primary [812]. Principal [406]. Principes [151]. Principle [369]. Principles [191, 134, 803, 1037, 754, 623, 856, 1055, 596]. Private [1083]. Prizeman [567]. Probabilités [316, 740, 206]. Probabilities [646]. Probability [883, 640, 828, 802]. Problem [586, 210, 687, 161, 186, 32, 654, 867, 887, 836, 266, 353, 208, 322, 327]. Probleme [213, 528]. Problèmes [435]. Problems [30, 245, 82, 744, 290, 655, 867, 944, 162, 12, 786, 139, 138, 896]. Proceedings [72, 77, 98, 179, 224, 284, 309, 402, 499, 517, 611, 607, 735, 739, 767, 855, 965, 969, 993, 1078, 299]. Producer [464]. Product [254, 257]. Production [614, 992, 397, 138, 212, 228, 83, 501]. Productivity [994]. Products [310, 254, 487, 297, 812]. Prof. [674]. Professional [838]. Professions [892]. Professor [326, 302, 607, 698, 903, 474, 587]. Profit [924, 533]. Profitableness [236]. Profits [496, 995, 560, 278, 535]. Programme [693, 375]. Progress [138, 271, 620, 657, 595]. Progressive [508]. Prokopovich [14]. Proneness [407]. Properity [864]. Property [1083, 521]. Proposals [433]. Proposed [481]. Prospect [396]. Prosperity [752, 244, 240]. Prospettive [540]. Protectionism [663]. Protectorates [1034]. Provision [214]. Psychological [999]. Psychology [776]. Public [394, 378, 785, 850, 326, 366, 1086, 752, 1083]. Publications [23, 43, 63, 90, 114, 140, 164, 192, 219, 248, 272, 300, 330, 357, 387, 419, 443, 465, 489, 509, 544, 574, 599, 629, 664, 695, 725, 755, 789, 818, 844, 869, 899, 929, 952, 984, 1014, 1043, 1068, 1090, 226]. Punjab [81]. Purchasing [109, 210, 437]. Pure [182]. Purpose [533]. Purposive [495].

Quadratic [543]. Qualities [999]. Quality [226, 257, 707, 310]. Quantitative [30, 12, 200]. Quantitativo [941]. Quantity [226, 343]. Quelques [435, 151]. Quincey [346]. Quinquennial [584]. Quo [158]. Quotas [622]. Quotient [52]. Qureshi [722].

R [57, 40, 263, 264, 132, 207, 480, 714, 1002, 186, 297, 623, 835, 838, 103, 181, 1037, 205, 206, 749, 379, 500, 998, 1030, 1056, 18, 239, 345, 352, 382, 832, 1000, 1064, 1079, 214, 318, 1035, 923, 943, 944, 376, 487, 573, 658, 39, 454, 1029]. Rôle [267]. Ragnar [404, 405, 102]. Rail [294, 694]. Railway [839, 265]. Railways [365, 949, 1086, 660]. Raleigh [131]. Ralph [781, 207, 560]. Ramie [511]. Ranadive [161]. Random [877, 706, 878, 883]. Randomness [877]. Rankin [617, 191]. Rao [81, 379]. Rappard [213]. Rate [826, 232, 508, 1000]. Rates [734, 608, 881, 1026, 944]. Rather [867]. Rational [713]. Rationalization [246, 38]. Raw [138]. Raymond [128, 867, 1057]. Re [311]. Re-Statement [311]. Real [51, 328, 110]. Reality [240]. Received [423, 514, 636, 762, 873, 989, 1097]. Reconstruction [591, 375, 834, 531]. Record [652]. Recorded [769]. Recovery [35, 1004, 210, 260]. Recueil [235]. Reduced [543]. Reedman [536]. Reference [481, 161, 643, 826, 541, 860, 214, 81]. Reflections [438]. Reflexion [886]. Regard [586]. Regarding [584]. Regional [860]. Registration [132]. Regression [737, 453]. Regressive [177]. Regulating [742]. Regulation [261]. Reich [1022]. Reinvestment [1038]. Relating [271, 73, 123, 323, 1076, 426]. Relation [188]. Relations [58, 532, 462, 620]. Relative [10, 148]. Relief [295]. Remedy [375]. Remittances [497]. Rendered [108]. Rene [886]. Renne [487]. Rents [73]. Reorganization [783]. Reply [705]. Report [75, 76, 178, 283, 401, 498, 610, 738, 854, 968, 1077, 976, 236, 270, 22, 783, 1012, 945, 294, 895, 185, 651, 830]. Reports [312, 1034]. Representation [205]. Representative [495]. Reproduction [454]. Republique [983]. Research [693, 322, 1029, 826, 1030]. Researches [126, 453]. Reserve [786]. Reserves [481]. Response [817]. Responsibility [1085]. Results [741, 236, 270, 851, 282]. Retail [582, 891]. Retrospect [396]. Returns [28, 48, 120, 144, 223, 252, 336, 362, 448, 469, 549, 579, 669, 701, 795, 823, 906, 934, 1021, 1049, 609]. Revenue [703]. Reversions [589]. Review [34, 57, 108, 136, 162, 209, 230, 265, 294, 329, 348, 780, 1086, 40, 58, 59, 111, 139, 191, 218, 243, 245, 381, 406, 418, 485, 530, 570, 588, 589, 590, 591, 614, 628, 647, 693, 711, 743, 810, 813, 816, 817, 840, 883, 884, 895, 1060, 453, 741, 82, 477, 12, 15, 19, 37, 54, 86, 106, 109, 127, 129, 152, 180, 184, 263, 264, 370, 371, 646, 691, 692, 708, 709, 719, 723, 740, 745, 752, 753, 776, 777, 779, 781, 786, 973, 350, 866, 976, 78, 84, 85, 126, 130, 131, 132, 134, 135, 138, 157, 160, 207, 210, 211, 212, 213, 244, 271, 291, 344, 346, 375, 411, 433, 434, 435, 457, 480, 481, 538, 567, 618, 619, 620, 687, 713, 714]. Review [744, 928, 948, 1002, 1054, 598, 649, 684, 804, 863, 946, 972, 977, 1061, 801, 11, 13, 14, 38, 55, 89, 150, 153, 161, 183, 185, 186, 235, 236, 285, 297, 349, 354, 378, 380, 383, 384, 385, 416, 437, 488, 501, 502, 536, 558, 592, 623, 626, 627, 645, 648, 650, 686, 710, 721, 746, 747, 748, 782, 785, 803, 815, 835, 837, 838, 839, 889, 891, 894, 920, 921, 927, 950, 951, 981, 1003, 1009, 1010, 1011, 1034, 1042, 1058, 1059, 1062, 1063, 1065, 1066, 1082, 1083, 1085, 1088, 17, 79, 103, 104, 105, 128, 133, 154, 155, 181, 182, 188, 189, 190, 217, 231, 232, 233, 262, 270, 289, 290, 319, 326, 356, 377, 407, 417, 430, 442, 456, 478, 504, 505, 506]. Review [525, 595, 613, 654, 655, 685, 688, 742, 750, 806, 833, 858, 1001, 1036, 1037, 1081, 8, 53, 151, 203, 204, 205, 206, 229, 257, 314, 315, 316, 317, 543, 559, 707, 749, 831, 856, 898, 258, 292, 293, 379, 541, 859, 860, 266, 500, 842, 885, 886, 916, 917, 970, 974, 975, 998, 1030, 1031, 1032, 1055, 1056, 287, 323, 408, 482, 537, 653, 660, 694, 751, 754, 890, 949, 1008, 158, 18, 35, 61, 87, 107, 156, 163, 208, 237, 238, 239, 240, 241, 242, 247, 259, 260, 261, 268, 298, 299, 324, 345, 347, 352, 355, 382, 386, 410, 412, 413, 414, 432, 438, 440, 459, 460, 461, 475, 476, 479, 483, 503, 529, 531, 532, 534, 539, 562, 563, 565, 566, 568]. Review [571, 593, 594, 596, 612, 615, 616, 617, 622, 625, 652, 656, 657, 659, 663, 689, 690, 712, 715, 716, 717, 718, 724, 787, 788, 805, 807, 808, 811, 832, 834, 867, 868, 888, 893, 918, 919, 925, 926, 947, 1000, 1004, 1006, 1039, 1040, 1041, 1064, 1079, 1087, 1089, 941, 942, 971, 1028, 214, 295, 296, 16, 215, 267, 288, 439, 286, 318, 404, 405, 431, 436, 535, 542, 561, 569, 887, 1035, 1057, 9, 10, 36, 83, 88, 113, 320, 533, 560, 651, 683, 774, 775, 802, 809, 814, 836, 857, 861, 862, 864, 882, 892, 922, 923, 943, 944, 945, 978, 979, 980, 982, 999, 1005, 1038, 1084, 896, 22, 41, 42, 60, 62, 112, 234, 322, 351, 353, 376, 409, 441, 486]. Review [487, 527, 528, 572, 573, 597, 624, 658, 661, 662, 720, 722, 778, 783, 812, 841, 843, 897, 983, 1012, 1013, 1067, 20, 21, 39, 56, 110, 137, 159, 216, 246, 269, 327, 328, 462, 784, 80, 81, 102, 321, 325, 369, 372, 373, 374, 403, 415, 454, 455, 458, 463, 464, 484, 507, 508, 526, 540, 564, 621, 830, 865, 924, 1007, 1080, 187, 915, 1029, 1033]. Revival [837]. Revolution [112]. Reynolds [751]. Rhodes [683, 775, 455]. Rice [105]. Richard [218, 1011]. Richards [753, 18]. Riches [291]. Richesse [1009]. Rider [1056]. Riggleman [264]. Riley [814]. Risser [500]. Rist [843, 435]. Road [294, 553, 566, 659, 694]. Roads [1086, 660]. Robbins [867, 505, 62, 458]. Robert [816, 132, 375, 714, 79, 749, 503, 1064, 41, 484, 556]. Robertson [247, 324]. Robinson [863, 133, 506, 980]. Roland [487]. Roll [37, 1042, 563]. Romann [205]. Ronald [440]. Roos [614, 923]. Rosenheim [928]. Roth [268, 802]. Rotterdam [232]. Rountree [839]. Rowntree [661]. Rowse [788]. Roxburgh [236]. Roy [595]. Royal [136, 616, 77, 76, 179, 178, 224, 284, 283, 402, 401, 472, 499, 498, 611, 607, 610, 524, 739, 738, 705, 855, 854, 969, 968, 1078, 1077, 826]. Rubber [910]. Rupert [893, 21]. Rural [486]. Russell [187]. Russia [112, 353, 113, 384, 113]. Ruth [572, 720]. Ryoichi [982].

S [765, 406, 570, 628, 759, 350, 84, 85, 157, 481, 567, 744, 1002, 1054, 648, 835, 1063, 356, 707, 916, 1032, 1008, 87, 476, 596, 656, 1089, 542, 83, 814, 922, 1005, 22, 41, 487, 216]. sa [213, 456]. Saheb [81]. Saisonnier [978]. Salter [260]. Salusbury [270]. same [827]. Sample [643, 73, 70]. Sampled [6]. Samples [6]. Sampling [877, 495, 282, 6, 878]. Samuel [857]. Sanatorium [270]. Sanders [232]. Sanford [190]. Sanger [45]. Sanitary [1034]. Sargant [12, 1064, 439]. Sarkar [786]. Sasuly [778]. Saunders [711, 920, 1064, 858]. Saving [483]. Say [816]. Sayers [1002, 835]. Scale [709]. Scammon [181]. Scandinavian [798]. Scanlan [620]. Scarborough [152]. Scares [410]. Scatter [102]. Schauman [262]. Schmidt [502]. Schoolchildren [875]. Schultz [487]. Science [30, 12, 151, 543, 299, 126, 458, 785, 203]. Sciences [151, 693]. Scotland [236, 523]. Sealy [902]. Seasonal [711]. Second [739, 738, 473, 53, 260]. Section [58]. Secular [83]. Security [657]. Seebohm [661]. Seigneur [817]. seine [951]. Select [62]. Selection [495, 103]. Selkirk [236]. Semi [6]. Semi-Invariants [6]. Series [225, 770, 53, 706, 180, 1028, 80]. Service [570]. Services [815, 850, 123, 895]. ses [981]. Session [307, 674, 473, 773]. Sessional [76, 178, 283, 401, 498, 610, 738, 854, 968, 1077]. Seventh [179, 178, 918]. Seventy [133]. Shackle [1005]. Shaffner [836]. Shah [719]. Shares [889]. Shaw [134]. Sheets [927]. Shepherd [403]. Sheppard [939, 9]. Shewhart [257]. Shifts [860]. Shipping [876, 1023, 1038]. Shirras [785]. Shoe [1007]. Short [957, 1086]. Shortage [418]. Shrigley [617]. Sickness [3, 1]. Significance [586, 770, 1061, 458]. Silver [290, 413]. Simmonds [207]. Simon [1040, 83, 945]. Simple [853, 414, 924, 626]. Since [50, 184, 553, 944, 365, 501, 652, 919]. Sinha [438]. Sir [332, 445, 728, 758, 632, 166]. Situation [290, 811, 234, 409, 624, 897]. Six [588, 866, 594]. Sixth [77, 76]. Skandia [78]. Slum [217]. Slump [1004]. Slumps [807]. Small [188, 256]. Small-Pox [188]. Smart [39]. Smith [723, 567, 623, 868, 812, 713, 747, 891, 189, 417]. Smoothing [180]. Snider [349]. Social [30, 58, 111, 693, 895, 12, 650, 815, 920, 850, 1050, 105, 319, 123, 751, 352, 476, 651, 21, 269, 327, 328, 591]. Sociali [941]. Socialism [566, 921]. Societies [122]. Society [77, 76, 179, 178, 284, 283, 402, 401, 499, 498, 611, 610, 739, 738, 855, 854, 969, 968, 1078, 1077, 485, 659, 814, 77, 76, 179, 178, 224, 284, 283, 402, 401, 472, 499, 498, 611, 607, 610, 524, 739, 738, 705, 855, 854, 969, 968, 1078, 1077, 585, 644, 476, 659]. Sociological [826]. Solution [213, 377, 217]. Somary [184]. Some [586, 244, 1054, 798, 1037, 340, 53, 964, 851, 438, 825, 676, 1075, 426, 126, 532, 944, 865]. Sommes [291]. Sommes-nous [291]. son [981, 267]. Soper [116]. Sources [406]. South [1087, 21]. Southampton [526]. Southworth [293]. Sovereignty [57]. Soviet [113, 746]. Special [481, 161, 826, 541, 214, 81, 860]. Spencer [626]. Spero [216]. Spinning [691, 496]. Springer [645]. Square [9]. Squares [369, 369]. Stabilisation [89, 615, 925]. Stability [450, 474, 568]. Stabilization [346]. Stable [375]. Stagnation [816]. Stamp [244, 803, 289, 258, 949, 259, 926]. Standard [329, 433, 481, 1010, 238, 239, 412, 562, 1079, 376, 433, 712]. Standardization [707, 310]. Stanford [322]. Stanislas [384]. Stanley [504, 704]. State [772, 967, 508, 1087]. Statement [311]. States [59, 61, 944, 722, 110, 58, 385, 190, 806, 287, 561, 560, 415, 591]. Statesman [61]. Stationary [806, 1028]. Statistica [941]. Statistical [30, 77, 76, 26, 46, 65, 92, 179, 178, 118, 142, 167, 194, 225, 224, 284, 283, 221, 250, 274, 303, 402, 401, 334, 360, 389, 421, 472, 499, 498, 446, 467, 491, 512, 611, 607, 610, 524, 547, 577, 602, 634, 545, 575, 600, 630, 739, 738, 705, 667, 699, 729, 760, 665, 696, 726, 756, 855, 854, 793, 821, 847, 871, 790, 819, 845, 870, 969, 968, 904, 932, 958, 987, 900, 930, 953, 985, 1078, 1077, 1027, 1019, 1047, 1071, 1095, 1015, 1044]. Statistical [1069, 1091, 614, 122, 12, 585, 11, 236, 838, 226, 53, 707, 831, 916, 974, 1056, 61, 476, 926, 310, 405, 74, 639, 177, 102, 455, 681, 676, 776, 1030, 1029]. Statistician [704]. Statisticians [317]. Statisticiens [205]. Statistico [150]. Statistics [588, 913, 427, 582, 126, 132, 585, 798, 648, 838, 920, 583, 99, 1052, 175, 255, 428, 519, 736, 912, 970, 520, 1076, 641, 470, 308, 966, 910, 882, 429, 527, 778, 1075, 426, 766, 100, 202, 473, 773, 997, 838, 37, 129, 264, 973, 285, 349, 710, 128, 53, 229, 856, 898, 917, 970, 215, 1035, 857, 403, 187, 105, 323, 288, 318]. Statistiek [153]. Statistik [314, 998, 779, 645, 204, 314, 370]. Statistique [151, 500, 475, 647]. Statistiques [884, 235, 206]. Statistische [942]. Statistischen [971]. Steel [50, 397, 830]. Steffensen [126]. Sterling [237, 216, 244]. Steuerreform [951].

Stevens [949]. Stevenson [333]. Stewart [813]. Stock [55, 689, 944, 254, 135, 979]. Stocks [781]. Story [217, 377, 504, 659]. Strange [972]. Stratified [495]. Street [76, 178, 283, 401, 498, 610, 738, 854, 968, 1077]. Strong [233]. Structure [184, 920]. Struggle [750]. Stuart [1086, 105]. Stuart-Williams [1086]. Stubs [1089]. Students [648, 838, 1062]. Studies [614, 322, 105, 724, 926, 662, 1060, 649, 11, 592, 460]. Studio [941]. Study [162, 86, 135, 210, 804, 416, 889, 442, 742, 970, 412, 625, 659, 888, 1028, 256, 1084, 722, 21, 784, 894]. Stuffs [138]. Sugar [185, 487]. Suggested [801, 944]. Suicide [104, 442]. Summary [918]. Supply [138, 71, 405, 161]. Surface [777]. Surplus [263]. Surrounding [784]. Survey [111, 679, 650, 920, 155, 231, 319, 430, 478, 525, 613, 70, 123, 310, 639, 651, 1013, 526, 815, 351]. Sweden [946, 649]. Swedish [78, 804]. Sykes [378, 1084]. Synthetic [1033]. Syrett [948]. Syrie [42]. System [130, 481, 89, 286, 269, 716, 717].

T [570, 708, 719, 207, 244, 457, 1061, 281, 161, 502, 536, 290, 559, 163, 238, 476, 562, 625, 690, 888, 1089, 282, 187]. Tables [154, 543, 1030, 78, 406, 885, 974, 588, 708, 317, 559, 975, 1031]. Talks [479]. Tangle [244]. Tappenden [589]. Tariff [487, 20, 541]. Tariffs [243]. Taussig [356]. Taxation [803, 101, 623, 258, 379]. Taylor [89, 662]. Tchecoslovaquie [213]. Teaching [1075]. Tebbutt [1035]. Technik [370]. Technique [778]. Technology [718]. Temperament [189]. Tendencies [561, 177]. Terence [504]. Terry [1032]. Test [281, 878]. Tests [586, 770, 1085, 709, 407]. Text [207]. Text-Book [207]. Textbook [1062]. thee [111]. Their [83, 529, 256, 1038]. Theo [487]. Theodore [860]. Theorems [125]. Theoretical [614, 944]. Theoretische [314]. Theoretischen [779]. Theorie [646, 886, 434, 316]. Theories [529, 437, 131, 534]. Theory [453, 54, 126, 134, 344, 684, 863, 103, 182, 833, 898, 916, 1055, 239, 382, 715, 1079, 318, 436, 535, 569, 980, 62, 376, 778, 924, 1007, 1080, 1061, 1037, 532, 433]. Theresa [191, 617]. Third [674, 855, 854, 773, 1022, 228]. Thirty [659]. Thomas [816, 536, 556, 157, 804, 333]. Thompson [207]. Thought [1042, 868]. Three [659]. Throop [747]. Tientsin [60]. Time [586, 180, 1028]. Time-Correlation [586]. Times [59, 298]. Timoshenko [14, 353]. Tintner [743]. Tippett [229, 882, 878]. Titmus [1011]. Tjardus [411]. To-day [1074, 113]. To-morrow [439, 113]. Todd [183]. Todhunter [207]. Together [441]. Topley [774]. Totalisator [99]. touchant [817]. Tourette [817]. Tourist [408]. Town [896]. Towns [866]. Townshend [1001]. Tracts [298]. Trade [394, 743, 582, 344, 31, 437, 721, 200, 654, 366, 292, 4, 1023, 470, 158, 239, 242, 534, 565, 522, 807, 832, 277, 296, 148, 201, 938, 173, 587, 948, 380, 621, 480, 837, 1008]. Trader [464]. Trading [415, 676]. Tramp [876]. Transport [162, 265, 329, 7, 312, 553, 266, 136, 162, 294]. Transportation [754]. Trasferimenti [564]. Treasuries [618]. Treatise [30, 12, 182]. Treatment [741, 270]. Tregaskiss [463]. Trenchard [558]. Trend [673, 995, 675, 778]. Trends [501, 937, 858]. Trial [866]. Tribe [940]. Truman [974]. Truptil [714]. Trust [241, 976]. Trusts [928, 484]. Tschuprov [1055]. Tuberculosis [270]. Tuesday [77, 179, 284, 402, 499, 611, 607, 739, 855, 969, 1078]. Turin [628]. Turned [154]. Twenty [473, 773, 278]. Twenty-Second [473]. Twenty-Third [773]. Twilight [922]. Two [52, 827, 706, 495].

U.S.A [501]. Übererzeugung [268]. Udny [898]. Umriss [842]. Unakar [81]. Unbalanced [536]. Unchanging [6]. Under-Consumption [437]. Unemployment [590, 427, 457, 156, 436, 245, 892]. Unit [928]. United [58, 59, 591, 385, 190, 287, 919, 561, 560, 944, 722, 110, 415, 339, 277, 228, 608, 881, 1026]. Units [6]. University [516, 1050, 322]. Upon [78, 879, 415, 83, 1012]. Urban [638]. Ursula [1003]. Urwick [38, 439]. Usage [205]. Use [13, 583]. Used [494, 53, 471]. Uses [310]. Utility [404]. Uven [741].

V [750, 8, 204, 316, 292, 379, 459, 1004, 81]. Vaccination [32, 188]. Vadimus [158]. Vakil [292, 541]. Valedictory [734]. Valeur [1009]. Valid [939]. Valuation [263]. Value [411, 826, 742, 497, 800, 261, 787, 1037]. Vance [21]. Vane [79]. Variables [883, 102]. Variance [913, 125]. Variates [52, 827]. Variation [174, 280]. Variations [711, 978]. Various [3, 1, 101]. Veenstra [860]. Venn [441]. Venture [293]. Vera [713]. Verbal [999]. Verhulst [1053]. Vernon [999]. Versicherungswesen [130, 286]. Versicherungswesens [286]. Versicherungswirtschaft [130, 286]. Versus [450, 474, 921]. Vickers [375]. View [6]. Vigand [302]. Vincent [375, 488]. Vinci [150, 941]. Viner [346]. Virginia [355]. Viscount [661]. Vital [426]. Vladimir [14, 353]. Vol [708, 155, 182, 319, 543, 559]. Volkseinkommen [482, 482]. Volkswirtschaftslehre [230]. Volume [68, 170, 198, 276, 392, 425, 493, 604, 671, 731, 763, 849, 961, 990, 1073, 138, 648, 918, 1064, 215, 523, 784]. vom [842]. vor [482]. Vorgange [942]. Vorlesungen [204].

W [209, 348, 1018, 530, 711, 813, 19, 78, 210, 928, 948, 977, 1061, 186, 297, 536, 748, 815, 356, 688, 257, 408, 754, 949, 35, 413, 689, 690, 807, 867, 16, 1035, 9, 113, 774, 862, 487, 812, 374, 915]. W.C [76, 178, 283, 401, 498, 610, 738, 854, 968, 1077]. Wadsworth [242, 689]. Wage [381, 189, 608, 881, 1026]. Wage-Earning [189]. Wages [59, 966, 51, 110, 924, 382, 569, 530, 691, 567, 919]. Wahrheit [998]. Wahrscheinlichkeit [998]. Wahrscheinlichkeitsrechnung [779]. Waight [1082]. Wales [920, 673, 638, 176, 675, 852, 800]. Walker [11, 440]. Walks [479]. Wallace [372]. Walras [1009]. Walter [453, 723, 625, 215, 892, 1088]. War [719, 496, 384, 888, 50, 365, 553, 71, 184, 135, 258, 107, 259, 441]. Wargentin [78]. Warren [371, 890, 431, 36]. Warriner [1067, 246]. Warrington [1089]. Waterlow [1018]. Watson [728]. Way [348, 244, 480, 239]. Wealth [526, 1060]. Weather [81]. Wedgwood [106]. Wehner [1059]. Weigert [590]. Weights [1025]. Weinberger [230]. Weldon [838]. Welfare [15]. Welthandelsware [268]. Werner [566]. Wernette [507]. West [940]. Westergaard [792, 285]. Western [810]. Whale [135, 380]. Wheat [14, 322, 353, 662, 1012, 81, 14, 322, 1012]. which [939]. White [82, 537]. Wholesale [33, 124, 227, 338, 341, 451, 554, 677, 799, 911, 1024, 522]. Whooping [340]. Whooping-Cough [340]. Wickens [1094, 20]. Wicksell [742]. Wickwar [815]. Widow [840]. Wie [942]. Wigglesworth [244]. Wilbert [860]. Wilford [109]. Wilhelm [314]. Wilkinson [82]. Wilks [916]. Willcox [215]. William [243, 903, 213, 618, 728, 758, 1065, 1088, 893, 372, 187, 902, 704]. Williams [1086, 326, 346]. Willis [346]. Willsmore [374]. Wilson [894, 479, 774]. Wilzin [841]. Windett [464]. Winkler [314]. Winston [570, 190]. Without [976]. Wm [134]. Wohl [694]. Wold [1028]. Wolfenden [328]. Wolkswohlstand [482]. Woods [187].

Work

[418, 913, 976, 704, 565, 966, 928, 1054, 526]. Worker [1010, 256, 839, 216]. Workers [37, 1029]. Working [555, 414]. Working-Class [555]. Works [752]. World [82, 184, 244, 375, 14, 158, 61, 412, 440, 460, 811, 322, 624, 897, 414, 459, 858, 890, 351]. Wright [346, 415]. Wynnard [633].

Xénophon [434].

Yamato [385]. Yates [1030]. Year

[76, 178, 283, 401, 498, 610, 738, 854, 968, 1077, 1034, 61, 113, 527]. Year-Book [61]. Years [588, 155, 689, 278]. Yields [944]. York [976, 267]. Young [504, 673, 675, 131]. Yule [898].

Zahlen [942]. Zahn [475]. Zealand [111, 812]. Zeitraum [268]. Zizek [942]. Zolotas [434]. zum [842]. Zweig [718].

References

Hill:1929:ISV

[1] A. Bradford Hill. An investigation of sickness in various industrial occupations. Journal of the Royal Statistical Society, 92(2):183–238, ???? 1929. ISSN 0952-8385. URL http://www.

                                              jstor.org/stable/2341507. See cor-

rigendum [3].

Anonymous:1930:FMa

[2] Anonymous. Front matter. Journal of the Royal Statistical Society, 93 (1):i–viii, ???? 1930. ISSN 0952-

                                              8385. URL http://www.jstor.org/
                                              stable/2342143.

Anonymous:1930:CIS

[3] Anonymous. Corrigendum: An investigation of sickness in various industrial occupations. Journal of the Royal Statistical Society, 93(1):??, ???? 1930. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342144. See [1].

Jones:1930:PPB

[4] J. H. Jones. The present position of the British coal trade. Journal of the Royal Statistical Society, 93 (1):1–63, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342145.

Hawtrey:1930:MIN

[5] R. G. Hawtrey. Money and indexnumbers. Journal of the Royal Statistical Society, 93(1):64–103, ???? 1930. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342146.

Soper:1930:SMM

[6] H. E. Soper. Sampling moments of moments of samples of n units each drawn from an unchanging sampled population, from the point of view of semiinvariants. Journal of the Royal Statistical Society, 93(1):104–114, ???? 1930. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342147.

Anonymous:1930:TC

[7] Anonymous. The transport of coal. Journal of the Royal Statistical Society, 93(1):115–117, ???? 1930. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342148.

I:1930:BRBa

[8] J. O. I. Book review: “A New Form of the Frequency Function.”, by C. V. L. Charlier. Journal of the Royal Statistical Society, 93(1):118–120, ???? 1930.

ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2342149.

S:1930:BRBa

[9] H. E. S. Book review: Interpolation with Least Mean Square of Error, by W. F. Sheppard. Journal of the Royal Statistical Society, 93(1):120, ???? 1930. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2342150.

S:1930:BRBb

[10] H. E. S. Book review: Les Coefficients d’Intensité Relative de Kőrösy, by Charles Jordan. Journal of the Royal Statistical Society, 93(1):120–121, ???? 1930. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2342151.

G:1930:BRBa

[11] M. G. Book review: Studies in the History of Statistical Method, by Helen M. Walker. Journal of the Royal Statistical Society, 93(1):121–122, ???? 1930. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2342152.

C:1930:BRBa

[12] L. R. C. Book review: The Statistical Method in Economics and Political Science: A Treatise on the Quantitative and Institutional Approach to Social and Industrial Problems, by P. Sargant Florence. Journal of the Royal Statistical Society, 93(1):122, ???? 1930. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342153. See corrigenda

[30].

G:1930:BRBb

[13] R. F. G. Book review: Crop Forecasting and the Use of Meteorological Data in Its Improvement, by J. O. Irwin. Journal of the Royal Statistical Society, 93 (1):123–124, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342154.

G:1930:BRBc

[14] R. F. G. Book review: Wheat Prices and the World Wheat Market, by Vladimir Prokopovich Timoshenko. Journal of the Royal Statistical Society, 93(1):124–125, ???? 1930. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342155.

C:1930:BRBb

[15] L. R. C. Book review: The Economics of Welfare, by A. C. Pigou. Journal of the Royal Statistical Society, 93 (1):125–126, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342156.

P:1930:BRB

[16] T. P. Book review: Emigration from the British Isles, by W. A. Carrothers. Journal of the Royal Statistical Society, 93(1):126–128, ???? 1930. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342157.

H:1930:BRBa

[17] R. G. H. Book review: The French Franc 1914–28, by E. L. Dulles. Journal of the Royal Statistical Society, 93 (1):128–131, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342158.

M:1930:BRBa

[18] H. W. M. Book review: The Early History of Banking in England, by R. D.

Richards. Journal of the Royal Statistical Society, 93(1):131–132, ???? 1930. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342159.

C:1930:BRBc

[19] L. R. C. Book review: Corporate Earning Power, by W. L. Crum. Journal of the Royal Statistical Society, 93 (1):132–134, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342160.

V:1930:BRBa

[20] J. W. V. Book review: The Australian Tariff: An Economic Enquiry, by J. B. Brigden; D. B. Copland; E. C. Dyason; L. F. Giblin; C. H. Wickens. Journal of the Royal Statistical Society, 93 (1):134–135, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342161.

V:1930:BRBb

[21] J. W. V. Book review: Human Factors in Cotton Culture: A Study in the Social Geography of the American South, by Rupert B. Vance. Journal of the Royal Statistical Society, 93 (1):135–136, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342162.

T:1930:BRBa

[22] C. T. Book review: The Cotton Industry of Japan and China: Being the Report of the Journey to Japan and China, by Arno S. Pearse. Journal of the Royal Statistical Society, 93(1): 136–138, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342163.

Anonymous:1930:ONPa

[23] Anonymous. Other new publications. Journal of the Royal Statistical Society, 93(1):138–140, ???? 1930. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342164.

Anonymous:1930:CNa

[24] Anonymous. Current notes. Journal of the Royal Statistical Society, 93 (1):141–154, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342165.

Anonymous:1930:MPG

[25] Anonymous. Major P. G. Craigie. Journal of the Royal Statistical Society, 93(1):155–158, ???? 1930. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342166.

Anonymous:1930:SEAa

[26] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 93 (1):159–166, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342167.

Anonymous:1930:LALa

[27] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 93(1):167–176, ???? 1930. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342168.

Anonymous:1930:PRa

[28] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 93 (1):177–184, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342169.

Anonymous:1930:FMb

[29] Anonymous. Front matter. Journal of the Royal Statistical Society, 93 (2):i–viii, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342170.

Anonymous:1930:CSM

[30] Anonymous. Corrigenda: The statistical method in economics and political science: A treatise on the quantitative and institutional approach to social and industrial problems. Journal of the Royal Statistical Society, 93(2):??, ???? 1930. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342171. See

[12].

Ellinger:1930:JCC

[31] Barnard Ellinger and Hugh Ellinger. Japanese competition in the cotton trade. Journal of the Royal Statistical Society, 93(2):185–232, ???? 1930. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342172.

Greenwood:1930:VP

[32] M. Greenwood. The vaccination problem. Journal of the Royal Statistical Society, 93(2):233–270, ???? 1930. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342173.

Anonymous:1930:WPC

[33] Anonymous. Wholesale prices of commodities in 1929. Journal of the Royal Statistical Society, 93(2):271–287, ???? 1930. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342174.

A:1930:BRBa

[34] J. E. A. Book review: British Budgets, 1913–14 to 1920–21, by Bernard Mallet; C. Oswald George. Journal of the Royal Statistical Society, 93(2):288– 290, 1930. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2342175.

M:1930:BRBb

[35] J. M. Book review: The Recovery of Germany, by James W. Angell. Journal of the Royal Statistical Society, 93 (2):290–292, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342176.

S:1930:BRBc

[36] E. C. S. Book review: Financial and Business Forecasting, by Warren F. Hickernell. Journal of the Royal Statistical Society, 93(2):292–293, ???? 1930. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342177.

C:1930:BRBd

[37] L. R. C. Book review: Earnings of Factory Workers 1899 to 1927: An Analysis of Pay-Roll Statistics, by Paul F. Brissenden. Journal of the Royal Statistical Society, 93(2):293–294, ???? 1930. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342178.

G:1930:BRBd

[38] A. P. L. G. Book review: The Meaning of Rationalization, by L. Urwick. Journal of the Royal Statistical Society, 93(2):294–295, ???? 1930. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342179.

V:1930:BRBc

[39] J. W. V. Book review: The Economics of the Coal Industry, by R. C. Smart. Journal of the Royal Statistical Society, 93(2):295–296, ???? 1930. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342180.

B:1930:BRBa

[40] W. A. B. Book review: The Economics of Australian Coal, by F. R. E. Mauldon. Journal of the Royal Statistical Society, 93(2):296–298, ???? 1930. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2342181.

T:1930:BRBb

[41] C. T. Book review: Economic Democracy, by Robert S. Brookings. Journal of the Royal Statistical Society, 93 (2):298–299, ???? 1930. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342182.

T:1930:BRBc

[42] C. T. Book review: La Syrie Criminelle. Essai sur la Criminalite en Syrie, au Liban, dans l’Etat des Alaouites et en Palestine Anglaise, by Fouad Ammoun. Journal of the Royal Statistical Society, 93(2):299–300, ???? 1930. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2342183.

Anonymous:1930:ONPb

[43] Anonymous. Other new publications. Journal of the Royal Statistical Society, 93(2):300–302, ???? 1930. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342184.

Anonymous:1930:CNb

[44] Anonymous. Current notes. Journal of the Royal Statistical Society, 93 (2):303–315, ???? 1930. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342185.

Anonymous:1930:CPS

[45] Anonymous. Charles Percy Sanger. Journal of the Royal Statistical Society, 93(2):316, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342186.

Anonymous:1930:SEAb

[46] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 93 (2):317–323, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342187.

Anonymous:1930:LALb

[47] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 93(2):324–331, ???? 1930. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342188.

Anonymous:1930:PRb

[48] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 93 (2):332–341, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342189.

Anonymous:1930:FMc

[49] Anonymous. Front matter. Journal of the Royal Statistical Society, 93 (3):i–viii, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342067.

Birkett:1930:ISI

[50] M. S. Birkett. The iron and steel industry since the War. Journal of the Royal Statistical Society, 93(3):343–397, ???? 1930. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342068.

Richardson:1930:ICR

[51] J. H. Richardson. International comparisons of real wages. Journal of the Royal Statistical Society, 93(3): 398–441, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342069.

Geary:1930:FDQ

[52] R. C. Geary. The frequency distribution of the quotient of two normal variates. Journal of the Royal Statistical Society, 93(3):442–446, ???? 1930. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342070.

I:1930:BRBb

[53] J. O. I. Book review: On Some Classes of Series Used in Mathematical Statistics, by Harald Cramér; On the Composition of Elementary Errors. First Paper: Mathematical Deductions. Second Paper: Statistical Applications by Harald Cramér. Journal of the Royal Statistical Society, 93(3):447–450, ???? 1930. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342071.

C:1930:BRBe

[54] L. R. C. Book review: The Theory of Interest, by Irving Fisher. Journal of the Royal Statistical Society, 93 (3):450–452, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342072.

G:1930:BRBe

[55] A. P. L. G. Book review: The Stock Market Crash and After, by Irving Fisher. Journal of the Royal Statistical Society, 93(3):452–453, ???? 1930. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342073.

V:1930:BRBd

[56] J. W. V. Book review: The Economic Development of the British Overseas Empire, by L. C. A. Knowles; C. M. Knowles. Journal of the Royal Statistical Society, 93(3):453–456, ???? 1930. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342074.

A:1930:BRBb

[57] P. A. Book review: Economic Aspects of Sovereignty, by R. G. Hawtrey. Journal of the Royal Statistical Society, 93(3):456–457, ???? 1930. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342075.

B:1930:BRBb

[58] W. A. B. Book review: The Labor Banking Movement in the United States, by Industrial Relations Section, Department of Economics and Social Institutions. Journal of the Royal Statistical Society, 93(3):457–459, ???? 1930. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342076.

B:1930:BRBc

[59] W. A. B. Book review: History of Wages in the United States from Colonial Times to 1928. Journal of the Royal Statistical Society, 93(3): 460–461, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342077.

T:1930:BRBd

[60] C. T. Book review: Effects of Industrialisation in China, by Franklin L. Ho; Hsien Ding Fong; Tientsin Carpet Industry, by H. D. Fong. Journal of the Royal Statistical Society, 93 (3):461–462, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342078.

M:1930:BRBc

[61] H. W. M. Book review: The Statesman’s Year-Book: Statistical and Historical Annual of the States of the World for the Year 1930, by M. Epstein. Journal of the Royal Statistical Society, 93(3):462–463, ???? 1930. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342079.

T:1930:BRBe

[62] C. T. Book review: A Select Bibliography of Modern Economic Theory, 1870–1929, by Harold E. Batson; Lionel Robbins. Journal of the Royal Statistical Society, 93(3):463, ???? 1930. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342080.

Anonymous:1930:ONPc

[63] Anonymous. Other new publications. Journal of the Royal Statistical Society, 93(3):464–466, ???? 1930. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342081.

Anonymous:1930:CNc

[64] Anonymous. Current notes. Journal of the Royal Statistical Society, 93 (3):467–476, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342082.

Anonymous:1930:SEAc

[65] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 93 (3):477–480, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342083.

Anonymous:1930:LALc

[66] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 93(3):481–488, ???? 1930. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342084.

Anonymous:1930:BMa

[67] Anonymous. Back matter. Journal of the Royal Statistical Society, 93(3):??, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342085.

Anonymous:1930:VI

[68] Anonymous. Volume information. Journal of the Royal Statistical Society, 93(4):648–vii, ???? 1930. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342335.

Anonymous:1930:FMd

[69] Anonymous. Front matter. Journal of the Royal Statistical Society, 93 (4):i–viii, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342336.

Jones:1930:HLS

[70] D. Caradog Jones and Colin G. Clark. Housing in Liverpool: A survey by sample of present conditions. Journal of the Royal Statistical Society, 93 (4):489–537, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342337.

Flux:1930:OFS

[71] A. W. Flux. Our food supply before and after the War. Journal of the Royal Statistical Society, 93(4):538–556, ???? 1930. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342338.

Anonymous:1930:PM

[72] Anonymous. Proceedings of the meeting. Journal of the Royal Statistical Society, 93(4):556–560, ???? 1930. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342339.

Jones:1930:DRR

[73] D. Caradog Jones. Data relating to rents and incomes in a sample of overcrowded families in Liverpool. Journal of the Royal Statistical Society, 93 (4):561–568, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342340.

Rhodes:1930:FPC

[74] E. C. Rhodes. On the fitting of parabolic curves to statistical data. Journal of the Royal Statistical Society, 93(4):569–572, ???? 1930. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342341.

Anonymous:1930:RC

[75] Anonymous. Report on the census. Journal of the Royal Statistical Society, 93(4):573–575, ???? 1930. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342342.

Anonymous:1930:RCF

[76] Anonymous. Report of the Council for the financial year ended December 31, 1929, and for the sessional year ending June 17, 1930, presented at the ninetysixth Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts, John Street, Adelphi, W.C. 2, on June 17, 1930. Journal of the Royal Statistical Society, 93(4):576–591, ???? 1930.

ISSN 0952-8385. URL http://www.

jstor.org/stable/2342343.

Anonymous:1930:PNS

[77] Anonymous. Proceedings of the ninetysixth Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts on Tuesday, June 17, 1930. Journal of the Royal Statistical Society, 93(4): 592–594, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342344.

E:1930:BRBa

[78] W. P. E. Book review: Försäkringsaktiebolaget Skandia, 1855–1930, by Harald Cramér; Tables of Mortality Based Upon the Swedish Population, by P. W. Wargentin. Journal of the Royal Statistical Society, 93(4):595–598, ???? 1930. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342345.

H:1930:BRBb

[79] A. B. H. Book review: Causes of Death, by Occupation. Occupational Mortality Experience of the Metropolitan Life Insurance Company, Industrial Department, 1922–24 by Louis I. Dublin; Robert J. Vane. Journal of the Royal Statistical Society, 93(4):598–599, ???? 1930. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342346.

W:1930:BRBa

[80] J. W. Book review: Determination of the Degree of Credibility of Normal Series, by Eilif Gjermoe. Journal of the Royal Statistical Society, 93 (4):599–600, ???? 1930. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342347.

W:1930:BRBb

[81] J. W. Book review: Correlation Between Weather and Crops, with Special Reference to Punjab Wheat, by Rao Saheb Mukund V. Unakar. Journal of the Royal Statistical Society, 93 (4):600–601, ???? 1930. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342348.

C-S:1930:BRB

[82] A. M. C.-S. Book review: The World’s Population Problems and A White Australia, by H. L. Wilkinson. Journal of the Royal Statistical Society, 93(4): 601–603, ???? 1930. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342349.

S:1930:BRBd

[83] E. C. S. Book review: Secular Movements in Production and Prices: Their Nature and Bearing upon Cyclical Fluctuations, by Simon S. Kuznets. Journal of the Royal Statistical Society, 93 (4):603–605, ???? 1930. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342350.

E:1930:BRBb

[84] W. A. E. Book review: The Imperial Banks, by A. S. J. Baster. Journal of the Royal Statistical Society, 93 (4):605–607, ???? 1930. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342351.

E:1930:BRBc

[85] W. A. E. Book review: The Assignats, by S. E. Harris. Journal of the Royal Statistical Society, 93(4):607–609, ???? 1930. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2342352.

C:1930:BRBf

[86] L. R. C. Book review: Mergers in Industry: A Study of Certain Economic Aspects of Industrial Combination. Journal of the Royal Statistical Society, 93(4):609–611, ???? 1930. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342353.

M:1930:BRBd

[87] H. W. M. Book review: The Cotton Industry of India, by Arno S. Pearse. Journal of the Royal Statistical Society, 93(4):611–613, ???? 1930. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342354.

S:1930:BRBe

[88] E. C. S. Book review: Investment, by L. L. B. Angas. Journal of the Royal Statistical Society, 93(4):613–614, ???? 1930. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342355.

G:1930:BRBf

[89] A. P. L. G. Book review: The Dual System of Stabilisation, by J. Taylor Peddie. Journal of the Royal Statistical Society, 93(4):614, ???? 1930. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342356.

Anonymous:1930:ONPd

[90] Anonymous. Other new publications. Journal of the Royal Statistical Society, 93(4):614–617, ???? 1930. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342357.

Anonymous:1930:CNd

[91] Anonymous. Current notes. Journal of the Royal Statistical Society, 93 (4):618–629, ???? 1930. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342358.

Anonymous:1930:SEAd

[92] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 93 (4):630–633, ???? 1930. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342359.

Anonymous:1930:LALd

[93] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 93(4):634–637, ???? 1930. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2342360.

Anonymous:1930:ALA

[94] Anonymous. Annual list of additions to the library. Journal of the Royal Statistical Society, 93(4):638–647, ???? 1930. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2342361.

Anonymous:1930:BMb

[95] Anonymous. Back matter. Journal of the Royal Statistical Society, 93(4):??, ???? 1930. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342362.

Anonymous:1931:FMa

[96] Anonymous. Front matter. Journal of the Royal Statistical Society, 94 (1):i–viii, ???? 1931. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341816.

Stamp:1931:NC

[97] Josiah Charles Stamp. The national capital. Journal of the Royal Statistical Society, 94(1):1–25, ???? 1931. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341817.

Anonymous:1931:PM

[98] Anonymous. Proceedings of the meeting. Journal of the Royal Statistical Society, 94(1):25–30, ???? 1931. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341818.

Gordon:1931:STB

[99] A. P. L. Gordon. Statistics of totalisator betting. Journal of the Royal Statistical Society, 94(1):31–76, ???? 1931. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2341819.

Bowley:1931:IIS

[100] A. L. Bowley. The International Institute of Statistics. Journal of the Royal Statistical Society, 94(1):77–82, ???? 1931. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2341820.

Sandral:1931:BTV

[101] D. M. Sandral. The burden of taxation on the various classes of the community. Journal of the Royal Statistical Society, 94(1):83–94, ???? 1931. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341821.

W:1931:BRB

[102] J. W. Book review: Frisch (Ragnar): Correlation and Scatter in Statistical Variables. Journal of the Royal Statistical Society, 94(1):95–98, ???? 1931. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341822.

H:1931:BRBa

[103] A. B. H. Book review: The Genetical Theory of Natural Selection, by R. A. Fisher. Journal of the Royal Statistical Society, 94(1):98–100, ???? 1931. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341823.

H:1931:BRBb

[104] A. B. H. Book review: Les Causes du Suicide, by Maurice Halbwachs. Journal of the Royal Statistical Society, 94 (1):100–101, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341824.

H:1931:BRBc

[105] A. B. H. Book review: Statistics in Social Studies, by Stuart A. Rice. Journal of the Royal Statistical Society, 94 (1):101–102, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341825.

C:1931:BRBa

[106] L. R. C. Book review: The Economics of Inheritance, by Josiah Wedgwood. Journal of the Royal Statistical Society, 94(1):102–104, ???? 1931. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341826.

M:1931:BRBa

[107] H. W. M. Book review: The Economic War, by George Peel. Journal of the Royal Statistical Society, 94(1):104, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341827.

A:1931:BRBa

[108] J. E. A. Book review: Account Rendered, 1900–1930, by Ernest Benn.

Journal of the Royal Statistical Society, 94(1):105–106, ???? 1931. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341828.

C:1931:BRBb

[109] L. R. C. Book review: The National Income and its Purchasing Power, by Wilford Isbell King; L. Epstein. Journal of the Royal Statistical Society, 94(1):106, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341829.

V:1931:BRBa

[110] J. W. V. Book review: Real Wages in the United States, 1890–1926, by Paul H. Douglas. Journal of the Royal Statistical Society, 94(1):107–108, ???? 1931. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341830.

B:1931:BRBa

[111] W. A. B. Book review: New Zealand in thee Making: A Survey of Economic and Social Development, by J. B. Condliffe. Journal of the Royal Statistical Society, 94(1):108–110, ???? 1931. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341831.

T:1931:BRB

[112] R. J. T. Book review: Agricultural Russia on the Eve of the Revolution, by George Pavlovsky. Journal of the Royal Statistical Society, 94(1):110–111, ???? 1931. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341832.

S:1931:BRB

[113] E. C. S. Book review: The Economic Policy of Soviet Russia, by Paul Haensel; Russia To-day and Tomorrow, by Maurice Dobb; The Five-Year Plan, by W. P. Coates. Journal of the Royal Statistical Society, 94 (1):111–113, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341833.

Anonymous:1931:ONPa

[114] Anonymous. Other new publications. Journal of the Royal Statistical Society, 94(1):113–116, ???? 1931. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341834.

Anonymous:1931:CNa

[115] Anonymous. Current notes. Journal of the Royal Statistical Society, 94 (1):117–134, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341835.

Greenwood:1931:HES

[116] M. Greenwood. Herbert Edward Soper. Journal of the Royal Statistical Society, 94(1):135–141, ???? 1931. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341836.

Anonymous:1931:JH

[117] Anonymous. John Hooper. Journal of the Royal Statistical Society, 94 (1):141–143, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341837.

Anonymous:1931:SEAa

[118] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 94 (1):144–152, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341838.

Anonymous:1931:LALa

[119] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 94(1):153–164, ???? 1931. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341839.

Anonymous:1931:PRa

[120] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 94 (1):165–172, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341840.

Anonymous:1931:FMb

[121] Anonymous. Front matter. Journal of the Royal Statistical Society, 94 (2):i–viii, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342363.

Brace:1931:SAB

[122] James Brace. A statistical analysis of building societies. Journal of the Royal Statistical Society, 94(2):173–217, ???? 1931. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342364.

Jones:1931:SSM

[123] D. Caradog Jones. The social survey of Merseyside: An analysis of material relating to poverty, overcrowding, and the social services. Journal of the Royal Statistical Society, 94(2):218–266, ???? 1931. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342365.

Anonymous:1931:WPC

[124] Anonymous. Wholesale prices of commodities in 1930. Journal of the Royal Statistical Society, 94(2):267–283, ???? 1931. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342366.

Irwin:1931:MTI

[125] J. O. Irwin. Mathematical theorems involved in the analysis of variance. Journal of the Royal Statistical Society, 94 (2):284–300, ???? 1931. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342367.

E:1931:BRBa

[126] W. P. E. Book review: Some Recent Researches in the Theory of Statistics and Actuarial Science, by J. F. Steffensen. Journal of the Royal Statistical Society, 94(2):301–302, ???? 1931. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2342368.

C:1931:BRBc

[127] L. R. C. Book review: Methods of Correlation Analysis, by Mordecai Ezekiel. Journal of the Royal Statistical Society, 94(2):302–303, ???? 1931. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342369.

H:1931:BRBd

[128] A. B. H. Book review: Introduction to Medical Biometry and Statistics, by Raymond Pearl. Journal of the Royal Statistical Society, 94(2):303–304, ???? 1931. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2342370.

C:1931:BRBd

[129] L. R. C. Book review: Elements of Practical Statistics, by F. H. Harper. Journal of the Royal Statistical Society, 94(2):304, ???? 1931. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342371.

E:1931:BRBb

[130] W. P. E. Book review: Versicherungswesen: System der Versicherungswirtschaft, by Alfred Manes. Journal of the Royal Statistical Society, 94(2):304–305, ???? 1931. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342372.

E:1931:BRBc

[131] P. G. E. Book review: Theories of Population from Raleigh to Arthur Young, by James Bonar. Journal of the Royal Statistical Society, 94(2):305–306, ???? 1931. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342373.

E:1931:BRBd

[132] P. G. E. Book review: Birth Registration and Birth Statistics in Canada, by Robert R. Kuczynski. Journal of the Royal Statistical Society, 94(2): 307–308, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342374.

H:1931:BRBe

[133] A. B. H. Book review: Seventy Birth Control Clinics, by Caroline Hadley Robinson. Journal of the Royal Statistical Society, 94(2):308–309, ???? 1931. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342375.

E:1931:BRBe

[134] W. A. E. Book review: The Theory and Principles of Central Banking, by Wm. A. Shaw. Journal of the Royal Statistical Society, 94(2):309–311, ???? 1931. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342376.

E:1931:BRBf

[135] W. A. E. Book review: Joint-Stock Banking in Germany: A Study of the German Credit Banks before and After the War, by P. Barrett Whale; Ernest Cassel. Journal of the Royal Statistical Society, 94(2):311–314, ???? 1931. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342377.

A:1931:BRBb

[136] J. E. A. Book review: The Royal Commission on Transport. Journal of the Royal Statistical Society, 94(2): 314–316, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342378.

V:1931:BRBb

[137] J. W. V. Book review: The Industrial Development of Birmingham and the Black Country, 1860–1927, by G. C. Allen. Journal of the Royal Statistical Society, 94(2):316–317, ???? 1931. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342379.

E:1931:BRBg

[138] B. E. Book review: Problems of the Japanese Exchange, 1914–26, by Junnosuke Inouye; Japan’s Economic Position: The Progress of Industrialization, by John E. Orchard; Food Supply and Raw Materials in Japan. An Index of the Physical Volume of Production of Food-Stuffs, Industrial Crops, and Minerals, 1894–1927, by E. F. Penrose. Journal of the Royal Statistical Society, 94(2):317–320, ???? 1931. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342380.

B:1931:BRBb

[139] W. A. B. Book review: Problems of the Pacific, 1929, by J. B. Condliffe. Journal of the Royal Statistical Society, 94(2):320–321, ???? 1931. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342381.

Anonymous:1931:ONPb

[140] Anonymous. Other new publications. Journal of the Royal Statistical Society, 94(2):321–324, ???? 1931. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342382.

Anonymous:1931:CNb

[141] Anonymous. Current notes. Journal of the Royal Statistical Society, 94 (2):325–334, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342383.

Anonymous:1931:SEAb

[142] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 94 (2):335–339, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342384.

Anonymous:1931:LALb

[143] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 94(2):340–348, ???? 1931. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342385.

Anonymous:1931:PRb

[144] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 94 (2):349–358, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342386.

Anonymous:1931:BMa

[145] Anonymous. Back matter. Journal of the Royal Statistical Society, 94(2):??, ???? 1931. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342387.

Anonymous:1931:FMc

[146] Anonymous. Front matter. Journal of the Royal Statistical Society, 94 (3):i–viii, ???? 1931. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341841.

Vivian:1931:FC

[147] S. P. Vivian. The forthcoming census. Journal of the Royal Statistical Society, 94(3):359–372, ???? 1931. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341842.

Snow:1931:RIE

[148] E. C. Snow. The relative importance of export trade. Journal of the Royal Statistical Society, 94(3):373–431, ???? 1931. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341843.

Douglas:1931:ACD

[149] Iris Douglas. The American census of distribution, 1930. Journal of the Royal Statistical Society, 94(3):432–437, ???? 1931. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341844.

G:1931:BRBa

[150] M. G. Book review: Introduzione al Metodo Statistico, by Felice Vinci. Journal of the Royal Statistical Society, 94(3):438–439, ???? 1931. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341845.

I:1931:BRB

[151] J. O. I. Book review: Les Principes de la Méthode Statistique avec Quelques Applications aux Sciences Naturelles et à la Science des Affaires, by Lucien March. Journal of the Royal Statistical Society, 94(3):439–442, ???? 1931. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341846.

C:1931:BRBe

[152] L. R. C. Book review: Numerical Mathematical Analysis, by James B. Scarborough. Journal of the Royal Statistical Society, 94(3):442–443, ???? 1931. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341847.

G:1931:BRBb

[153] M. G. Book review: Doodsoorzaak en Statistiek, by M. G. Neurdenburg. Journal of the Royal Statistical Society, 94(3):443, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341848.

H:1931:BRBf

[154] H. H. Book review: The Tables Turned, by James Bonar. Journal of the Royal Statistical Society, 94 (3):444, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341849.

H:1931:BRBg

[155] A. B. H. Book review: New Survey of London Life and Labour. Vol. I. Forty Years of Change. Journal of the Royal Statistical Society, 94(3):445–447, ???? 1931. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341850.

M:1931:BRBb

[156] H. W. M. Book review: Gold, Credit and Unemployment, by G. D. H. Cole. Journal of the Royal Statistical Society, 94(3):447–449, ???? 1931. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341851.

E:1931:BRBh

[157] W. A. E. Book review: British Banks and the Finance of Industry, by S. Evelyn Thomas. Journal of the Royal Statistical Society, 94(3):449–450, ???? 1931. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341852.

Ll:1931:BRB

[158] G. I. H. Ll. Book review: Britain-and World Trade: Quo Vadimus and Other Economic Essays, by A. Loveday. Journal of the Royal Statistical Society, 94 (3):450–454, ???? 1931. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341853.

V:1931:BRBc

[159] J. W. V. Book review: American Industry and Commerce, by Edward Dana Durand. Journal of the Royal Statistical Society, 94(3):454–456, ???? 1931. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341854.

E:1931:BRBi

[160] B. E. Book review: How to Compete with Foreign Cloth, by M. P. Gandhi. Journal of the Royal Statistical Society, 94(3):456, ???? 1931. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341855.

G:1931:BRBc

[161] R. F. G. Book review: Population Problem of India with Special Reference to Food Supply, by B. T. Ranadive. Journal of the Royal Statistical Society, 94(3):457, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341856.

A:1931:BRBc

[162] J. E. A. Book review: Transport Co-Ordination: A Study of Present-day Transport Problems, by K. G. Fenelon. Journal of the Royal Statistical Society, 94(3):457–459, ???? 1931. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341857.

M:1931:BRBc

[163] H. W. M. Book review: History of the Financial Administration of Ireland to 1817, by T. J. Kiernan. Journal of the Royal Statistical Society, 94 (3):459–460, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341858.

Anonymous:1931:ONPc

[164] Anonymous. Other new publications. Journal of the Royal Statistical Society, 94(3):460–462, ???? 1931. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341859.

Anonymous:1931:CNc

[165] Anonymous. Current notes. Journal of the Royal Statistical Society, 94 (3):463–472, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341860.

M:1931:SHB

[166] H. W. M. Sir Hugh Bell. Journal of the Royal Statistical Society, 94 (3):472–473, ???? 1931. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341861.

Anonymous:1931:SEAc

[167] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 94 (3):474–478, ???? 1931. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341862.

Anonymous:1931:LALc

[168] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 94(3):479–486, ???? 1931. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2341863.

Anonymous:1931:BMb

[169] Anonymous. Back matter. Journal of the Royal Statistical Society, 94(3):??, ???? 1931. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341864.

Anonymous:1931:VI

[170] Anonymous. Volume information. Journal of the Royal Statistical Society, 94(4):680–vii, ???? 1931. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341929.

Anonymous:1931:FMd

[171] Anonymous. Front matter. Journal of the Royal Statistical Society, 94 (4):i–viii, ???? 1931. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341930.

Rhodes:1931:LOC

[172] E. C. Rhodes. Labour and output in the coal-mining industry in Great Britain. Journal of the Royal Statistical Society, 94(4):487–539, ???? 1931. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341931.

Sprague:1931:MMT

[173] O. M. W. Sprague. Major and minor trade fluctuations. Journal of the Royal Statistical Society, 94(4):540–563, ???? 1931. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341932.

McKay:1931:DEC

[174] A. T. McKay. The distribution of the estimated coefficient of variation. Journal of the Royal Statistical Society, 94 (4):564–567, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341933.

Irwin:1931:RAM

[175] J. O. Irwin. Recent advances in mathematical statistics. Journal of the Royal Statistical Society, 94(4):568–578, ???? 1931. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341934.

George:1931:PPE

[176] R. F. George. Potato prices in England and Wales: 1909/10–1913/14 and 1926/7–1930/1. Journal of the Royal Statistical Society, 94(4):579–590, ???? 1931. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341935.

Secrist:1931:SER

[177] Horace Secrist. Statistical evidences of regressive tendencies in distributive costs. Journal of the Royal Statistical Society, 94(4):591–598, ???? 1931. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341936.

Anonymous:1931:RCF

[178] Anonymous. Report of the Council for the financial year ended December 31, 1930, and for the sessional year ending June 16, 1931, presented at the ninetyseventh Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts, John Street, Adelphi, W.C. 2, on June 16, 1931. Journal of the Royal Statistical Society, 94(4):599–611, ???? 1931. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341937.

Anonymous:1931:PNS

[179] Anonymous. Proceedings of the ninetyseventh Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts on Tuesday, June 16, 1931. Journal of the Royal Statistical Society, 94 (4):612–613, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341938.

C:1931:BRBf

[180] L. R. C. Book review: Smoothing of Time Series, by Frederick G. Macaulay. Journal of the Royal Statistical Society, 94(4):614–616, ???? 1931. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341939.

H:1931:BRBh

[181] A. B. H. Book review: The Measurement of Man, by J. A. Harris; C. M. Jackson; D. G. Paterson; R. E. Scammon. Journal of the Royal Statistical Society, 94(4):616–617, ???? 1931. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341940.

H:1931:BRBi

[182] R. G. H. Book review: A Treatise on Money. Vol. I. The Pure Theory of Money, by J.M. Keynes; A Treatise on Money. Vol. II. The Applied Theory of Money, by J. M. Keynes. Journal of the Royal Statistical Society, 94 (4):618–623, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341941.

G:1931:BRBd

[183] R. G. Book review: The Fall in Prices, by John A. Todd. Journal of the Royal Statistical Society, 94(4):623–624, ???? 1931. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341942.

C:1931:BRBg

[184] L. R. C. Book review: Changes in the Structure of World Economics Since the War, by Felix Somary. Journal of the Royal Statistical Society, 94(4): 624–625, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341943.

G:1931:BRBe

[185] R. F. G. Book review: Report on the Sugar Beet Industry at Home and Abroad. Journal of the Royal Statistical Society, 94(4):625–626, ???? 1931. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341944.

G:1931:BRBf

[186] R. F. G. Book review: The Japanese Population Problem–The Coming Crisis, by W. R. Crocker. Journal of the Royal Statistical Society, 94(4): 626–628, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341945.

Y:1931:BRB

[187] G. U. Y. Book review: An Introduction to Medical Statistics, by Hilda M. Woods; William T. Russell. Journal of the Royal Statistical Society, 94 (4):628–630, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341946.

H:1931:BRBj

[188] W. H. H. Book review: A Review of Certain Present Aspects of Small-Pox Prevention, in Relation Particularly to the Vaccination Acts, 1867–1907. Journal of the Royal Statistical Society, 94 (4):630–631, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341947.

H:1931:BRBk

[189] A. B. H. Book review: The Incidence of Illness among Wage-Earning Adults, by Dean K. Brundage; The Nervous Temperament by Millais Culpin; May Smith. Journal of the Royal Statistical Society, 94(4):631–633, ???? 1931. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341948.

H:1931:BRBl

[190] A. B. H. Book review: Illiteracy in the United States, by Sanford Winston. Journal of the Royal Statistical Society, 94(4):633–634, ???? 1931. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341949.

B:1931:BRBc

[191] W. A. B. Book review: Arbitration Principles and the Industrial Court, by Mary Theresa Rankin. Journal of the Royal Statistical Society, 94(4): 634–635, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341950.

Anonymous:1931:ONPd

[192] Anonymous. Other new publications. Journal of the Royal Statistical Society, 94(4):635–640, ???? 1931. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341951.

Anonymous:1931:CNd

[193] Anonymous. Current notes. Journal of the Royal Statistical Society, 94 (4):641–657, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341952.

Anonymous:1931:SEAd

[194] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 94 (4):658–663, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341953.

Anonymous:1931:LALd

[195] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 94(4):664–669, ???? 1931. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341954.

Anonymous:1931:ALA

[196] Anonymous. Annual list of additions to the library. Journal of the Royal Statistical Society, 94(4):670–679, ????

1931. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341955.

Anonymous:1931:BMc

[197] Anonymous. Back matter. Journal of the Royal Statistical Society, 94(4):??, ???? 1931. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341956.

Anonymous:1932:VIa

[198] Anonymous. Volume information. Journal of the Royal Statistical Society, 95(1):v–vii, ???? 1932. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341957.

Anonymous:1932:FM

[199] Anonymous. Front matter. Journal of the Royal Statistical Society, 95(1):i–iii, ???? 1932. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341958.

Glenday:1932:BFQ

[200] R. G. Glenday. Business forecasting: A quantitative investigation of the influence of money on trade development. Journal of the Royal Statistical Society, 95(1):1–75, ???? 1932. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341959.

Snow:1932:BT

[201] E. C. Snow. The balance of trade. Journal of the Royal Statistical Society, 95 (1):76–104, ???? 1932. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341960.

Bowley:1932:IIS

[202] A. L. Bowley. The International Institute of Statistics, Madrid, 1931. Journal of the Royal Statistical Society, 95 (1):105–107, ???? 1932. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341961.

I:1932:BRBa

[203] L. I. Book review: British Association for the Advancement of Science. Journal of the Royal Statistical Society, 95 (1):108–109, ???? 1932. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341962.

I:1932:BRBb

[204] J. O. I. Book review: Vorlesungen über die Grundzüge der Mathematischen Statistik, by von C. V. L. Charlier. Journal of the Royal Statistical Society, 95(1):109–110, ???? 1932. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341963.

I:1932:BRBc

[205] J. O. I. Book review: Representation des lois Empiriques par des Formules Approchees, a l’Usage des Chimistes, des Physiciens, des Ingenieurs et des Statisticiens, by M. Frechet; R. Romann. Journal of the Royal Statistical Society, 95(1):110, ???? 1932. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341964.

I:1932:BRBd

[206] J. O. I. Book review: Probabilites et Statistiques, by R. de Montessus de Ballore. Journal of the Royal Statistical Society, 95(1):110–111, ???? 1932. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341965.

E:1932:BRBa

[207] W. P. E. Book review: The Institute of Actuaries’ Text-Book on Compound Interest and Annuities-Certain, by Ralph Todhunter; R. C. Simmonds; T. P. Thompson. Journal of the Royal Statistical Society, 95(1):112, ???? 1932. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2341966.

M:1932:BRBa

[208] H. W. M. Book review: The International Gold Problem. Journal of the Royal Statistical Society, 95(1): 112–115, ???? 1932. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341967.

A:1932:BRBa

[209] J. E. A. Book review: This Gold Crisis, by F. W. Pethick-Lawrence. Journal of the Royal Statistical Society, 95 (1):115–116, ???? 1932. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341968.

E:1932:BRBb

[210] W. A. E. Book review: The Problem of Maintaining Purchasing Power: A Study of Industrial Depression and Recovery, by P. W. Martin. Journal of the Royal Statistical Society, 95 (1):116–118, ???? 1932. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341969.

E:1932:BRBc

[211] W. A. E. Book review: Credit and Currency Control, by D. B. Copland. Journal of the Royal Statistical Society, 95 (1):118–119, ???? 1932. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341970.

E:1932:BRBd

[212] W. A. E. Book review: Exchange, Prices and Production in Hyper-Inflation: Germany, 1920–23, by Frank Graham. Journal of the Royal Statistical Society, 95(1):119–122, ???? 1932. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341971.

E:1932:BRBe

[213] W. A. E. Book review: Le Probleme Monetaire d’Apres-Guerre et sa Solution en Pologne, en Autriche, et en Tchecoslovaquie, by Michel A. Heilperin; William E. Rappard. Journal of the Royal Statistical Society, 95 (1):122–123, ???? 1932. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341972.

O:1932:BRBa

[214] C. S. O. Book review: The Provision of Credit, with Special Reference to Agriculture, by H. Belshaw; R. R. Enfield. Journal of the Royal Statistical Society, 95(1):123–125, ???? 1932. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341973.

P:1932:BRBa

[215] T. P. Book review: International Migrations. Volume I, Statistics, by I. Ferenczi; International Migrations. Volume II, Interpretations. by Walter F. Willcox. Journal of the Royal Statistical Society, 95(1):125–128, ???? 1932. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341974.

V:1932:BRBa

[216] J. W. V. Book review: The Black Worker, by Sterling D. Spero; Abram S. Harris. Journal of the Royal Statistical Society, 95(1):128–130, ???? 1932. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341975.

H:1932:BRBa

[217] A. B. H. Book review: The Slum: Its Story and Solution, by Harry Barnes. Journal of the Royal Statistical Society, 95(1):130–131, ???? 1932. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341976.

B:1932:BRBa

[218] J. B. Book review: Essai sur la Nature du Commerce en General, by Richard Cantillon; Henry Higgs. Journal of the Royal Statistical Society, 95 (1):131–133, ???? 1932. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341977.

Anonymous:1932:ONPa

[219] Anonymous. Other new publications. Journal of the Royal Statistical Society, 95(1):133–135, ???? 1932. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341978.

Anonymous:1932:CNa

[220] Anonymous. Current notes. Journal of the Royal Statistical Society, 95 (1):136–157, ???? 1932. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341979.

Anonymous:1932:SEAa

[221] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 95 (1):158–166, ???? 1932. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341980.

Anonymous:1932:ALa

[222] Anonymous. Additions to the library. Journal of the Royal Statistical Society, 95(1):167–178, ???? 1932. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341981.

Anonymous:1932:PRa

[223] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 95 (1):179–186, ???? 1932. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341982.

Anonymous:1932:PMR

[224] Anonymous. Proceedings of the meeting of the Royal Statistical Society 19 January, 1932. Journal of the Royal Statistical Society, 95(2):187–223, ???? 1932. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342086.

Anonymous:1932:EAM

[225] Anonymous. An experiment in the application of mathematical methods of statistical analysis to series of economic data. Journal of the Royal Statistical Society, 95(2):224–278, ???? 1932. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342087.

Greenwood:1932:DQQ

[226] M. Greenwood. Discussion on the quantity and quality of official statistical publications. Journal of the Royal Statistical Society, 95(2):279–302, ???? 1932. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342088.

Anonymous:1932:WPC

[227] Anonymous. Wholesale prices of commodities in 1931. Journal of the Royal Statistical Society, 95(2):303–319, ???? 1932. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342089. See

errata [338].

Macrosty:1932:TCP

[228] Henry W. Macrosty. The third census of production of the United Kingdom. Journal of the Royal Statistical Society, 95(2):320–322, ???? 1932. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342090.

I:1932:BRBe

[229] J. O. I. Book review: The Methods of Statistics, by L. H. C. Tippett. Journal of the Royal Statistical Society, 95 (2):323–325, ???? 1932. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342091.

A:1932:BRBb

[230] R. G. D. A. Book review: Mathematische Volkswirtschaftslehre. Eine Einführung, by Otto Weinberger. Journal of the Royal Statistical Society, 95 (2):325–326, ???? 1932. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342092.

H:1932:BRBb

[231] A. B. H. Book review: New Survey of London Life and Labour, by Charles Booth. Journal of the Royal Statistical Society, 95(2):326–328, ???? 1932. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2342093.

H:1932:BRBc

[232] A. B. H. Book review: The Declining Birth-Rate in Rotterdam, by J. Sanders. Journal of the Royal Statistical Society, 95(2):328–329, ???? 1932. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2342094.

H:1932:BRBd

[233] A. B. H. Book review: Change of Interests with Age, by Edward K. Strong. Journal of the Royal Statistical Society, 95(2):329–330, ???? 1932. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342095.

T:1932:BRB

[234] R. J. T. Book review: The Agricultural Situation in 1930–31. Journal of the Royal Statistical Society, 95 (2):330–332, ???? 1932. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342096.

G:1932:BRBa

[235] R. F. G. Book review: Recueil de Statistiques: Basées sur les Données de la Comptabilité Agricole, 1927–8. Journal of the Royal Statistical Society, 95 (2):332–333, ???? 1932. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342097.

G:1932:BRBb

[236] R. F. G. Book review: The Profitableness of Farming in Scotland: Report on the Financial Results Obtained on Certain Groups of Farms in Scotland in 1928–9, with a Statistical Account of the Farms in the Counties of Berwick, Roxburgh, and Selkirk. Journal of the Royal Statistical Society, 95 (2):334–335, ???? 1932. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342098.

M:1932:BRBb

[237] H. W. M. Book review: The Pound Sterling, a History of English Money, by A. E. Feavearyear. Journal of the Royal Statistical Society, 95(2):

335–337, ???? 1932. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342099.

M:1932:BRBc

[238] H. W. M. Book review: The Gold Standard and Its Future, by T. E. Gregory. Journal of the Royal Statistical Society, 95(2):337–339, ???? 1932. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342100.

M:1932:BRBd

[239] H. W. M. Book review: The Gold Standard in Theory and Practice, by R. G. Hawtrey; Trade Depression and the Way Out by R. G. Hawtrey. Journal of the Royal Statistical Society, 95 (2):339–340, ???? 1932. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342101.

M:1932:BRBe

[240] H. W. M. Book review: Prosperity: Myth and Reality in American Economic Life, by M. J. Bonn. Journal of the Royal Statistical Society, 95 (2):340–341, ???? 1932. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342102.

M:1932:BRBf

[241] H. W. M. Book review: The Federal Trust Policy, by John D. Clark. Journal of the Royal Statistical Society, 95 (2):341–342, ???? 1932. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342103.

M:1932:BRBg

[242] H. W. M. Book review: The Cotton Trade and Industrial Lancashire, 1600– 1780, by Alfred P. Wadsworth; Julia de Lacy Mann. Journal of the Royal Statistical Society, 95(2):342–343, ???? 1932. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342104.

B:1932:BRBb

[243] W. A. B. Book review: Tariffs: The Case Examined, by William Beveridge. Journal of the Royal Statistical Society, 95(2):343–344, ???? 1932. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342105.

E:1932:BRBf

[244] D. C. T. E. Book review: Capital in Sterling, by A. P. L. Gordon; Josiah Stamp; The Crisis of 1931 by D. T. Jack; The Key to Prosperity: Abundance: Leisure: Contentment by G. A. Cram; World Depression and World Prosperity by A. G. McGregor; The Gold Tangle and the Way Out by Frederick Wigglesworth; Alfred Wigglesworth; Some Lessons of the World Depression by Johan Åkerman. Journal of the Royal Statistical Society, 95 (2):344–346, ???? 1932. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342106.

B:1932:BRBc

[245] W. A. B. Book review: Unemployment Problems in 1931, by International Labour Office. Journal of the Royal Statistical Society, 95(2):346–347, ???? 1932. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342107.

V:1932:BRBb

[246] J. W. V. Book review: Combines and Rationalization in Germany, 1924–1928, by D. Warriner. Journal of the Royal Statistical Society, 95 (2):347–350, ???? 1932. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342108.

M:1932:BRBh

[247] H. W. M. Book review: Economic Essays and Addresses, by A. C. Pigou; Dennis H. Robertson. Journal of the Royal Statistical Society, 95(2): 350–351, ???? 1932. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342109.

Anonymous:1932:ONPb

[248] Anonymous. Other new publications. Journal of the Royal Statistical Society, 95(2):351–353, ???? 1932. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342110.

Anonymous:1932:CNb

[249] Anonymous. Current notes. Journal of the Royal Statistical Society, 95 (2):354–371, ???? 1932. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342111.

Anonymous:1932:SEAb

[250] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 95 (2):372–377, ???? 1932. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342112.

Anonymous:1932:ALb

[251] Anonymous. Additions to the library. Journal of the Royal Statistical Society, 95(2):378–386, ???? 1932. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342113.

Anonymous:1932:PRb

[252] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 95 (2):387–394, ???? 1932. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342114.

Burnett-Hurst:1932:LIM

[253] A. R. Burnett-Hurst. Lancashire and the Indian market. Journal of the Royal Statistical Society, 95(3):395–454, ???? 1932. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342409.

White:1932:LSP

[254] G. Ronald White. Live-stock byproducts and by-product industries. Journal of the Royal Statistical Society, 95(3):455–497, ???? 1932. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342410.

Irwin:1932:RAM

[255] J. O. Irwin. Recent advances in mathematical statistics. Journal of the Royal Statistical Society, 95(3):498–530, ???? 1932. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342411.

Parker:1932:IWS

[256] H. J. H. Parker. The independent worker and the small family business: A study of their importance on Merseyside. Journal of the Royal Statistical Society, 95(3):531–545, ???? 1932. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342412.

I:1932:BRBf

[257] J. O. I. Book review: The Economic Control of Quality of Manufactured Product, by W. A. Shewhart.

Journal of the Royal Statistical Society, 95(3):546–549, ???? 1932. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342413.

J:1932:BRBa

[258] H. S. J. Book review: Taxation During the War, by Josiah Stamp. Journal of the Royal Statistical Society, 95 (3):549–552, ???? 1932. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342414.

M:1932:BRBi

[259] H. W. M. Book review: The Financial Aftermath of War, by Josiah Stamp. Journal of the Royal Statistical Society, 95(3):552–553, ???? 1932. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342415.

M:1932:BRBj

[260] H. W. M. Book review: Recovery, the Second Effort, by Arthur Salter. Journal of the Royal Statistical Society, 95 (3):553–554, ???? 1932. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342416.

M:1932:BRBk

[261] H. W. M. Book review: Modern Currency and the Regulation of its Value, by Edwin Cannon. Journal of the Royal Statistical Society, 95(3):554–555, ???? 1932. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342417.

H:1932:BRBe

[262] H. H. Book review: The National Gain, by Anders Chydenius; Georg Schauman. Journal of the Royal Statistical Society, 95(3):556–557, ???? 1932.

ISSN 0952-8385. URL http://www.

jstor.org/stable/2342418.

C:1932:BRBa

[263] C. R. V. C. Book review: Valuation and Surplus, by R. K. Lochhead. Journal of the Royal Statistical Society, 95(3):557, ???? 1932. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342419.

C:1932:BRBb

[264] L. R. C. Book review: Business Statistics, by John R. Riggleman; Ira N. Frisbee. Journal of the Royal Statistical Society, 95(3):557–558, ???? 1932. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342420.

A:1932:BRBc

[265] J. E. A. Book review: Railway Economics, by K. G. Fenelon; Transport and Communications by K. G. Fenelon. Journal of the Royal Statistical Society, 95(3):558–560, ???? 1932. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342421.

K:1932:BRB

[266] A. E. K. Book review: London’s Passenger Transport Problem, by Gilbert J. Ponsonby. Journal of the Royal Statistical Society, 95(3):560–561, ???? 1932. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342422.

P:1932:BRBb

[267] T. P. Book review: Le Port de New-York dans son Rôle Économique, by Gaston Eyskens. Journal of the Royal Statistical Society, 95(3):561–563, ???? 1932. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342423.

M:1932:BRBl

[268] H. W. M. Book review: Die Übererzeugung in der Welthandelsware Kaffee im Zeitraum von 1790–1929, by Hans Roth. Journal of the Royal Statistical Society, 95(3):563–564, ???? 1932. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342424.

V:1932:BRBc

[269] J. W. V. Book review: The British System of Social Insurance, by Percy Cohen; Neville Chamberlain. Journal of the Royal Statistical Society, 95 (3):564–565, ???? 1932. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342425.

H:1932:BRBf

[270] A. B. H. Book review: A Report on Tuberculosis, Including an Examination of the Results of Sanatorium Treatment, by A. Salusbury MacNalty. Journal of the Royal Statistical Society, 95(3):565–566, ???? 1932. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342426.

E:1932:BRBg

[271] P. G. E. Book review: Papers Relating to the Health and Progress of Native Populations in Certain Parts of the Empire. Journal of the Royal Statistical Society, 95(3):567–569, ???? 1932. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342427.

Anonymous:1932:ONPc

[272] Anonymous. Other new publications. Journal of the Royal Statistical Society, 95(3):569–571, ???? 1932. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342428.

Anonymous:1932:CNc

[273] Anonymous. Current notes. Journal of the Royal Statistical Society, 95 (3):572–590, ???? 1932. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342429.

Anonymous:1932:SEAc

[274] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 95 (3):591–598, ???? 1932. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342430.

Anonymous:1932:ALc

[275] Anonymous. Additions to the library. Journal of the Royal Statistical Society, 95(3):599–606, ???? 1932. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342431.

Anonymous:1932:VIb

[276] Anonymous. Volume information. Journal of the Royal Statistical Society, 95(4):783–788, ???? 1932. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342037.

Macrosty:1932:OTU

[277] Henry W. Macrosty. The overseas trade of the United Kingdom, 1924–31. Journal of the Royal Statistical Society, 95 (4):607–657, ???? 1932. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342038.

Stamp:1932:IPP

[278] Josiah Stamp. Industrial profits in the past twenty years — a new index number. Journal of the Royal Statistical Society, 95(4):658–683, ???? 1932. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342039.

Exley:1932:IMA

[279] H. J. Exley. Infantile mortality in Australia, 1906–1930. Journal of the Royal Statistical Society, 95(4):684–694, ???? 1932. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342040.

McKay:1932:DCV

[280] A. T. McKay. Distribution of the coefficient of variation and the extended “t” distribution. Journal of the Royal Statistical Society, 95(4):695–698, ???? 1932. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342041.

Fieller:1932:NTA

[281] E. C. Fieller. A numerical test of the adequacy of A. T. McKay’s approximation. Journal of the Royal Statistical Society, 95(4):699–702, ???? 1932. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342042.

Pearson:1932:CMA

[282] E. S. Pearson. Comparison of A. T. McKay’s approximation with experimental sampling results. Journal of the Royal Statistical Society, 95(4): 703–704, ???? 1932. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342043.

Anonymous:1932:RCF

[283] Anonymous. Report of the Council for the financial year ended December 31, 1931, and for the sessional year ending June 21, 1932, presented at the ninety-eighth Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts, John Street, Adelphi, W.C. 2, on June 21, 1932. Journal of the Royal Statistical Society, 95(4):705–717, ???? 1932. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342044.

Anonymous:1932:PNE

[284] Anonymous. Proceedings of the ninetyeighth Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts on Tuesday, June 21st, 1932. Journal of the Royal Statistical Society, 95(4): 718–719, ???? 1932. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342045.

G:1932:BRBc

[285] M. G. Book review: Contributions to the History of Statistics, by Harald Westergaard. Journal of the Royal Statistical Society, 95(4):720–722, ???? 1932. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342046.

R:1932:BRB

[286] G. W. R. Book review: Versicherungswesen: System der Versicherungswirtschaft, by Alfred Manes; Grundzüge des Versicherungswesens, by Alfred Manes. Journal of the Royal Statistical Society, 95(4):722–724, ???? 1932. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342047.

L:1932:BRB

[287] H. L. Book review: Measures of Exports of the United States, by Dudley J. Cowden. Journal of the Royal Statistical Society, 95(4):724–726, ???? 1932. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342048.

P:1932:BRBc

[288] T. P. Book review: Statistics of Migration. Definitions–Methods– Classifications. Journal of the Royal Statistical Society, 95(4):726–727, ???? 1932. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2342049.

H:1932:BRBg

[289] R. G. H. Book review: Papers on Gold and the Price Level, by Josiah Stamp. Journal of the Royal Statistical Society, 95(4):728–729, ???? 1932. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342050.

H:1932:BRBh

[290] R. G. H. Book review: The Silver Situation: Problems and Possibilities, by T. E. Gregory. Journal of the Royal Statistical Society, 95(4):729–730, ???? 1932. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2342051.

E:1932:BRBh

[291] W. A. E. Book review: Le Franc d’Aujourd’hui, 1913–1931. Sommesnous Riches?, by Georges Grand. Journal of the Royal Statistical Society, 95 (4):730–731, ???? 1932. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342052.

J:1932:BRBb

[292] H. S. J. Book review: Growth of Trade and Industry in Modern India, by C. N. Vakil; P. V. Deolalkar. Journal of the Royal Statistical Society, 95 (4):731–732, ???? 1932. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342053.

J:1932:BRBc

[293] H. S. J. Book review: The French Colonial Venture, by Constant Southworth. Journal of the Royal Statistical Society, 95(4):733–734, ???? 1932. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342054.

A:1932:BRBd

[294] J. E. A. Book review: Report of the Conference on Rail and Road Transport. Journal of the Royal Statistical Society, 95(4):734–735, ???? 1932. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342055.

O:1932:BRBb

[295] C. S. O. Book review: Agricultural Depression and Farm Relief in England, 1813–1852, by Leonard P. Adams. Journal of the Royal Statistical Society, 95(4):735–737, ???? 1932. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342056.

O:1932:BRBc

[296] C. S. O. Book review: Britain’s Trade and Agriculture, by Montague Fordham. Journal of the Royal Statistical Society, 95(4):737, ???? 1932. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342057.

G:1932:BRBd

[297] R. F. G. Book review: Co-Operation in Marketing Dairy Products, by R. W. Bartlett. Journal of the Royal Statistical Society, 95(4):737–738, ???? 1932. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342058.

M:1932:BRBm

[298] H. W. M. Book review: Economic Tracts for the Times, by G. D. H. Cole. Journal of the Royal Statistical Society, 95(4):738–739, ???? 1932. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342059.

M:1932:BRBn

[299] H. W. M. Book review: Papers and Proceedings of the Fourth Annual Meeting of the Canadian Political Science Association. Journal of the Royal Statistical Society, 95(4):739–740, ???? 1932. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2342060.

Anonymous:1932:ONPd

[300] Anonymous. Other new publications. Journal of the Royal Statistical Society, 95(4):740–743, ???? 1932. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342061.

Anonymous:1932:CNd

[301] Anonymous. Current notes. Journal of the Royal Statistical Society, 95 (4):744–758, ???? 1932. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342062.

Anonymous:1932:PVA

[302] Anonymous. Professor Vigand Andreas Falbe-Hansen. Journal of the Royal Statistical Society, 95(4):759, ???? 1932. ISSN 0952-8385. URL http:/

                                        /www.jstor.org/stable/2342063.

Anonymous:1932:SEAd

[303] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 95 (4):760–763, ???? 1932. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342064.

Anonymous:1932:ALd

[304] Anonymous. Additions to the library. Journal of the Royal Statistical Society, 95(4):764–770, ???? 1932. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342065.

Anonymous:1932:ALAa

[305] Anonymous. Annual list of additions to the library. Journal of the Royal Statistical Society, 95(4):771–782, ???? 1932. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342066.

Anonymous:1933:FMa

[306] Anonymous. Front matter. Journal of the Royal Statistical Society, 96 (1):i–viii, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341865.

Anonymous:1933:FMS

[307] Anonymous. First meeting of session 1932–33, 15th November, 1932. Journal of the Royal Statistical Society, 96(1):1, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341866.

Meston:1933:SI

[308] Lord Meston. Statistics in India. Journal of the Royal Statistical Society, 96 (1):1–14, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341867.

Anonymous:1933:PM

[309] Anonymous. Proceedings of the meeting. Journal of the Royal Statistical Society, 96(1):14–20, ???? 1933. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341868.

Pearson:1933:SUS

[310] E. S. Pearson. A survey of the uses of statistical method in the control and standardization of the quality of manufactured products. Journal of the Royal Statistical Society, 96(1):21–75, ???? 1933. ISSN 0952-8385. URL http:/

                                        /www.jstor.org/stable/2341869.

Greenwood:1933:GPR

[311] M. Greenwood. Graunt and Petty — a re-statement. Journal of the Royal Statistical Society, 96(1):76–81, ???? 1933. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2341870.

Anonymous:1933:RRT

[312] Anonymous. Recent reports on transport. Journal of the Royal Statistical Society, 96(1):82–86, ???? 1933. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341871.

Trachtenberg:1933:EJP

[313] H. L. Trachtenberg. Estimate of the Jewish population of London in 1929. Journal of the Royal Statistical Society, 96(1):87–98, ???? 1933. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341872.

I:1933:BRBa

[314] J. O. I. Book review: Grundriss der Statistik.–I. Theoretische Statistik, by Wilhelm Winkler. Journal of the Royal Statistical Society, 96(1):99–100, ???? 1933. ISSN 0952-8385. URL http:/

                                        /www.jstor.org/stable/2341873.

I:1933:BRBb

[315] J. O. I. Book review: The Combination of Observations, by David Brunt. Journal of the Royal Statistical Society, 96(1):100–101, ???? 1933. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341874.

I:1933:BRBc

[316] J. O. I. Book review: Applications de la Théorie des Probabilités à l’Astronomie, by C. V. L. Charlier. Journal of the Royal Statistical Society, 96(1):101–103, ???? 1933. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341875.

I:1933:BRBd

[317] J. O. I. Book review: Tables for Statisticians and Biometricians, by Karl Pearson. Journal of the Royal Statistical Society, 96(1):103–106, ???? 1933. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341876.

R:1933:BRBa

[318] E. C. R. Book review: Statistics in Theory and Practice, by L. R. Connor. Journal of the Royal Statistical Society, 96(1):106–108, ???? 1933. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341877.

H:1933:BRBa

[319] A. B. H. Book review: The New Survey of London Life and Labour. Vol. III. Survey of Social Conditions. I. The Eastern Area ; The New Survey of London Life and Labour. Vol. IV. Social Survey. I. Eastern Area Maps. Journal of the Royal Statistical Society, 96 (1):108–109, ???? 1933. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341878.

S:1933:BRB

[320] E. C. S. Book review: The National Income, 1924–1931, by Colin Clark. Journal of the Royal Statistical Society, 96(1):110–114, ???? 1933. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341879.

W:1933:BRBa

[321] G. R. W. Book review: The Economic Lessons of 1929–1931. Journal of the Royal Statistical Society, 96 (1):114–115, ???? 1933. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341880.

T:1933:BRBa

[322] R. J. T. Book review: Wheat Studies of the Food Research Institute, Stanford University, California World Wheat Problem. Journal of the Royal Statistical Society, 96(1):115–116, ???? 1933. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2341881.

L:1933:BRBa

[323] H. L. Book review: Statistics Relating to Mortality in the Mercantile Marine. Journal of the Royal Statistical Society, 96(1):116–118, ???? 1933. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341882.

M:1933:BRBa

[324] H. W. M. Book review: Banking Policy and the Price Level, by D. H. Robertson. Journal of the Royal Statistical Society, 96(1):118–120, ???? 1933. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341883.

W:1933:BRBb

[325] G. R. W. Book review: Business Forecasting, by Lewis H. Haney. Journal of the Royal Statistical Society, 96 (1):120–122, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341884.

H:1933:BRBb

[326] W. H. H. Book review: A Century of Public Health in Britain, 1832–1929, by J. H. Harley Williams; Professor Jameson. Journal of the Royal Statistical Society, 96(1):122–124, ???? 1933. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341885.

V:1933:BRBa

[327] J. W. V. Book review: The Social and Economic Aspects of the Drink Problem. Journal of the Royal Statistical Society, 96(1):124–125, ???? 1933. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341886.

V:1933:BRBb

[328] J. W. V. Book review: The Real Meaning of Social Insurance, by Hugh H. Wolfenden. Journal of the Royal Statistical Society, 96(1):125–126, ???? 1933. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341887.

A:1933:BRBa

[329] J. E. A. Book review: Back to the Coal Standard: The Future of Transport and Power, by Bernard Acworth. Journal of the Royal Statistical Society, 96(1):126–127, ???? 1933. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341888.

Anonymous:1933:ONPa

[330] Anonymous. Other new publications. Journal of the Royal Statistical Society, 96(1):128–130, ???? 1933. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341889.

Anonymous:1933:CNa

[331] Anonymous. Current notes. Journal of the Royal Statistical Society, 96 (1):131–147, ???? 1933. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341890.

Anonymous:1933:SBM

[332] Anonymous. Sir Bernard Mallet. Journal of the Royal Statistical Society, 96 (1):148–151, ???? 1933. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341891.

G:1933:THC

[333] M. G. Thomas Henry Craig Stevenson. Journal of the Royal Statistical Society, 96(1):151–156, ???? 1933. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341892.

Anonymous:1933:SEAa

[334] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 96 (1):157–164, ???? 1933. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341893.

Anonymous:1933:ALa

[335] Anonymous. Additions to the library. Journal of the Royal Statistical Society, 96(1):165–174, ???? 1933. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341894.

Anonymous:1933:PRa

[336] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 96 (1):175–182, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341895.

Anonymous:1933:FMb

[337] Anonymous. Front matter. Journal of the Royal Statistical Society, 96 (2):i–viii, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341789.

Anonymous:1933:EWP

[338] Anonymous. Errata: Wholesale prices of commodities in 1931. Journal of the Royal Statistical Society, 96(2):??, ???? 1933. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341790. See

[227].

Leak:1933:MUK

[339] H. Leak and T. Priday. Migration from and to the United Kingdom. Journal of the Royal Statistical Society, 96 (2):183–239, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341791.

Hill:1933:SAM

[340] A. Bradford Hill. Some aspects of the mortality from whooping-cough. Journal of the Royal Statistical Society, 96 (2):240–285, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341792.

Anonymous:1933:WP

[341] Anonymous. Wholesale prices in 1932. Journal of the Royal Statistical Society, 96(2):286–302, ???? 1933. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341793.

Koshal:1933:AMM

[342] R. S. Koshal. Application of the method of maximum likelihood to the improvement of curves fitted by the method of moments. Journal of the Royal Statistical Society, 96(2): 303–313, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341794.

Rhodes:1933:QIN

[343] E. C. Rhodes. A quantity indexnumber. Journal of the Royal Statistical Society, 96(2):314–318, ???? 1933. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341795.

E:1933:BRBa

[344] W. A. E. Book review: Monetary Theory and the Trade Cycle, by Friedrich A. Hayek; N. Kalder; H. M. Croome. Journal of the Royal Statistical Society, 96(2):319–321, ???? 1933. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341796.

M:1933:BRBb

[345] H. W. M. Book review: The Art of Central Banking, by R. G. Hawtrey. Journal of the Royal Statistical Society, 96(2):321–323, ???? 1933. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341797.

E:1933:BRBb

[346] W. A. E. Book review: Gold and Monetary Stabilization, by Quincey Wright; Jacob Viner; Gottfried Haberler; H. Parker Willis; Lionel D. Edie; John H.

Williams. Journal of the Royal Statistical Society, 96(2):323–325, ???? 1933. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341798.

M:1933:BRBc

[347] H. W. M. Book review: Planned Money, by Basil P. Blackett. Journal of the Royal Statistical Society, 96 (2):325–326, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341799.

A:1933:BRBb

[348] J. E. A. Book review: The Money Muddle and the Way Out, by F. W. Pethick-Lawrence. Journal of the Royal Statistical Society, 96(2):327, ???? 1933. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341800.

G:1933:BRBa

[349] R. G. Book review: Business Statistics, by Joseph Lyons Snider. Journal of the Royal Statistical Society, 96 (2):327–329, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341801.

D:1933:BRB

[350] I. D. Book review: The Distribution of Consumable Goods, by Dorothea Braithwaite; S. P. Dobbs. Journal of the Royal Statistical Society, 96(2): 329–331, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341802.

T:1933:BRBb

[351] R. J. T. Book review: World Agriculture: An International Survey. Journal of the Royal Statistical Society, 96 (2):331–332, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341803.

M:1933:BRBd

[352] H. W. M. Book review: The Corn Laws and Social England, by C. R. Fay. Journal of the Royal Statistical Society, 96 (2):332–334, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341804.

T:1933:BRBc

[353] R. J. T. Book review: Agricultural Russia and the Wheat Problem, by Vladimir P. Timoshenko. Journal of the Royal Statistical Society, 96 (2):334–335, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341805.

G:1933:BRBb

[354] R. F. G. Book review: Animal Industry in the British Empire, by A. N. Duckham; J. B. Orr. Journal of the Royal Statistical Society, 96(2):335–336, ???? 1933. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341806.

M:1933:BRBe

[355] H. W. M. Book review: American Population Before the Federal Census of 1790, by Evarts B. Greene; Virginia D. Harrington. Journal of the Royal Statistical Society, 96(2):336, ???? 1933. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341807.

H:1933:BRBc

[356] A. B. H. Book review: American Business Leaders, by F. W. Taussig; Henry Lee; C. S. Joslyn. Journal of the Royal Statistical Society, 96(2):336–337, ????

1933. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341808.

Anonymous:1933:ONPb

[357] Anonymous. Other new publications. Journal of the Royal Statistical Society, 96(2):337–339, ???? 1933. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341809.

Anonymous:1933:CNb

[358] Anonymous. Current notes. Journal of the Royal Statistical Society, 96 (2):340–353, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341810.

G:1933:EMN

[359] M. G. Ethel May newbold. Journal of the Royal Statistical Society, 96 (2):354–357, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341811.

Anonymous:1933:SEAb

[360] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 96 (2):358–364, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341812.

Anonymous:1933:ALb

[361] Anonymous. Additions to the library. Journal of the Royal Statistical Society, 96(2):365–370, ???? 1933. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341813.

Anonymous:1933:PRb

[362] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 96 (2):371–380, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341814.

Anonymous:1933:BMa

[363] Anonymous. Back matter. Journal of the Royal Statistical Society, 96(2):??, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341815.

Anonymous:1933:FMc

[364] Anonymous. Front matter. Journal of the Royal Statistical Society, 96 (3):i–viii, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342115.

Fenelon:1933:BRS

[365] K. G. Fenelon. British Railways since the War. Journal of the Royal Statistical Society, 96(3):381–437, ???? 1933. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342116.

Hawtrey:1933:PET

[366] R. G. Hawtrey. Public expenditure and trade depression. Journal of the Royal Statistical Society, 96(3):438–477, ???? 1933. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342117. See

corrigendum [394].

Yule:1933:NNJ

[367] G. Udny Yule. Note on the number of the Jews in Germany. Journal of the Royal Statistical Society, 96 (3):478–480, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342118.

Martin:1933:EIM

[368] W. J. Martin. The estimation of infant mortality. Journal of the Royal Statistical Society, 96(3):481–486, ???? 1933. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342119.

W:1933:BRBc

[369] J. W. Book review: Approximation and Graduation According to the Principle of Least Squares by Orthogonal Polynomials, by C. Jordan; On the Graduation of Data by the Orthogonal Polynomials of Least Squares, by A. C. Aitken. Journal of the Royal Statistical Society, 96(3):487–491, ???? 1933. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342120.

C:1933:BRBa

[370] L. R. C. Book review: Statistik im Dienste der Technik, by Emil Kohlweiler. Journal of the Royal Statistical Society, 96(3):492, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342121.

C:1933:BRBb

[371] L. R. C. Book review: Forecasting Business Cycles, by Warren M. Persons. Journal of the Royal Statistical Society, 96(3):492–493, ???? 1933. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342122.

W:1933:BRBd

[372] G. R. W. Book review: Business Forecasting and its Practical Application, by William Wallace. Journal of the Royal Statistical Society, 96(3): 493–494, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342123.

W:1933:BRBe

[373] G. R. W. Book review: Booms and Depressions, by Irving Fisher. Journal of the Royal Statistical Society, 96 (3):494–495, ???? 1933. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342124.

W:1933:BRBf

[374] G. R. W. Book review: Business Budgets and Budgetary Control, by A. W. Willsmore. Journal of the Royal Statistical Society, 96(3):495–496, ???? 1933. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2342125.

E:1933:BRBc

[375] W. A. E. Book review: Stable Money. The Remedy for the Economic World Crisis. A Programme of Financial Reconstruction for the International Conference, 1933, by Robert Eisler; Vincent C. Vickers. Journal of the Royal Statistical Society, 96(3):496–498, ???? 1933. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2342126.

T:1933:BRBd

[376] C. T. Book review: The Gold Standard in Theory and Practice, by R. G. Hawtrey. Journal of the Royal Statistical Society, 96(3):498–500, ???? 1933. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2342127.

H:1933:BRBd

[377] R. G. H. Book review: The Portuguese Bank Note Case: The Story and Solution of a Financial Perplexity, by Cecil H. Kisch. Journal of the Royal Statistical Society, 96(3):500–502, ???? 1933. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2342128.

G:1933:BRBc

[378] C. O. G. Book review: British Public Expenditure, 1921–31, by Joseph Sykes. Journal of the Royal Statistical Society, 96(3):502–504, ???? 1933. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342129.

J:1933:BRB

[379] H. S. J. Book review: Taxation of Income in India, by V. K. R. V. Rao. Journal of the Royal Statistical Society, 96(3):504–506, ???? 1933. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342130.

G:1933:BRBd

[380] R. F. G. Book review: International Trade, by P. Barrett Whale. Journal of the Royal Statistical Society, 96 (3):506–507, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342131.

B:1933:BRBa

[381] W. A. B. Book review: International Wage Comparisons. Journal of the Royal Statistical Society, 96(3): 507–508, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342132.

M:1933:BRBf

[382] H. W. M. Book review: The Theory of Wages, by J. R. Hicks. Journal of the Royal Statistical Society, 96 (3):508–509, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342133.

G:1933:BRBe

[383] R. F. G. Book review: Agricultural Credit, by Alexander J. Boyazoglu.

Journal of the Royal Statistical Society, 96(3):509–510, ???? 1933. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342134.

G:1933:BRBf

[384] R. F. G. Book review: The Cost of the War to Russia, by Stanislas Kohn; A. F. Meyendorff. Journal of the Royal Statistical Society, 96(3):510–511, ???? 1933. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342135.

G:1933:BRBg

[385] R. F. G. Book review: Japanese in the United States, by Yamato Ichihashi. Journal of the Royal Statistical Society, 96(3):511–512, ???? 1933. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342136.

M:1933:BRBg

[386] H. W. M. Book review: Essays in Biography, by John Maynard Keynes. Journal of the Royal Statistical Society, 96 (3):512–514, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342137.

Anonymous:1933:ONPc

[387] Anonymous. Other new publications. Journal of the Royal Statistical Society, 96(3):514–518, ???? 1933. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342138.

Anonymous:1933:CNc

[388] Anonymous. Current notes. Journal of the Royal Statistical Society, 96 (3):519–532, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342139.

Anonymous:1933:SEAc

[389] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 96 (3):533–538, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342140.

Anonymous:1933:ALc

[390] Anonymous. Additions to the library. Journal of the Royal Statistical Society, 96(3):539–544, ???? 1933. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342141.

Anonymous:1933:BMb

[391] Anonymous. Back matter. Journal of the Royal Statistical Society, 96 (3):xi–xvi, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342142.

Anonymous:1933:VI

[392] Anonymous. Volume information. Journal of the Royal Statistical Society, 96(4):755–vii, ???? 1933. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341896.

Anonymous:1933:FMd

[393] Anonymous. Front matter. Journal of the Royal Statistical Society, 96 (4):i–viii, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341897.

Anonymous:1933:CPE

[394] Anonymous. Corrigendum: Public expenditure and trade depression. Journal of the Royal Statistical Society, 96(4):??, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341898. See [366].

Jevons:1933:CFI

[395] H. Stanley Jevons. The causes of fluctuations of industrial activity and the price-level. Journal of the Royal Statistical Society, 96(4):545–605, ???? 1933. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2341899.

Flux:1933:MPC

[396] A. W. Flux. The measurement of price changes: Retrospect and prospect. Journal of the Royal Statistical Society, 96(4):606–636, ???? 1933. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341900.

Clarke:1933:POP

[397] R. W. B. Clarke. Production, output per head, prices and costs in the iron and steel industry, 1924–1931. Journal of the Royal Statistical Society, 96 (4):637–650, ???? 1933. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341901.

Clark:1933:NIN

[398] Colin G. Clark. The national income and the net output of industry. Journal of the Royal Statistical Society, 96 (4):651–659, ???? 1933. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341902.

Zaiman:1933:CP

[399] A. Zaiman. Census of Palestine, 1931. Journal of the Royal Statistical Society, 96(4):660–662, ???? 1933. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341903.

Neale:1933:MD

[400] E. P. Neale. Migration and depression. Journal of the Royal Statistical Society, 96(4):663–665, ???? 1933. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341904.

Anonymous:1933:RCF

[401] Anonymous. Report of the Council for the financial year ended December 31, 1932, and for the sessional year ending June 20, 1933, presented at the ninetyninth Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts, John Street, Adelphi, W.C. 2, on June 20, 1933. Journal of the Royal Statistical Society, 96(4):666–677, ???? 1933. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341905.

Anonymous:1933:PNN

[402] Anonymous. Proceedings of the ninetyninth Annual General Meeting of the Royal Statistical Society held in the Hall of the Royal Society of Arts on Tuesday, June 20th, 1933. Journal of the Royal Statistical Society, 96(4): 678–679, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341906.

W:1933:BRBg

[403] J. W. Book review: An Introduction to the Computation of Statistics, by Shepherd Dawson. Journal of the Royal Statistical Society, 96(4):680–682, ???? 1933. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341907.

R:1933:BRBb

[404] E. C. R. Book review: New Methods of Measuring Marginal Utility, by Ragnar Frisch. Journal of the Royal Statistical Society, 96(4):682–683, ???? 1933.

ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2341908.

R:1933:BRBc

[405] E. C. R. Book review: Pitfalls in the Statistical Construction of Demand and Supply Curves, by Ragnar Frisch. Journal of the Royal Statistical Society, 96(4):683–684, ???? 1933. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341909.

B:1933:BRBb

[406] C. H. L. B. Book review: Sources and Characteristics of the Principal Mortality Tables, by J. S. Elston. Journal of the Royal Statistical Society, 96 (4):684–685, ???? 1933. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341910.

H:1933:BRBe

[407] A. B. H. Book review: Tests for Accident Proneness, by E. Farmer; E. G. Chambers; F. J. Kirk. Journal of the Royal Statistical Society, 96(4): 685–686, ???? 1933. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341911.

L:1933:BRBb

[408] H. L. Book review: The Tourist Movement, by F. W. Ogilvie. Journal of the Royal Statistical Society, 96 (4):686–688, ???? 1933. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341912.

T:1933:BRBe

[409] R. J. T. Book review: The Agricultural Situation in 1931–32. Journal of the Royal Statistical Society, 96 (4):688–690, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341913.

M:1933:BRBh

[410] H. W. M. Book review: Economic Scares, by Edwin Cannan. Journal of the Royal Statistical Society, 96(4): 690–691, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341914.

E:1933:BRBd

[411] W. A. E. Book review: The Value of Money, by Tjardus Greidanus. Journal of the Royal Statistical Society, 96 (4):691–693, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341915.

M:1933:BRBi

[412] H. W. M. Book review: Great Britain and the Gold Standard, a Study of the Present World Depression, by H. F. Fraser. Journal of the Royal Statistical Society, 96(4):693–694, ???? 1933. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341916.

M:1933:BRBj

[413] H. W. M. Book review: Money, Gold, Silver and Paper, by Francis W. Hirst. Journal of the Royal Statistical Society, 96(4):694–696, ???? 1933. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341917.

M:1933:BRBk

[414] H. W. M. Book review: This Money Business, A Simple Account of the Institutions and Working of the Banking and Financial World, by Barnard Ellinger. Journal of the Royal Statistical Society, 96(4):696–697, ???? 1933. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341918.

W:1933:BRBh

[415] G. R. W. Book review: Future Trading Upon Organized Commodity Markets in the United States, by G. Wright Hoffman. Journal of the Royal Statistical Society, 96(4):697–699, ???? 1933. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341919.

G:1933:BRBh

[416] R. F. G. Book review: A Contribution to the Study of International Comparisons of Costs of Living. Journal of the Royal Statistical Society, 96 (4):699–700, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341920.

H:1933:BRBf

[417] A. B. H. Book review: The Industries of Greater London, by Douglas H. Smith. Journal of the Royal Statistical Society, 96(4):700–701, ???? 1933. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341921.

B:1933:BRBc

[418] W. A. B. Book review: The Incidence of Work Shortage, by Margaret H. Hogg. Journal of the Royal Statistical Society, 96(4):701–702, ???? 1933. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341922.

Anonymous:1933:ONPd

[419] Anonymous. Other new publications. Journal of the Royal Statistical Society, 96(4):702–706, ???? 1933. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341923.

Anonymous:1933:CNd

[420] Anonymous. Current notes. Journal of the Royal Statistical Society, 96 (4):707–723, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341924.

Anonymous:1933:SEAd

[421] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 96 (4):724–728, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341925.

Anonymous:1933:ALd

[422] Anonymous. Additions to the library. Journal of the Royal Statistical Society, 96(4):729–737, ???? 1933. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341926.

Anonymous:1933:PRL

[423] Anonymous. Periodicals received by the Library. Journal of the Royal Statistical Society, 96(4):738–753, ???? 1933. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341927.

Anonymous:1933:BMc

[424] Anonymous. Back matter. Journal of the Royal Statistical Society, 96 (4):xi–xvi, ???? 1933. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341928.

Anonymous:1934:VIa

[425] Anonymous. Volume information. Journal of the Royal Statistical Society, 97(1):i–vii, ???? 1934. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342013.

Yule:1934:SPR

[426] G. Udny Yule. On some points relating to vital statistics, more especially statistics of occupational mortality. Journal of the Royal Statistical Society, 97(1):1–84, ???? 1934. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342014.

Dale:1934:ISU

[427] J. A. Dale. The interpretation of the statistics of unemployment. Journal of the Royal Statistical Society, 97 (1):85–113, ???? 1934. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342015.

Irwin:1934:RAM

[428] J. O. Irwin. Recent advances in mathematical statistics (1932). Journal of the Royal Statistical Society, 97(1): 114–154, ???? 1934. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342016.

Sinha:1934:IAS

[429] H. Sinha. Indian agricultural statistics. Journal of the Royal Statistical Society, 97(1):155–162, ???? 1934. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342017.

H:1934:BRBa

[430] A. B. H. Book review: The New Survey of London Life and Labour. Journal of the Royal Statistical Society, 97 (1):163–165, ???? 1934. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342018.

R:1934:BRBa

[431] E. C. R. Book review: Prices, by George F. Warren; Frank A. Pearson. Journal of the Royal Statistical Society, 97(1):165–166, ???? 1934. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342019.

M:1934:BRBa

[432] H. W. M. Book review: Monetary Policy and the Depression, by Humphrey Milford. Journal of the Royal Statistical Society, 97(1):166–168, ???? 1934. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342020.

E:1934:BRBa

[433] W. A. E. Book review: Theory of the Credit Standard, by Hans Glückstadt; The Mechanism of the Credit Standard. Practical Proposals for a New Monetary Standard by Hans Glückstadt. Journal of the Royal Statistical Society, 97(1):168–169, ???? 1934. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342021.

E:1934:BRBb

[434] W. A. E. Book review: L’Étalon-Or, en Théorie et en Pratique, by Xénophon Zolotas. Journal of the Royal Statistical Society, 97(1):169–170, ???? 1934. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342022.

E:1934:BRBc

[435] W. A. E. Book review: Essais sur Quelques Problèmes Économiques et Monétaires, by Charles Rist. Journal of the Royal Statistical Society, 97 (1):170–172, ???? 1934. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342023.

R:1934:BRBb

[436] J. H. R. Book review: The Theory of Unemployment, by A. C. Pigou. Journal of the Royal Statistical Society, 97 (1):172–175, ???? 1934. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342024.

G:1934:BRBa

[437] R. G. Book review: Purchasing Power and Trade Depression: A Critique of Under-Consumption Theories, by E. F. M. Durbin. Journal of the Royal Statistical Society, 97(1):175–177, ???? 1934. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342025.

M:1934:BRBb

[438] H. W. M. Book review: Early European Banking in India, with Some Reflections on Present Conditions, by H. Sinha. Journal of the Royal Statistical Society, 97(1):177–178, ???? 1934. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342026.

P:1934:BRB

[439] W. P. Book review: The Logic of Industrial Organization, by P. Sargant Florence; Management of To-morrow, by L. Urwick. Journal of the Royal Statistical Society, 97(1):178–180, ???? 1934. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342027.

M:1934:BRBc

[440] H. W. M. Book review: Australia in the World Depression, by E. Ronald Walker. Journal of the Royal Statistical Society, 97(1):180–182, ???? 1934.

ISSN 0952-8385. URL http://www.

jstor.org/stable/2342028.

T:1934:BRBa

[441] R. J. T. Book review: The Foundations of Agricultural Economics, Together with an Economic History of British Agriculture During and After the Great War, by J. A. Venn. Journal of the Royal Statistical Society, 97 (1):182–183, ???? 1934. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342029.

H:1934:BRBb

[442] A. B. H. Book review: To be or Not to be. A Study of Suicide, by Louis I. Dublin; Bessie Bunzel. Journal of the Royal Statistical Society, 97 (1):183–184, ???? 1934. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342030.

Anonymous:1934:ONPa

[443] Anonymous. Other new publications. Journal of the Royal Statistical Society, 97(1):185–187, ???? 1934. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342031.

Anonymous:1934:CNa

[444] Anonymous. Current notes. Journal of the Royal Statistical Society, 97 (1):188–211, ???? 1934. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342032.

Anonymous:1934:SEH

[445] Anonymous. Sir Edgar Harper. Journal of the Royal Statistical Society, 97(1):212, ???? 1934. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342033.

Anonymous:1934:SEAa

[446] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 97 (1):213–220, ???? 1934. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342034.

Anonymous:1934:LALa

[447] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 97(1):221–232, ???? 1934. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2342035.

Anonymous:1934:PRa

[448] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 97 (1):233–240, ???? 1934. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342036.

Crowther:1934:EIB

[449] Geoffrey Crowther. The economist index of business activity. Journal of the Royal Statistical Society, 97(2): 241–276, ???? 1934. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342388.

Jones:1934:ESV

[450] J. H. Jones. Exchange stability versus internal price stability. Journal of the Royal Statistical Society, 97(2): 277–312, ???? 1934. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342389.

Anonymous:1934:WP

[451] Anonymous. Wholesale prices in 1933. Journal of the Royal Statistical Society, 97(2):313–330, ???? 1934. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342390.

Elderton:1934:ICF

[452] W. Palin Elderton and G. H. Hansmann. Improvement of curves fitted by the method of moments. Journal of the Royal Statistical Society, 97 (2):331–333, ???? 1934. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342391.

Ba:1934:BRB

[453] M. S. Ba. Book review: Researches Into the Theory of Regression, by Walter Anderson. Journal of the Royal Statistical Society, 97(2):334–336, ???? 1934. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342392.

W:1934:BRBa

[454] J. W. Book review: Fertility and Reproduction. Methods of Measuring the Balance of Births and Deaths, by R. R. Kuczynski. Journal of the Royal Statistical Society, 97(2):336, ???? 1934. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342393.

W:1934:BRBb

[455] J. W. Book review: Elementary Statistical Methods, by E. C. Rhodes. Journal of the Royal Statistical Society, 97 (2):337–338, ???? 1934. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342394.

H:1934:BRBc

[456] R. G. H. Book review: L’Or et sa Distribution Mondiale, by Albert Aftalion. Journal of the Royal Statistical Society, 97(2):338–341, ???? 1934. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342395.

E:1934:BRBd

[457] W. A. E. Book review: Gold Unemployment and Capitalism, by T. E. Gregory; Ernest Cassel. Journal of the Royal Statistical Society, 97(2): 342–343, ???? 1934. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342396.

W:1934:BRBc

[458] H. W. M. Book review: An Essay on the Nature and Significance of Economic Science, by Lionel Robbins. Journal of the Royal Statistical Society, 97(2):343–344, ???? 1934. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342397.

M:1934:BRBd

[459] H. W. M. Book review: Economics of a Changing World, by H. V. Hodson. Journal of the Royal Statistical Society, 97(2):345–346, ???? 1934. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342398.

M:1934:BRBe

[460] H. W. M. Book review: Studies in World Economics, by G. D. H. Cole. Journal of the Royal Statistical Society, 97(2):346–348, ???? 1934. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342399.

M:1934:BRBf

[461] H. W. M. Book review: International Combines in Modern Industry, by Alfred Plummer. Journal of the Royal Statistical Society, 97(2):348–349, ????

1934. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342400.

V:1934:BRB

[462] J. W. V. Book review: Industrial Relations in Great Britain. Journal of the Royal Statistical Society, 97(2): 349–350, ???? 1934. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342401.

W:1934:BRBd

[463] N. W. Book review: Fundamentals of Industrial Administration, by Edward Tregaskiss Elbourne. Journal of the Royal Statistical Society, 97(2): 350–351, ???? 1934. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342402.

W:1934:BRBe

[464] N. W. Book review: Australia as Producer and Trader, 1920–1932, by Nancy Windett. Journal of the Royal Statistical Society, 97(2):351–353, ???? 1934. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342403.

Anonymous:1934:ONPb

[465] Anonymous. Other new publications. Journal of the Royal Statistical Society, 97(2):353–355, ???? 1934. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342404.

Anonymous:1934:CNb

[466] Anonymous. Current notes. Journal of the Royal Statistical Society, 97 (2):356–371, ???? 1934. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342405.

Anonymous:1934:SEAb

[467] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 97 (2):372–379, ???? 1934. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342406.

Anonymous:1934:LALb

[468] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 97(2):380–388, ???? 1934. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2342407.

Anonymous:1934:PRb

[469] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 97 (2):389–398, ???? 1934. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342408.

Lindsay:1934:ITI

[470] H. A. F. Lindsay. India’s trade and industrial statistics, past, present and future. Journal of the Royal Statistical Society, 97(3):399–422, ???? 1934. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2342256.

Stamp:1934:MUD

[471] Josiah Stamp. Methods used in different countries for estimating national income. Journal of the Royal Statistical Society, 97(3):423–466, ???? 1934. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2342257.

Anonymous:1934:CRS

[472] Anonymous. The centenary of the Royal Statistical Society. Journal of the Royal Statistical Society, 97(3):

467–472, ???? 1934. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342258.

Bowley:1934:TSS

[473] A. L. Bowley. The twenty-second session of the International Institute of Statistics. Journal of the Royal Statistical Society, 97(3):473–477, ???? 1934. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342259.

Jones:1934:IPS

[474] J. H. Jones and Robert Eisler. Internal price stability versus exchange stability: A note on Professor J. H. Jones’s paper. Journal of the Royal Statistical Society, 97(3):478–483, ???? 1934. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342260.

M:1934:BRBg

[475] H. W. M. Book review: 50 Années de l’Institut International de Statistique, by Friedrich Zahn. Journal of the Royal Statistical Society, 97(3):484–486, ???? 1934. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342261.

M:1934:BRBh

[476] H. W. M. Book review: Economic and Social Investigations in Manchester, 1833–1933; A Centenary History of the Manchester Statistical Society, by T. S. Ashton. Journal of the Royal Statistical Society, 97(3):487–488, ???? 1934. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342262.

C-S:1934:BRB

[477] A. M. C-S. Book review: Census of India, 1931, by J. H. Hutton. Journal of the Royal Statistical Society, 97 (3):488–490, ???? 1934. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342263.

H:1934:BRBd

[478] A. B. H. Book review: The New Survey of London Life and Labour. Journal of the Royal Statistical Society, 97 (3):490–492, ???? 1934. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342264.

M:1934:BRBi

[479] H. W. M. Book review: Walks and Talks: The Diary of a Member of Parliament in 1933–4, by Arnold Wilson. Journal of the Royal Statistical Society, 97(3):492–494, ???? 1934. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342265.

E:1934:BRBe

[480] W. A. E. Book review: Trade Depression and the Way Out, by R. G. Hawtrey. Journal of the Royal Statistical Society, 97(3):494–496, ???? 1934. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342266.

E:1934:BRBf

[481] W. A. E. Book review: Gold Reserves and the Monetary Standard, with Special Reference to the Position of the Bank of England: and an Outline of a Proposed Monetary System for the British Commonwealth of Nations, by D. S. Edwards. Journal of the Royal Statistical Society, 97(3):496–497, ???? 1934. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342267.

L:1934:BRB

[482] H. M. L. Book review: Das Deutsche Volkseinkommen vor und nach dem Kriege ; Wolkswohlstand und Volkseinkommen, by M. J. Elsas. Journal of the Royal Statistical Society, 97 (3):498–500, ???? 1934. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342268.

M:1934:BRBj

[483] H. W. M. Book review: The Economics of Saving, by J. Harry Jones. Journal of the Royal Statistical Society, 97 (3):500–501, ???? 1934. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342269.

W:1934:BRBf

[484] G. R. W. Book review: Cartels, Concerns and Trusts, by Robert Liefmann. Journal of the Royal Statistical Society, 97(3):501–502, ???? 1934. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342270.

B:1934:BRB

[485] H. B. Book review: Building Society Finance, by J. L. Cohen. Journal of the Royal Statistical Society, 97 (3):502–504, ???? 1934. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342271.

T:1934:BRBb

[486] R. J. T. Book review: The Rural Exodus in Germany, by International Labour Office; International Institute of Agriculture. Journal of the Royal Statistical Society, 97(3):504–506, ???? 1934. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342272.

T:1934:BRBc

[487] R. J. T. Book review: The Tariff on Sugar, by Lippert S. Ellis; The Tariff on Dairy Products, by Roland R.

Renne; The Tariff on Barley, Oats and Corn, by Theo. W. Schultz. Journal of the Royal Statistical Society, 97 (3):506–507, ???? 1934. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342273.

G:1934:BRBb

[488] E. H. G. Book review: Communications in the Far East, by Frederick Vincent de Fellner. Journal of the Royal Statistical Society, 97(3):507–510, ???? 1934. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2342274.

Anonymous:1934:ONPc

[489] Anonymous. Other new publications. Journal of the Royal Statistical Society, 97(3):510–514, ???? 1934. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342275.

Anonymous:1934:CNc

[490] Anonymous. Current notes. Journal of the Royal Statistical Society, 97 (3):515–529, ???? 1934. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342276.

Anonymous:1934:SEAc

[491] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 97 (3):530–535, ???? 1934. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342277.

Anonymous:1934:LALc

[492] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 97(3):536–540, ???? 1934. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2342278.

Anonymous:1934:VIb

[493] Anonymous. Volume information. Journal of the Royal Statistical Society, 97(4):720–724, ???? 1934. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342190.

Anonymous:1934:DMU

[494] Anonymous. Discussion on methods used in different countries for estimating national income. Journal of the Royal Statistical Society, 97(4): 541–557, ???? 1934. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342191.

Neyman:1934:TDA

[495] Jerzy Neyman. On the two different aspects of the representative method: The method of stratified sampling and the method of purposive selection. Journal of the Royal Statistical Society, 97(4):558–625, ???? 1934. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342192.

Campion:1934:PWF

[496] H. Campion. Pre-war fluctuations of profits in the cotton-spinning industry. Journal of the Royal Statistical Society, 97(4):626–632, ???? 1934. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342193.

Rowson:1934:VRA

[497] S. Rowson. The value of remittances abroad for cinematograph films. Journal of the Royal Statistical Society, 97 (4):633–640, ???? 1934. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342194.

Anonymous:1934:RCF

[498] Anonymous. Report of the Council for the financial year ended December 31, 1933, and for the sessional year ending June 19, 1934, presented at the one hundredth Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts, John Street, Adelphi, W.C. 2, on June 19, 1934. Journal of the Royal Statistical Society, 97(4):641–653, ???? 1934. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342195.

Anonymous:1934:POH

[499] Anonymous. Proceedings of the onehundredth Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts on Tuesday, June 19th, 1934. Journal of the Royal Statistical Society, 97 (4):654–657, ???? 1934. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342196.

K:1934:BRB

[500] O. C. J. K. Book review: Applications de la Statistique à la Démographie et à la Biologie, by R. Risser; Emile Borel. Journal of the Royal Statistical Society, 97(4):658–660, ???? 1934. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342197.

G:1934:BRBc

[501] R. G. Book review: Production Trends in the U.S.A. since 1870, by Arthur F. Burns. Journal of the Royal Statistical Society, 97(4):660–661, ???? 1934. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342198.

G:1934:BRBd

[502] R. G. Book review: German Business Cycles, 1924–1933, by Carl T. Schmidt. Journal of the Royal Statistical Society, 97(4):661–662, ???? 1934. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342199.

M:1934:BRBk

[503] H. W. M. Book review: Economic Conditions in France, by Robert Cahill. Journal of the Royal Statistical Society, 97(4):662–665, ???? 1934. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342200.

H:1934:BRBe

[504] A. B. H. Book review: Becontree and Dagenham: The Story of the Growth of a Housing Estate, by Terence Young; Stanley Baldwin. Journal of the Royal Statistical Society, 97(4):665–666, ???? 1934. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342201.

H:1934:BRBf

[505] R. G. H. Book review: The Great Depression, by L. Robbins. Journal of the Royal Statistical Society, 97(4): 666–671, ???? 1934. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342202.

H:1934:BRBg

[506] R. G. H. Book review: The Economics of Imperfect Competition, by Joan Robinson. Journal of the Royal Statistical Society, 97(4):671–674, ???? 1934. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342203.

W:1934:BRBg

[507] G. R. W. Book review: Money, Business and Prices, by J. P. Wernette. Journal of the Royal Statistical Society, 97(4):674–675, ???? 1934. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342204.

W:1934:BRBh

[508] G. R. W. Book review: The Rate of Interest in a Progressive State, by J. E. Meade. Journal of the Royal Statistical Society, 97(4):675–676, ???? 1934. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2342205.

Anonymous:1934:ONPd

[509] Anonymous. Other new publications. Journal of the Royal Statistical Society, 97(4):676–678, ???? 1934. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342206.

Anonymous:1934:CNd

[510] Anonymous. Current notes. Journal of the Royal Statistical Society, 97 (4):679–692, ???? 1934. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342207.

Anonymous:1934:HRB

[511] Anonymous. Henry Ramie Beeton. Journal of the Royal Statistical Society, 97(4):693–694, ???? 1934. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342208.

Anonymous:1934:SEAd

[512] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 97 (4):695–700, ???? 1934. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342209.

Anonymous:1934:LALd

[513] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 97(4):701–705, ???? 1934. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342210.

Anonymous:1934:PRL

[514] Anonymous. Periodicals received by the Library. Journal of the Royal Statistical Society, 97(4):706–719, ???? 1934. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342211.

Anonymous:1935:FMa

[515] Anonymous. Front matter. Journal of the Royal Statistical Society, 98 (1):i–viii, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342432.

Anonymous:1935:UER

[516] Anonymous. University education: Its recent history and function. Journal of the Royal Statistical Society, 98 (1):1–33, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342433.

Anonymous:1935:PM

[517] Anonymous. Proceedings of the meeting. Journal of the Royal Statistical Society, 98(1):33–38, ???? 1935. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342434.

Fisher:1935:LII

[518] R. A. Fisher. The logic of inductive inference. Journal of the Royal Statistical Society, 98(1):39–82, ???? 1935. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342435.

Irwin:1935:RAM

[519] J. O. Irwin. Recent advances in mathematical statistics (1933). Journal of the Royal Statistical Society, 98 (1):83–127, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342436.

Koshal:1935:AMM

[520] R. S. Koshal. Application of the method of maximum likelihood to the derivation of efficient statistics for fitting frequency curves. Journal of the Royal Statistical Society, 98(1):128, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342437.

Subramanian:1935:PPC

[521] S. Subramanian. On a property of partial correlation. Journal of the Royal Statistical Society, 98(1):129, ???? 1935. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2342438.

M:1935:NBT

[522] H. W. M. The new board of trade wholesale prices index number. Journal of the Royal Statistical Society, 98 (1):130–133, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342439.

Shaul:1935:CVA

[523] J. R. H. Shaul. Changes in the volume of the agricultural output of Scotland. Journal of the Royal Statistical Society, 98(1):134–139, ???? 1935. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342440.

Anonymous:1935:RSS

[524] Anonymous. The Royal Statistical Society. Journal of the Royal Statistical Society, 98(1):140–151, ???? 1935. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342441.

H:1935:BRBa

[525] A. B. H. Book review: The New Survey of London Life and Labour. Journal of the Royal Statistical Society, 98 (1):152–154, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342442.

W:1935:BRBa

[526] N. W. Book review: Work and Wealth in a Modern Port: An Economic Survey of Southampton, by P. Ford. Journal of the Royal Statistical Society, 98 (1):154–156, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342443.

T:1935:BRBa

[527] R. J. T. Book review: International Year Book of Agricultural Statistics, 1932–33. Journal of the Royal Statistical Society, 98(1):156–157, ???? 1935. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342444.

T:1935:BRBb

[528] R. J. T. Book review: Le Problème Mondiale du Blé, by Paul de Hevesy. Journal of the Royal Statistical Society, 98(1):157–158, ???? 1935. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342445.

M:1935:BRBa

[529] H. W. M. Book review: Population Theories and their Applications, by E.

F. Penrose. Journal of the Royal Statistical Society, 98(1):158–160, ???? 1935. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342446.

B:1935:BRBa

[530] A. L. B. Book review: Wages in Eighteenth-Century England, by Elizabeth W. Gilboy. Journal of the Royal Statistical Society, 98(1):160–161, ???? 1935. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342447.

M:1935:BRBb

[531] H. W. M. Book review: Economic Reconstruction. Journal of the Royal Statistical Society, 98(1):161–163, ???? 1935. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342448.

M:1935:BRBc

[532] H. W. M. Book review: Some Relations Between Political and Economic Theory, by G. D. H. Cole. Journal of the Royal Statistical Society, 98 (1):164–166, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342449.

S:1935:BRBa

[533] E. C. S. Book review: Enterprise, Purpose and Profit, by D. H. MacGregor. Journal of the Royal Statistical Society, 98(1):166–169, ???? 1935. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342450.

M:1935:BRBd

[534] H. W. M. Book review: Theories of the Trade Cycle, by Alec L. Macfie. Journal of the Royal Statistical Society, 98 (1):169–172, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342451.

R:1935:BRBa

[535] M. A. R. Book review: The Theory of Profits, by P. Ch. Ghosh. Journal of the Royal Statistical Society, 98 (1):172–173, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342452.

G:1935:BRBa

[536] C. O. G. Book review: Unbalanced Budgets, by Hugh Dalton; Brinley Thomas; J. N. Reedman; T. J. Hughes; W. J. Leaning. Journal of the Royal Statistical Society, 98(1):173–175, ???? 1935. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342453.

L:1935:BRB

[537] H. M. L. Book review: The French International Accounts, 1880–1913, by Harry D. White. Journal of the Royal Statistical Society, 98(1):175–177, ???? 1935. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342454.

E:1935:BRBa

[538] W. A. E. Book review: L’Inflation du Crédit et la Prévention des Crises, by Maurice Ansiaux. Journal of the Royal Statistical Society, 98(1):177–178, ???? 1935. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342455.

M:1935:BRBe

[539] H. W. M. Book review: The Nemesis of Economic Nationalism and Other Lectures in Applied Economics, by Joseph Johnston. Journal of the Royal Statistical Society, 98(1):178–180, ???? 1935.

ISSN 0952-8385. URL http://www.

jstor.org/stable/2342456.

W:1935:BRBb

[540] K. W. Book review: Prospettive Economiche, by Giorgio Mortara. Journal of the Royal Statistical Society, 98 (1):180–181, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342457.

J:1935:BRB

[541] H. S. J. Book review: Industrial Policy of India, with Special Reference to Customs Tariff, by C. N. Vakil; M. C. Munshi. Journal of the Royal Statistical Society, 98(1):181–182, ???? 1935. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342458.

R:1935:BRBb

[542] M. A. R. Book review: Banking in India, by S. G. Panandikar. Journal of the Royal Statistical Society, 98 (1):182–183, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342459.

I:1935:BRBa

[543] J. O. I. Book review: British Association for the Advancement of Science — Mathematical Tables: Vol. III. Minimum Decompositions into Fifth Powers, by L. E. Dickson; British Association for the Advancement of Science — Mathematical Tables: Vol. IV. Cycles of Reduced Ideals in Quadratic Fields, by E. L. Ince. Journal of the Royal Statistical Society, 98(1):184–185, ???? 1935. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342460.

Anonymous:1935:ONPa

[544] Anonymous. Other new publications. Journal of the Royal Statistical Society, 98(1):185–187, ???? 1935. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342461.

Anonymous:1935:SNa

[545] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 98 (1):188–204, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342462.

Anonymous:1935:CNa

[546] Anonymous. Current notes. Journal of the Royal Statistical Society, 98 (1):205–208, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342463.

Anonymous:1935:SEAa

[547] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 98 (1):209–215, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342464.

Anonymous:1935:LALa

[548] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 98(1):216–229, ???? 1935. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342465.

Anonymous:1935:PRa

[549] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 98 (1):230–237, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342466.

Anonymous:1935:BMa

[550] Anonymous. Back matter. Journal of the Royal Statistical Society, 98(1):??, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342467.

Anonymous:1935:FMb

[551] Anonymous. Front matter. Journal of the Royal Statistical Society, 98 (2):i–viii, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341983.

Snow:1935:LIE

[552] E. C. Snow. The limits of industrial employment (II). the influence of growth of population on the development of industry. Journal of the Royal Statistical Society, 98(2):239–292, ???? 1935. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341984.

Fenelon:1935:RTG

[553] K. G. Fenelon. Road transport in Great Britain since the War. Journal of the Royal Statistical Society, 98 (2):293–346, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341985.

Anonymous:1935:WP

[554] Anonymous. Wholesale prices in 1934. Journal of the Royal Statistical Society, 98(2):347–362, ???? 1935. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341986.

Bowley:1935:NCW

[555] A. L. Bowley. Number of children in working-class families in London, 1929– 30. Journal of the Royal Statistical Society, 98(2):363–375, ???? 1935. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341987.

Anonymous:1935:OTR

[556] Anonymous. Obituary: Thomas Robert Malthus: Died 29 December, 1834. Journal of the Royal Statistical Society, 98(2):376–379, ???? 1935. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341988.

Anonymous:1935:CC

[557] Anonymous. Calendar committee. Journal of the Royal Statistical Society, 98(2):380–382, ???? 1935. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341989.

G:1935:BRBb

[558] M. G. Book review: Isaac Newton: A Biography, by Louis Trenchard More. Journal of the Royal Statistical Society, 98(2):383–386, ???? 1935. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341990.

I:1935:BRBb

[559] J. O. I. Book review: Tables of the Higher Mathematical Functions, Vol. 1, by H. T. Davis. Journal of the Royal Statistical Society, 98(2):386–387, ???? 1935. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341991.

S:1935:BRBb

[560] J. C. S. Book review: Industrial Profits in the United States, by Ralph C. Epstein; Florence M. Clark. Journal of the Royal Statistical Society, 98 (2):388–390, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341992.

R:1935:BRBc

[561] M. A. R. Book review: Economic Tendencies in the United States, by Frederick C. Mills. Journal of the Royal Statistical Society, 98(2):390–391, ???? 1935. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341993.

M:1935:BRBf

[562] H. W. M. Book review: The Gold Standard and Its Future, by T. E. Gregory; E. Cassel. Journal of the Royal Statistical Society, 98(2):391–392, ???? 1935. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2341994.

M:1935:BRBg

[563] H. W. M. Book review: About Money, by Erich Roll. Journal of the Royal Statistical Society, 98(2):392–394, ???? 1935. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341995.

W:1935:BRBc

[564] K. W. Book review: I Trasferimenti Anormali dei Capitali e le Crisi, by Marco Fanno. Journal of the Royal Statistical Society, 98(2):394–395, ???? 1935. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341996.

M:1935:BRBh

[565] H. W. M. Book review: Credit and International Trade: How they Work in Practice, by Barnard Ellinger. Journal of the Royal Statistical Society, 98 (2):395–397, ???? 1935. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341997.

M:1935:BRBi

[566] H. W. M. Book review: The Road to Planned Economy: Capitalism and Socialism in Germany’s Development, by Werner Friedrich Bruck. Journal of the Royal Statistical Society, 98(2): 397–399, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341998.

E:1935:BRBb

[567] W. A. E. Book review: Wages Policy and the Price Level, by K. S. Isles; Adam Smith Prizeman. Journal of the Royal Statistical Society, 98 (2):399–400, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341999.

M:1935:BRBj

[568] H. W. M. Book review: The Foundation of Industrial Stability, by E. M. Daltroff. Journal of the Royal Statistical Society, 98(2):400–401, ???? 1935. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342000.

R:1935:BRBd

[569] M. A. R. Book review: The Theory of Wages, by Paul H. Douglas. Journal of the Royal Statistical Society, 98 (2):401–403, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342001.

B:1935:BRBb

[570] W. A. B. Book review: The Employment Exchange Service of Great Britain, by T. S. Chegwidden; G. Myrddin-Evans; Winston Churchill. Journal of the Royal Statistical Society, 98(2):403–405, ???? 1935. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342002.

M:1935:BRBk

[571] H. W. M. Book review: Miners and Management, by Mary Van Kleeck. Journal of the Royal Statistical Society, 98(2):405–406, ???? 1935. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342003.

T:1935:BRBc

[572] R. J. T. Book review: The Planning of Britain’s Food Imports, by K. A. H. Murray; Ruth L. Cohen. Journal of the Royal Statistical Society, 98 (2):406–407, ???? 1935. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342004.

T:1935:BRBd

[573] R. J. T. Book review: Agricultural Cooperation in Fascist Italy, by Freppel Cotta; C. R. Fay. Journal of the Royal Statistical Society, 98(2):407–408, ???? 1935. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2342005.

Anonymous:1935:ONPb

[574] Anonymous. Other new publications. Journal of the Royal Statistical Society, 98(2):408–409, ???? 1935. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342006.

Anonymous:1935:SNb

[575] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 98 (2):410–425, ???? 1935. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342007.

Anonymous:1935:CNb

[576] Anonymous. Current notes. Journal of the Royal Statistical Society, 98 (2):426–430, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342008.

Anonymous:1935:SEAb

[577] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 98 (2):431–437, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342009.

Anonymous:1935:LALb

[578] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 98(2):438–444, ???? 1935. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342010.

Anonymous:1935:PRb

[579] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 98 (2):445–454, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342011.

Anonymous:1935:BMb

[580] Anonymous. Back matter. Journal of the Royal Statistical Society, 98(2):??, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342012.

Anonymous:1935:FMc

[581] Anonymous. Front matter. Journal of the Royal Statistical Society, 98 (3):i–viii, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342279.

Douglas:1935:RTS

[582] Iris Douglas. Retail trade statistics in different countries. Journal of the Royal Statistical Society, 98(3): 455–496, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342280.

Glenday:1935:UME

[583] Roy Glenday and M. Greenwood. The use and misuse of economic statistics. Journal of the Royal Statistical Society, 98(3):497–522, ???? 1935. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342281.

Anonymous:1935:MRQ

[584] Anonymous. Memorandum regarding a quinquennial census. Journal of the Royal Statistical Society, 98(3): 523–530, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342282.

Ellinger:1935:MMS

[585] Barnard Ellinger, H. Campton, and H. G. Hughes. Memorandum by the Manchester Statistical Society on the for improvement and extension of official statistics. Journal of the Royal Statistical Society, 98(3):531–535, ???? 1935. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342283.

Bartlett:1935:SAT

[586] M. S. Bartlett. Some aspects of the time-correlation problem in regard to tests of significance. Journal of the Royal Statistical Society, 98 (3):536–543, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342284.

Todd:1935:PME

[587] G. F. Todd. Professor MacGregor on enterprise and the trade cycle. Journal of the Royal Statistical Society, 98 (3):544–552, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342285.

B:1935:BRBc

[588] C. H. L. B. Book review: Continuous Investigation into the Mortality of Assured Lives: (a) Statistics for the Six Years 1924–9 ; Continuous Investigation into the Mortality of Assured Lives: (b) Monetary Tables. Journal of the Royal Statistical Society, 98 (3):553–555, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342286.

B:1935:BRBd

[589] C. H. L. B. Book review: Reversions and Life Interests, by H. J. Tappenden. Journal of the Royal Statistical Society, 98(3):555–556, ???? 1935. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342287.

B:1935:BRBe

[590] M. B. Book review: Administration of Placement and Unemployment Insurance in Germany, by Oscar Weigert. Journal of the Royal Statistical Society, 98(3):556–558, ???? 1935. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342288.

B:1935:BRBf

[591] W. A. B. Book review: Social and Economic Reconstruction in the United States. Journal of the Royal Statistical Society, 98(3):558–559, ???? 1935. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342289.

G:1935:BRBc

[592] A. T. K. G. Book review: Studies in Capital and Investment, by G. D. H.

Cole. Journal of the Royal Statistical Society, 98(3):559–561, ???? 1935. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2342290.

M:1935:BRBl

[593] H. W. M. Book review: The Exchange Equalization Account, by N. F. Hall. Journal of the Royal Statistical Society, 98(3):561–563, ???? 1935. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342291.

M:1935:BRBm

[594] H. W. M. Book review: Economics in Practice, Six Lectures on Current Issues, by A. C. Pigon. Journal of the Royal Statistical Society, 98(3): 563–564, ???? 1935. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342292.

H:1935:BRBb

[595] R. G. H. Book review: The Economic Consequences of Progress, by Roy Glenday. Journal of the Royal Statistical Society, 98(3):564–566, ???? 1935. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2342293.

M:1935:BRBn

[596] H. W. M. Book review: Principles of Economic Planning, by G. D. H. Cole; Pleasures of Planning, by I. M. Horobin; A Planned Nationalism, Canada’s Effort, by E. S. Bates. Journal of the Royal Statistical Society, 98 (3):566–568, ???? 1935. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342294.

T:1935:BRBe

[597] C. T. Book review: The Correct Economy for the Machine Age, by A. G.

McGregor; John A. Hobson. Journal of the Royal Statistical Society, 98(3):568, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342295.

F:1935:BRB

[598] R. A. F. Book review: Heredity, Mainly Human, by Eldon Moore. Journal of the Royal Statistical Society, 98(3):569, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342296.

Anonymous:1935:ONPc

[599] Anonymous. Other new publications. Journal of the Royal Statistical Society, 98(3):569–571, ???? 1935. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342297.

Anonymous:1935:SNc

[600] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 98 (3):572–584, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342298.

Anonymous:1935:CNc

[601] Anonymous. Current notes. Journal of the Royal Statistical Society, 98 (3):585–586, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342299.

Anonymous:1935:SEAc

[602] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 98 (3):587–592, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342300.

Anonymous:1935:LALc

[603] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 98(3):593–599, ???? 1935. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2342301.

Anonymous:1935:VI

[604] Anonymous. Volume information. Journal of the Royal Statistical Society, 98(4):804–vii, ???? 1935. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342302.

Anonymous:1935:FMd

[605] Anonymous. Front matter. Journal of the Royal Statistical Society, 98 (4):i–viii, ???? 1935. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342303.

Carslaw:1935:LBO

[606] R. McG. Carslaw and P. E. Graves. The labour bill and output on arable farms. Journal of the Royal Statistical Society, 98(4):601–637, ???? 1935. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2342304.

Anonymous:1935:POM

[607] Anonymous. Proceedings of the ordinary meeting of the Royal Statistical Society, Tuesday, June 18th, 1935, the President, Professor M. Greenwood, F.R.S., in the chair. Journal of the Royal Statistical Society, 98 (4):638–639, ???? 1935. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342305.

Ramsbottom:1935:CWR

[608] E. C. Ramsbottom. The course of wage rates in the United Kingdom, 1921– 1934. Journal of the Royal Statistical Society, 98(4):639–694, ???? 1935. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342306.

Grunwald:1935:LRG

[609] Kurt Grunwald. On the law of returns in a growing industrial economy. Journal of the Royal Statistical Society, 98 (4):695–704, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342307.

Anonymous:1935:RCF

[610] Anonymous. Report of the Council for the financial year ended December 31, 1934, and for the sessional year ending June 18, 1935, presented at the one hundred and first Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts, John Street, Adelphi, W.C. 2, on June 18, 1935. Journal of the Royal Statistical Society, 98 (4):705–718, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342308.

Anonymous:1935:POH

[611] Anonymous. Proceedings of the one hundred and first Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts on Tuesday, June 18th, 1935. Journal of the Royal Statistical Society, 98(4):719, ???? 1935. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342309.

M:1935:BRBo

[612] H. W. M. Book review: Epidemics and Crowd Diseases, by Major Greenwood. Journal of the Royal Statistical Society, 98(4):720–724, ???? 1935. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342310.

H:1935:BRBc

[613] A. B. H. Book review: The New Survey of London Life and Labour. Journal of the Royal Statistical Society, 98 (4):724–726, ???? 1935. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342311.

B:1935:BRBg

[614] A. L. B. Book review: Dynamic Economics: Theoretical and Statistical Studies of Demand, Production, and Prices, by Charles Frederick Roos. Journal of the Royal Statistical Society, 98(4):726–728, ???? 1935. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2342312.

M:1935:BRBp

[615] H. W. M. Book review: Monetary Policy and Economic Stabilisation, by Arthur D. Gayer. Journal of the Royal Statistical Society, 98(4):728–731, ???? 1935. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2342313.

M:1935:BRBq

[616] H. W. M. Book review: The Future of Monetary Policy, by Royal Institute of International Affairs. Journal of the Royal Statistical Society, 98 (4):731–734, ???? 1935. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342314.

M:1935:BRBr

[617] H. W. M. Book review: Monetary Opinions and Policy 1924–34, by Mary Theresa Rankin; The Price of Gold by I. Shrigley. Journal of the Royal Statistical Society, 98(4):734–735, ???? 1935. ISSN 0952-8385. URL http://www.

jstor.org/stable/2342315.

E:1935:BRBc

[618] W. A. E. Book review: Treasuries and Central Banks, by David William Dodwell; Basil P. Blackett. Journal of the Royal Statistical Society, 98(4): 735–737, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342316.

E:1935:BRBd

[619] W. A. E. Book review: Free Banking. An Outline of a Policy of Individualism, by Henry Meulen. Journal of the Royal Statistical Society, 98 (4):737–738, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342317.

E:1935:BRBe

[620] W. A. E. Book review: The Pound’s Progress. British Financial Policy and Anglo-American Relations, by F. J. Scanlan. Journal of the Royal Statistical Society, 98(4):739, ???? 1935. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342318.

W:1935:BRBd

[621] G. R. W. Book review: Interregional and International Trade, by Bertil Ohlin. Journal of the Royal Statistical Society, 98(4):739–740, ???? 1935. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342319.

M:1935:BRBs

[622] H. W. M. Book review: French Import Quotas. A New Instrument of Commercial Policy, by F. A. Haight. Journal of the Royal Statistical Society, 98 (4):740–742, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342320.

G:1935:BRBd

[623] C. O. G. Book review: Principles and Methods of Taxation, by G. Armitage-Smith; R. G. Hawtrey. Journal of the Royal Statistical Society, 98(4): 742–743, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342321.

T:1935:BRBf

[624] R. J. T. Book review: The World Agricultural Situation in 1933–34. Journal of the Royal Statistical Society, 98 (4):743–744, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342322.

M:1935:BRBt

[625] H. W. M. Book review: An Introduction to the Study of Prices, by Walter T. Layton; Geoffrey Crowther. Journal of the Royal Statistical Society, 98 (4):744–745, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342323.

G:1935:BRBe

[626] R. F. G. Book review: Simple Economics, by F. H. Spencer. Journal of the Royal Statistical Society, 98(4): 745–746, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342324.

G:1935:BRBf

[627] R. G. Book review: Mechanization in Industry, by Harry Jerome. Journal of the Royal Statistical Society, 98 (4):746–747, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342325.

B:1935:BRBh

[628] M. B. Book review: From Peter The Great to Lenin, by S. P. Turin. Journal of the Royal Statistical Society, 98 (4):747–748, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342326.

Anonymous:1935:ONPd

[629] Anonymous. Other new publications. Journal of the Royal Statistical Society, 98(4):748–753, ???? 1935. ISSN

0952-8385. URL http://www.jstor.
org/stable/2342327.

Anonymous:1935:SNd

[630] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 98 (4):754–769, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342328.

Anonymous:1935:CNd

[631] Anonymous. Current notes. Journal of the Royal Statistical Society, 98 (4):770–774, ???? 1935. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2342329.

Hawtrey:1935:SBB

[632] R. G. Hawtrey. Sir Basil Blackett, K.C.B., K.C.S.I. Journal of the Royal Statistical Society, 98(4):775–777, ???? 1935. ISSN 0952-8385. URL http://

www.jstor.org/stable/2342330.

Anonymous:1935:WH

[633] Anonymous. Wynnard Hooper. Journal of the Royal Statistical Society, 98 (4):777–778, ???? 1935. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342331.

Anonymous:1935:SEAd

[634] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 98 (4):779–783, ???? 1935. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2342332.

Anonymous:1935:LALd

[635] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 98(4):784–789, ???? 1935. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2342333.

Anonymous:1935:PRL

[636] Anonymous. Periodicals received by the Library. Journal of the Royal Statistical Society, 98(4):790–803, ???? 1935. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2342334.

Anonymous:1936:FMa

[637] Anonymous. Front matter. Journal of the Royal Statistical Society, 99 (1):i–viii, ???? 1936. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980310.

Connor:1936:UHE

[638] L. R. Connor. Urban housing in England and Wales. Journal of the Royal Statistical Society, 99(1):1–66, ???? 1936. ISSN 0952-8385. URL http:/

                                        /www.jstor.org/stable/2980311.

Rowson:1936:SSC

[639] S. Rowson. A statistical survey of the cinema industry in Great Britain in 1934. Journal of the Royal Statistical Society, 99(1):67–129, ???? 1936. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980312.

Isserlis:1936:IP

[640] L. Isserlis. Inverse probability. Journal of the Royal Statistical Society, 99 (1):130–137, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980313.

Linder:1936:DBJ

[641] A. Linder. Daniel Bernoulli and J. H. Lambert on mortality statistics. Journal of the Royal Statistical Society, 99 (1):138–141, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980314.

Rhodes:1936:PINa

[642] E. C. Rhodes. The precision of index numbers. Journal of the Royal Statistical Society, 99(1):142–146, ???? 1936. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980315.

George:1936:SIP

[643] R. F. George. A sample investigation of the 1931 population census with reference to earners and non-earners. Journal of the Royal Statistical Society, 99 (1):147–161, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980316.

Hill:1936:EPM

[644] A. Bradford Hill. The enumeration of population at the meetings of the Society. Journal of the Royal Statistical Society, 99(1):162–164B, ???? 1936. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980317.

G:1936:BRBa

[645] M. G. Book review: Einfuhrung in die Mathematische Statistik, by Oskar N. Anderson; Julius Springer. Journal of the Royal Statistical Society, 99 (1):165–168, ???? 1936. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980318.

C:1936:BRBa

[646] W. G. C. Book review: La Theorie des Probabilities, by P. Van Deuren. Journal of the Royal Statistical Society, 99 (1):169–170, ???? 1936. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980319.

B:1936:BRBa

[647] H. J. B. Book review: Statistique et Applications, by Georges Darmois. Journal of the Royal Statistical Society, 99 (1):170–171, ???? 1936. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980320.

G:1936:BRBb

[648] R. F. G. Book review: Elementary Statistics for Indian Students (Volume 1), by D. S. Dubey; S. L. Agrawal. Journal of the Royal Statistical Society, 99(1):171–172, ???? 1936. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980321.

F:1936:BRBa

[649] A. W. F. Book review: Studies of Differential Fertility in Sweden, by Karl Arvid Edin; Edward P. Hutchinson; Folketellingen i Norge (Census of December 1st, 1930. Fertility of Marriages). Journal of the Royal Statistical Society, 99(1):172–174, ???? 1936.

ISSN 0952-8385. URL http://www.

jstor.org/stable/2980322.

G:1936:BRBc

[650] R. F. G. Book review: The Social Survey of Merseyside, by D. Caradog Jones. Journal of the Royal Statistical Society, 99(1):174–178, ???? 1936. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980323.

S:1936:BRBa

[651] F. S. Book review: A Social Survey of Plymouth, A Preliminary Report. Journal of the Royal Statistical Society, 99 (1):178–179, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980324.

M:1936:BRBa

[652] H. W. M. Book review: Britain in Depression: A Record of British Industries since 1929. Journal of the Royal Statistical Society, 99(1):179–181, ???? 1936. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980325.

L:1936:BRBa

[653] H. L. Book review: Controlling Depressions, by Paul H. Douglas. Journal of the Royal Statistical Society, 99 (1):181–182, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980326.

H:1936:BRBa

[654] R. G. H. Book review: The Trade Balance; A Problem in National Planning, by K. E. Edgeworth; The Price Level, by K. E. Edgeworth. Journal of the Royal Statistical Society, 99 (1):182–184, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980327.

H:1936:BRBb

[655] H. L. Book review: The Problems of the Foreign Exchange, by L. L. B. Angas. Journal of the Royal Statistical Society, 99(1):184–185, ???? 1936. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980328.

M:1936:BRBb

[656] H. W. M. Book review: The International Banks, by A. S. J. Baster. Journal of the Royal Statistical Society, 99 (1):185–186, ???? 1936. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980329.

M:1936:BRBc

[657] H. W. M. Book review: The Clash of Progress and Security, by Allan G. B. Fisher. Journal of the Royal Statistical Society, 99(1):186–188, ???? 1936. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980330.

T:1936:BRBa

[658] G. F. T. Book review: The Depreciation of Capital Analytically Considered, by R. F. Fowler. Journal of the Royal Statistical Society, 99(1):188–190, ???? 1936. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2980331.

M:1936:BRBd

[659] H. W. M. Book review: The Thirty Three Millions: A Study of the Building Society Movement and The Story of The Abbey Road Society, by Harold Bellman. Journal of the Royal Statistical Society, 99(1):190–191, ???? 1936. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980332.

L:1936:BRBb

[660] H. G. L. Book review: Railways and Roads in Pioneer Development Overseas, by J. E. Holmstrom. Journal of the Royal Statistical Society, 99(1): 191–192, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980333.

T:1936:BRBb

[661] R. J. T. Book review: The Agricultural Dilemma, by Viscount Astor; B. Seebohm Rowntree. Journal of the Royal Statistical Society, 99(1):192–193, ???? 1936. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980334.

T:1936:BRBc

[662] R. J. T. Book review: International Wheat Policy and Planning.– Wheat Studies, by Alanzo Taylor. Journal of the Royal Statistical Society, 99 (1):194–195, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980335.

M:1936:BRBe

[663] H. W. M. Book review: Considerations on the Present Evolution of Agricultural Protectionism. Journal of the Royal Statistical Society, 99(1): 195–196, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980336.

Anonymous:1936:ONPa

[664] Anonymous. Other new publications. Journal of the Royal Statistical Society, 99(1):196–198, ???? 1936. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980337.

Anonymous:1936:SNa

[665] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 99 (1):199–219, ???? 1936. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980338.

Anonymous:1936:CNa

[666] Anonymous. Current notes. Journal of the Royal Statistical Society, 99 (1):220–222, ???? 1936. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980339.

Anonymous:1936:SEAa

[667] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 99 (1):223–229, ???? 1936. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980340.

Anonymous:1936:LALa

[668] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 99(1):230–238, ???? 1936. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980341.

Anonymous:1936:PRa

[669] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 99 (1):239–246, ???? 1936. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980342.

Anonymous:1936:BM

[670] Anonymous. Back matter. Journal of the Royal Statistical Society, 99(1):??, ???? 1936. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980343.

Anonymous:1936:VIa

[671] Anonymous. Volume information. Journal of the Royal Statistical Society, 99(2):??, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980577.

Anonymous:1936:FMb

[672] Anonymous. Front matter. Journal of the Royal Statistical Society, 99(2):??, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980578.

Anonymous:1936:CRT

[673] Anonymous. Corrigenda: The recent trend in England and Wales of mortality from phthisis at Young adult ages. Journal of the Royal Statistical Society, 99(2):??, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980579. See [675].

Anonymous:1936:TMS

[674] Anonymous. Third meeting of the session 1935–36, 21st January, 1936, the President, Prof. M. Greenwood, F.R.S., in the Chair. Journal of the Royal Statistical Society, 99(2):247, ???? 1936. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980580.

Hill:1936:RTE

[675] A. Bradford Hill. The recent trend in England and Wales of mortality from phthisis at Young adult ages. Journal of the Royal Statistical Society, 99 (2):247–296, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/

stable/2980581. See corrigenda [673].

White:1936:SSA

[676] G. Ronald White. Some statistical aspects of future trading on a commodity exchange. Journal of the Royal Statistical Society, 99(2):297–342, ???? 1936. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980582.

Anonymous:1936:WP

[677] Anonymous. Wholesale prices in 1935. Journal of the Royal Statistical Society, 99(2):343–359, ???? 1936. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980583.

Anonymous:1936:MEC

[678] Anonymous. Memorial on earnings and cost of living. Journal of the Royal Statistical Society, 99(2):360–363, ???? 1936. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2980584.

Bowley:1936:NSL

[679] A. L. Bowley. The new survey of London life and labour effect of modifying the poverty line. Journal of the Royal Statistical Society, 99(2):364–366, ???? 1936. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2980585.

Rhodes:1936:PINb

[680] E. C. Rhodes. The precision of index Numbers–II. Journal of the Royal Statistical Society, 99(2):367–369, ???? 1936. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2980586.

White:1936:PES

[681] Terence White. A possible extension of statistical method. Journal of the Royal Statistical Society, 99(2):

370–371, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980587.

Kantorowitsch:1936:EJP

[682] Miron Kantorowitsch. Estimate of the Jewish population of London in 1929– 1933. Journal of the Royal Statistical Society, 99(2):372–379, ???? 1936. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980588.

S:1936:BRBb

[683] F. S. Book review: International Institute Examinations Enquiry: An Examination of Examinations, by Philip Hartog; E. C. Rhodes. Journal of the Royal Statistical Society, 99(2): 380–383, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980589.

F:1936:BRBb

[684] A. W. F. Book review: The General Theory of Employment, Interest and Money, by John Maynard Keynes. Journal of the Royal Statistical Society, 99(2):383–388, ???? 1936. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980590.

H:1936:BRBc

[685] R. G. H. Book review: 100 Per Cent. Money, by Irving Fisher. Journal of the Royal Statistical Society, 99 (2):388–390, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980591. See note [801].

G:1936:BRBd

[686] A. T. K. G. Book review: The Fundamentals of Money, by Henry Houston.

Journal of the Royal Statistical Society, 99(2):390–391, ???? 1936. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980592.

E:1936:BRBa

[687] W. A. E. Book review: The Problem of Credit Policy, by E. F. M. Durbin. Journal of the Royal Statistical Society, 99(2):391–393, ???? 1936. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980593.

H:1936:BRBd

[688] R. G. H. Book review: British International Gold Movements and Banking Policy, 1881–1913, by W. Edwards Beach. Journal of the Royal Statistical Society, 99(2):393–394, ???? 1936. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980594.

M:1936:BRBf

[689] H. W. M. Book review: A Hundred Years of Joint Stock Banking, by W. F. Crick; J. E. Wadsworth. Journal of the Royal Statistical Society, 99 (2):394–397, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980595.

M:1936:BRBg

[690] H. W. M. Book review: History of the London Discount Market, by W. T. C. King. Journal of the Royal Statistical Society, 99(2):397–399, ???? 1936. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980596.

C:1936:BRBb

[691] H. C. Book review: Wages and Labour in the Lancashire Cotton Spinning Industry, by J. Jewkes; E. M. Gray. Journal of the Royal Statistical Society, 99 (2):399–401, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980597.

C:1936:BRBc

[692] H. C. Book review: Management Planning and Control, by A. G. H. Dent. Journal of the Royal Statistical Society, 99(2):401–402, ???? 1936. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980598.

B:1936:BRBb

[693] W. A. B. Book review: Employment Research. An Introduction to the McGill Programme of Research in the Social Sciences, by Leonard C. Marsh. Journal of the Royal Statistical Society, 99(2):402–403, ???? 1936. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980599.

L:1936:BRBc

[694] H. G. L. Book review: Road and Rail in Forty Countries, by Paul Wohl; A. Albitreccia. Journal of the Royal Statistical Society, 99(2):403–404, ???? 1936. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980600.

Anonymous:1936:ONPb

[695] Anonymous. Other new publications. Journal of the Royal Statistical Society, 99(2):404–406, ???? 1936. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980601.

Anonymous:1936:SNb

[696] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 99 (2):407–419, ???? 1936. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980602.

Anonymous:1936:CNb

[697] Anonymous. Current notes. Journal of the Royal Statistical Society, 99 (2):420–423, ???? 1936. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980603.

Anonymous:1936:PKP

[698] Anonymous. Professor Karl Pearson, F.R.S., 1857–1936. Journal of the Royal Statistical Society, 99(2): 424–428, ???? 1936. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980604.

Anonymous:1936:SEAb

[699] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 99 (2):429–436, ???? 1936. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980605.

Anonymous:1936:LALb

[700] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 99(2):437–445, ???? 1936. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980606.

Anonymous:1936:PRb

[701] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 99 (2):446–455, ???? 1936. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980607.

Anonymous:1936:FMc

[702] Anonymous. Front matter. Journal of the Royal Statistical Society, 99 (3):i–viii, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980344.

Gibbon:1936:ERL

[703] Gwilym Gibbon. The expenditure and revenue of local authorities. Journal of the Royal Statistical Society, 99 (3):457–515, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980345.

Keynes:1936:WSJ

[704] J. M. Keynes. William Stanley Jevons 1835–1882: A centenary allocation on his life and work as economist and statistician. Journal of the Royal Statistical Society, 99(3):516–555, ???? 1936. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980346.

Anonymous:1936:RML

[705] Anonymous. Reply from the Minister of Labour to the memorial on earnings and cost of living addressed to him by the President and Council of the Royal Statistical Society. Journal of the Royal Statistical Society, 99(3):556–558, ???? 1936. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980347.

Nair:1936:NEL

[706] K. Raghavan Nair. A note on the existence of lagging correlations between two random series. Journal of the Royal Statistical Society, 99(3):559–566, ???? 1936. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980348.

I:1936:BRBa

[707] J. O. I. Book review: The Application of Statistical Methods to Industrial Standardization and Quality Control, by E. S. Pearson. Journal of the Royal Statistical Society, 99(3):567–569, ???? 1936. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2980349.

C:1936:BRBd

[708] W. G. C. Book review: Tables of the Higher Mathematical Functions. Vol. II, by H. T. Davis. Journal of the Royal Statistical Society, 99(3):569–570, ???? 1936. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2980350.

C:1936:BRBe

[709] E. G. C. Book review: A Point Scale of Performance Tests, by Grace Arthur. Journal of the Royal Statistical Society, 99(3):570–571, ???? 1936. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980351.

G:1936:BRBe

[710] R. F. G. Book review: A Course in Business Statistics, by Charles H. Bayliss. Journal of the Royal Statistical Society, 99(3):571–572, ???? 1936. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980352.

B:1936:BRBc

[711] W. A. B. Book review: Seasonal Variations in Employment, by Christopher Saunders; G. W. Daniels. Journal of the Royal Statistical Society, 99 (3):572–573, ???? 1936. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980353.

M:1936:BRBh

[712] H. W. M. Book review: The Downfall of the Gold Standard, by Gustav Cassel. Journal of the Royal Statistical Society, 99(3):574–575, ???? 1936. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980354.

E:1936:BRBb

[713] W. A. E. Book review: The Rational of Central Banking, by Vera C. Smith. Journal of the Royal Statistical Society, 99(3):575–576, ???? 1936. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980355.

E:1936:BRBc

[714] B. E. Book review: British Banks and the London Money Market, by R. J. Truptil; Robert Kindersley. Journal of the Royal Statistical Society, 99 (3):576–577, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980356.

M:1936:BRBi

[715] H. W. M. Book review: The Theory of Monetary Policy, by B. P. Adarkar. Journal of the Royal Statistical Society, 99(3):577–578, ???? 1936. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980357.

M:1936:BRBj

[716] H. W. M. Book review: The New Industrial System, by Hermann Levy. Journal of the Royal Statistical Society, 99(3):578–579, ???? 1936. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980358.

M:1936:BRBk

[717] H. W. M. Book review: The Framework of the Pricing System, by E. H. Phelps Brown. Journal of the Royal Statistical Society, 99(3):579–580, ???? 1936. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980359.

M:1936:BRBl

[718] H. W. M. Book review: Economics and Technology, by Ferdynand Zweig. Journal of the Royal Statistical Society, 99 (3):580–582, ???? 1936. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980360.

C:1936:BRBf

[719] A. C. C. Book review: The Consequences of Post-War Price Changes, by K. T. Shah. Journal of the Royal Statistical Society, 99(3):582–583, ???? 1936. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980361.

T:1936:BRBd

[720] R. J. T. Book review: The History of Milk Prices, by Ruth L. Cohen. Journal of the Royal Statistical Society, 99 (3):583–584, ???? 1936. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980362.

G:1936:BRBf

[721] R. F. G. Book review: Agriculture and The Trade Cycle, by John H. Kirk. Journal of the Royal Statistical Society, 99(3):584–585, ???? 1936. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980363.

T:1936:BRBe

[722] R. J. T. Book review: Agricultural Credit, being a Study of Recent Developments in Agricultural Credit Administration in the United States of America, by Anwar Iqbal Qureshi. Journal of the Royal Statistical Society, 99 (3):585–586, ???? 1936. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980364.

C:1936:BRBg

[723] A. C. C. Book review: Fluctuations in American Business 1790–1860, by Walter Buckingham Smith; Arthur Harrison Cole. Journal of the Royal Statistical Society, 99(3):586–588, ???? 1936. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980365.

M:1936:BRBm

[724] H. W. M. Book review: The Medical Dictator and Other Biographical Studies, by Major Greenwood. Journal of the Royal Statistical Society, 99 (3):588–589, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980366.

Anonymous:1936:ONPc

[725] Anonymous. Other new publications. Journal of the Royal Statistical Society, 99(3):589–592, ???? 1936. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980367.

Anonymous:1936:SNc

[726] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 99 (3):593–603, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980368.

Anonymous:1936:CNc

[727] Anonymous. Current notes. Journal of the Royal Statistical Society, 99 (3):604–609, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980369.

E:1936:SAW

[728] W. P. E. Sir Alfred William Watson, K.C.B., F.I.A., 1870–1936. Journal of the Royal Statistical Society, 99 (3):610–613, ???? 1936. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980370.

Anonymous:1936:SEAc

[729] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 99 (3):614–619, ???? 1936. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980371.

Anonymous:1936:LALc

[730] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 99(3):620–625, ???? 1936. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980372.

Anonymous:1936:VIb

[731] Anonymous. Volume information. Journal of the Royal Statistical Society, 99(4):863–viii, ???? 1936. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980087.

Anonymous:1936:FMd

[732] Anonymous. Front matter. Journal of the Royal Statistical Society, 99 (4):i–viii, ???? 1936. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980088.

Stamp:1936:IPL

[733] Josiah Stamp. The influence of the price level on the higher incomes. Journal of the Royal Statistical Society, 99 (4):627–673, ???? 1936. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980089.

Greenwood:1936:EDR

[734] Major Greenwood. English deathrates, past, present and future: A valedictory address. Journal of the Royal Statistical Society, 99(4):674–707, ???? 1936. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980090.

Anonymous:1936:PM

[735] Anonymous. Proceedings of the meeting. Journal of the Royal Statistical Society, 99(4):708–713, ???? 1936. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980091.

Irwin:1936:RAM

[736] J. O. Irwin. Recent advances in mathematical statistics (1934). Journal of the Royal Statistical Society, 99(4): 714–769, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980092.

Yule:1936:PBD

[737] G. Udny Yule. On a parallelism between differential coefficients and regression coefficients. Journal of the Royal Statistical Society, 99(4): 770–771, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980093.

Anonymous:1936:RCF

[738] Anonymous. Report of the Council for the financial year ended December 31, 1935, and for the sessional year ending June 16, 1936, presented at the one hundred and second Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts, John Street, Adelphi, W.C. 2, on June 16, 1936. Journal of the Royal Statistical Society, 99 (4):772–785, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980094.

Anonymous:1936:POH

[739] Anonymous. Proceedings of the one hundred and second Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts on Tuesday, June 16th, 1936. Journal of the Royal Statistical Society, 99(4):786, ???? 1936. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980095.

C:1936:BRBh

[740] W. G. C. Book review: Applications des Probabilites, by P. Van Deuren. Journal of the Royal Statistical Society, 99(4):787–788, ???? 1936. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980096.

Br:1936:BRB

[741] J. Br. Book review: Mathematical Treatment of the Results of Agricultural and other Experiments, by M. J. Van Uven. Journal of the Royal Statistical Society, 99(4):788–789, ???? 1936. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980097.

H:1936:BRBe

[742] R. G. H. Book review: Interest and Prices; A Study of the Causes Regulating the Value of Money, by Knut Wicksell. Journal of the Royal Statistical Society, 99(4):789–790, ???? 1936. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980098.

B:1936:BRBd

[743] A. L. B. Book review: Prices in the Trade Cycle, by Gerhard Tintner. Journal of the Royal Statistical Society, 99(4):790–793, ???? 1936. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980099.

E:1936:BRBd

[744] W. A. E. Book review: Monetary Problems of the British Empire, by S. E. Harris. Journal of the Royal Statistical Society, 99(4):793–794, ???? 1936. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980100.

C:1936:BRBi

[745] N. E. C. Book review: Paris as a Financial Centre, by Margaret G. Myers. Journal of the Royal Statistical Society, 99(4):794–796, ???? 1936. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980101.

G:1936:BRBg

[746] C. O. G. Book review: Soviet Money and Finance, by L. E. Hubbard. Journal of the Royal Statistical Society, 99 (4):796–797, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980102.

G:1936:BRBh

[747] C. O. G. Book review: Deficits and Depressions, by Dan Throop Smith. Journal of the Royal Statistical Society, 99(4):797, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980103.

G:1936:BRBi

[748] A. T. K. G. Book review: The Behaviour of Money, by James W. Angell. Journal of the Royal Statistical Society, 99(4):797–799, ???? 1936. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980104.

I:1936:BRBb

[749] J. O. I. Book review: The Measurement of Population Growth, by R. R. Kuczynski; Population Movements. by Robert R. Kuczynski. Journal of the Royal Statistical Society, 99 (4):799–802, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980105.

H:1936:BRBf

[750] A. B. H. Book review: The Struggle for Population, by D. V. Glass. Journal of the Royal Statistical Society, 99 (4):803–805, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980106.

L:1936:BRBd

[751] H. L. Book review: The British Immigrant: His Social and Economic Adjustment in Canada, by Lloyd G. Reynolds. Journal of the Royal Statistical Society, 99(4):805–806, ???? 1936. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980107.

C:1936:BRBj

[752] H. C. Book review: Public Works in Prosperity and Depression, by Arthur D. Gayer. Journal of the Royal Statistical Society, 99(4):806–807, ???? 1936. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980108.

C:1936:BRBk

[753] H. C. Book review: Cotton and the A.A.A, by Henry I. Richards. Journal of the Royal Statistical Society, 99 (4):808–809, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980109.

L:1936:BRBe

[754] H. G. L. Book review: Economic Principles of Transportation, by W. F. Jackman. Journal of the Royal Statistical Society, 99(4):809–812, ???? 1936. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980110.

Anonymous:1936:ONPd

[755] Anonymous. Other new publications. Journal of the Royal Statistical Society, 99(4):812–815, ???? 1936. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980111.

Anonymous:1936:SNd

[756] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 99 (4):816–829, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980112.

Anonymous:1936:CNd

[757] Anonymous. Current notes. Journal of the Royal Statistical Society, 99 (4):830–835, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980113.

G:1936:SWH

[758] M. G. Sir William Heaton Hamer, M.D., F.R.C.P. Journal of the Royal Statistical Society, 99(4):836–837, ???? 1936. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980114.

B:1936:HF

[759] J. B. H. S. Foxwell. Journal of the Royal Statistical Society, 99(4): 837–841, ???? 1936. ISSN 0952-

8385. URL http://www.jstor.org/

stable/2980115. See erratum [765].

Anonymous:1936:SEAd

[760] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 99 (4):842–844, ???? 1936. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980116.

Anonymous:1936:LALd

[761] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 99(4):845–848, ???? 1936. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980117.

Anonymous:1936:PRL

[762] Anonymous. Periodicals received by the Library. Journal of the Royal Statistical Society, 99(4):849–862, ???? 1936. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2980118.

Anonymous:1937:VIa

[763] Anonymous. Volume information. Journal of the Royal Statistical Society, 100(1):v–vii, ???? 1937. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980276.

Anonymous:1937:FM

[764] Anonymous. Front matter. Journal of the Royal Statistical Society, 100(1):i–iii, ???? 1937. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980277.

Anonymous:1937:EHF

[765] Anonymous. Erratum: H. S. Foxwell. Journal of the Royal Statistical Society, 100(1):??, ???? 1937. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980278. See [759].

oftheDene:1937:CS

[766] Lord Kennet of the Dene. The consumption of statistics. Journal of the Royal Statistical Society, 100 (1):1–11, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980279.

Anonymous:1937:PM

[767] Anonymous. Proceedings of the meeting. Journal of the Royal Statistical Society, 100(1):11–17, ???? 1937. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980280.

Rhodes:1937:CIB

[768] E. C. Rhodes. The construction of an index of business activity. Journal of the Royal Statistical Society, 100 (1):18–66, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980281.

Lewis-Faning:1937:CBE

[769] E. Lewis-Faning. Comparison between enumerated census populations and the population calculated from recorded births and deaths at ages under 5. Journal of the Royal Statistical Society, 100(1):67–68, ???? 1937. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980282.

Cochran:1937:EBS

[770] W. G. Cochran. The efficiencies of the binomial series tests of significance of a mean and of a correlation coefficient. Journal of the Royal Statistical Society, 100(1):69–73, ???? 1937. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980283.

George:1937:NCP

[771] R. F. George. A new calculation of the poverty line. Journal of the Royal Statistical Society, 100(1):74–95, ???? 1937. ISSN 0952-8385. URL http:/

                                        /www.jstor.org/stable/2980284.

Douglas:1937:IFS

[772] Iris Douglas. Irish Free State Census of Distribution, 1933. Journal of the Royal Statistical Society, 100 (1):96–99, ???? 1937. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980285.

Bowley:1937:TTS

[773] A. L. Bowley. The twenty-third session of the International Institute of Statistics, 1936. Journal of the Royal Statistical Society, 100(1):100–102, ???? 1937. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980286.

S:1937:BRBa

[774] P. S. Book review: Experimental Epidemiology, by Major Greenwood; A. Bradford Hill; W. W. C. Topley; J. Wilson. Journal of the Royal Statistical Society, 100(1):103–106, ???? 1937. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980287.

S:1937:BRBb

[775] F. S. Book review: The Marks of Examiners, by Philip Hartog; E. C. Rhodes. Journal of the Royal Statistical Society, 100(1):106–110, ???? 1937. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980288.

C:1937:BRBa

[776] W. G. C. Book review: Statistical Methods in Biology, Medicine and Psychology, by C. B. Davenport; Merle P. Ekas. Journal of the Royal Statistical Society, 100(1):110–111, ???? 1937. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980289.

C:1937:BRBb

[777] E. G. C. Book review: The Growth of the Surface Area of the Human Body, by Edith Boyd. Journal of the Royal Statistical Society, 100(1): 111–112, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980290.

T:1937:BRBa

[778] L. H. C. T. Book review: Trend Analysis of Statistics: Theory and Technique, by Max Sasuly. Journal of the Royal Statistical Society, 100(1):112, ???? 1937. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980291.

C:1937:BRBc

[779] W. G. C. Book review: Grundfragen der Angewandten Wahrscheinlichkeitsrechnung und Theoretischen Statistik, by K. Marbe. Journal of the Royal Statistical Society, 100(1):112–113, ???? 1937. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980292.

A:1937:BRB

[780] R. G. D. A. Book review: Les Methodes de Construction des Index-Numbers, by Jules Lejeune. Journal of the Royal Statistical Society, 100 (1):113–114, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980293.

C:1937:BRBd

[781] C. C. Book review: Cyclical Fluctuations in Commodity Stocks, by Ralph H. Blodgett. Journal of the Royal Statistical Society, 100(1):114–115, ???? 1937. ISSN 0952-8385. URL http:/

                                        /www.jstor.org/stable/2980294.

G:1937:BRBa

[782] R. F. G. Book review: The Home Market, by G. Harrison; F. C. Mitchell; Frank Pick. Journal of the Royal Statistical Society, 100(1):115–116, ???? 1937. ISSN 0952-8385. URL http:/

                                        /www.jstor.org/stable/2980295.

T:1937:BRBb

[783] R. J. T. Book review: Milk: Report of the Reorganization Commission for Great Britain. Journal of the Royal Statistical Society, 100(1): 116–118, ???? 1937. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980296.

V:1937:BRB

[784] J. W. V. Book review: A Study of Industrial Edinburgh and the Surrounding Area, 1923–1934. Volume I, by Nora Milnes. Journal of the Royal Statistical Society, 100(1):118–120, ???? 1937. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980297.

G:1937:BRBb

[785] C. O. G. Book review: Science of Public Finance, by G. Findlay Shirras. Journal of the Royal Statistical Society, 100(1):120–121, ???? 1937. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980298.

C:1937:BRBe

[786] A. C. C. Book review: Indian Currency and Reserve Bank Problems, by Benoy Kumar Sarkar. Journal of the Royal Statistical Society, 100(1): 121–122, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980299.

M:1937:BRBa

[787] H. W. M. Book review: An Essay on Economy and Value, by Alec L. Macfie. Journal of the Royal Statistical Society, 100(1):122–123, ???? 1937. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980300.

M:1937:BRBb

[788] H. W. M. Book review: Mr. Keynes and the Labour Movement, by A. L. Rowse. Journal of the Royal Statistical Society, 100(1):123–124, ???? 1937. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980301.

Anonymous:1937:ONPa

[789] Anonymous. Other new publications. Journal of the Royal Statistical Society, 100(1):124–126, ???? 1937. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980302.

Anonymous:1937:SNa

[790] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 100 (1):127–143, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980303.

Anonymous:1937:CNa

[791] Anonymous. Current notes. Journal of the Royal Statistical Society, 100 (1):144–148, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980304.

Anonymous:1937:HW

[792] Anonymous. Harald Westergaard, 1853–1937. Journal of the Royal Statistical Society, 100(1):149–150, ???? 1937. ISSN 0952-8385. URL http:/

                                        /www.jstor.org/stable/2980305.

Anonymous:1937:SEAa

[793] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 100 (1):151–161, ???? 1937. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980306.

Anonymous:1937:LALa

[794] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 100(1):162–172, ???? 1937. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980307.

Anonymous:1937:PRa

[795] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 100 (1):173–180, ???? 1937. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980308.

Anonymous:1937:BMa

[796] Anonymous. Back matter. Journal of the Royal Statistical Society, 100(1):??, ???? 1937. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980309.

Edge:1937:DBC

[797] P. Granville Edge. The demography of British colonial possessions: A note on the assembly and interpretation of data. Journal of the Royal Statistical Society, 100(2):181–231, ???? 1937.

ISSN 0952-8385. URL http://www.

jstor.org/stable/2984616.

Flux:1937:SSS

[798] Alfred W. Flux. Some Scandinavian statistics: Fragments from a note-book. Journal of the Royal Statistical Society, 100(2):232–273, ???? 1937. ISSN

0952-8385. URL http://www.jstor.
org/stable/2984617.

Anonymous:1937:WP

[799] Anonymous. Wholesale prices in 1936. Journal of the Royal Statistical Society, 100(2):274–290, ???? 1937. ISSN

0952-8385. URL http://www.jstor.
org/stable/2984618.

Thompson:1937:CVA

[800] R. J. Thompson. Changes in the value of the agricultural output of England and Wales. Journal of the Royal Statistical Society, 100(2):291–295, ???? 1937. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2984619.

Fisher:1937:NSR

[801] Irving Fisher. Note suggested by review of “100 per cent money”. Journal of the Royal Statistical Society, 100 (2):296–298, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2984620. See [685].

S:1937:BRBc

[802] F. S. Book review: Elements of Probability, by H. Levy; L. Roth. Journal of the Royal Statistical Society, 100 (2):299–300, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2984621.

G:1937:BRBc

[803] C. O. G. Book review: The Fundamental Principles of Taxation in the Light of Modern Developments, by Josiah Stamp. Journal of the Royal Statistical Society, 100(2):300–301, ???? 1937. ISSN 0952-8385. URL http://www.

jstor.org/stable/2984622.

F:1937:BRBa

[804] A. W. F. Book review: Monetary Policy and Crises, a Study of Swedish Experience, by Brinley Thomas. Journal of the Royal Statistical Society, 100 (2):301–302, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2984623.

M:1937:BRBc

[805] H. W. M. Book review: An Introduction to Economic Analysis and Policy, by J. E. Meade. Journal of the Royal Statistical Society, 100(2): 302–303, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2984624.

H:1937:BRBa

[806] R. G. H. Book review: The Economics of Stationary States, by A. C. Pigou. Journal of the Royal Statistical Society, 100(2):303–306, ???? 1937. ISSN

0952-8385. URL http://www.jstor.
org/stable/2984625.

M:1937:BRBd

[807] H. W. M. Book review: Abolish Slumps, a Diagnosis of the Trade Cycle, by A. W. Knight. Journal of the Royal Statistical Society, 100(2): 306–307, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2984626.

M:1937:BRBe

[808] H. W. M. Book review: The ABC of the Foreign Exchanges, a Practical Guide, by George Clare; Norman Crump. Journal of the Royal Statistical Society, 100(2):307, ???? 1937. ISSN

0952-8385. URL http://www.jstor.
org/stable/2984627.

S:1937:BRBd

[809] R. W. S. Book review: Investment for Appreciation, by L. L. B. Angas. Journal of the Royal Statistical Society, 100 (2):307–309, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2984628.

B:1937:BRBa

[810] A. L. B. Book review: Labour Conditions in Western Europe, 1820 to 1935, by Jürgen Kuczynski. Journal of the Royal Statistical Society, 100 (2):309–310, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2984629.

M:1937:BRBf

[811] H. W. M. Book review: The World Agricultural Situation in 1934– 35. Journal of the Royal Statistical Society, 100(2):310–312, ???? 1937. ISSN

0952-8385. URL http://www.jstor.
org/stable/2984630.

T:1937:BRBc

[812] R. J. T. Book review: The Marketing of Australian and New Zealand Primary Products, by W. Millar Smith; J. Coatmain. Journal of the Royal Statistical Society, 100(2):312–313, ???? 1937. ISSN 0952-8385. URL http://www.

jstor.org/stable/2984631.

B:1937:BRBb

[813] W. A. B. Book review: Mines, Machines and Men, by W. D. Stewart. Journal of the Royal Statistical Society, 100(2):313–314, ???? 1937. ISSN

0952-8385. URL http://www.jstor.
org/stable/2984632.

S:1937:BRBe

[814] R. W. S. Book review: Building Society Practice, by G. S. Riley. Journal of the Royal Statistical Society, 100 (2):314–315, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2984633.

G:1937:BRBd

[815] G. Book review: The Social Services: A Historical Survey, by W. Hardy Wickwar. Journal of the Royal Statistical Society, 100(2):315, ???? 1937. ISSN 0952-8385. URL http://www.

jstor.org/stable/2984634.

B:1937:BRBc

[816] J. B. Book review: Letters to Thomas Robert Malthus on Political Economy and Stagnation of Commerece, by Jean Baptiste Say; Harold J. Laski. Journal of the Royal Statistical Society, 100 (2):315–317, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2984635.

B:1937:BRBd

[817] J. B. Book review: Paradoxes Inédits du Seigneur de Malestroit touchant les Monnoyes avec la Response du Président de la Tourette, by Luigi Einaudi. Journal of the Royal Statistical Society, 100(2):317–319, ???? 1937. ISSN 0952-8385. URL http://www.

jstor.org/stable/2984636.

Anonymous:1937:ONPb

[818] Anonymous. Other new publications. Journal of the Royal Statistical Society, 100(2):319–321, ???? 1937. ISSN

0952-8385. URL http://www.jstor.
org/stable/2984637.

Anonymous:1937:SNb

[819] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 100 (2):322–337, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2984638.

Anonymous:1937:CNb

[820] Anonymous. Current note. Journal of the Royal Statistical Society, 100(2):338, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2984639.

Anonymous:1937:SEAb

[821] Anonymous. Statistical and economic articles in recent periodical. Journal of the Royal Statistical Society, 100 (2):339–345, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2984640.

Anonymous:1937:LALb

[822] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 100(2):346–351, ???? 1937. ISSN 0952-8385. URL http://www.

jstor.org/stable/2984641.

Anonymous:1937:PRb

[823] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 100 (2):352–361, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2984642.

Anonymous:1937:BMb

[824] Anonymous. Back matter. Journal of the Royal Statistical Society, 100(2):??, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2984643.

Meek:1937:SME

[825] D. B. Meek. Some measures of economic activity in India. Journal of the Royal Statistical Society, 100(3): 363–395, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980524.

Greenwood:1937:VRC

[826] M. Greenwood. On the value of Royal Commissions in Sociological Research, with special reference to the birth-rate. Journal of the Royal Statistical Society, 100(3):396–414, ???? 1937. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980525.

Irwin:1937:FDD

[827] J. O. Irwin. The frequency distribution of the difference between two independent variates following the same Poisson distribution. Journal of the Royal Statistical Society, 100 (3):415–416, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980526.

Wisniewski:1937:NIP

[828] Jan Wisniewski. A note on inverse probability. Journal of the Royal Statistical Society, 100(3):417–420, ???? 1937. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980527.

Bresciani-Turroni:1937:PL

[829] C. Bresciani-Turroni. On Pareto’s law. Journal of the Royal Statistical Society, 100(3):421–432, ???? 1937. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980528.

W:1937:BRBa

[830] H. W. M. Book review: Report of the Import Duties Advisory Committee on the Present Position and Future Development of the Iron and Steel Industry. Journal of the Royal Statistical Society, 100(3):433–435, ???? 1937. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980529.

I:1937:BRBa

[831] L. I. Book review: Maritime Statistical Handbook. International Labour Office (League of Nations), Geneva. Journal of the Royal Statistical Society, 100 (3):436–437, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980530.

M:1937:BRBg

[832] H. W. M. Book review: The Trade Cycle, by R. F. Harrod; Capital and Employment. by R. G. Hawtrey. Journal of the Royal Statistical Society, 100 (3):438–440, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980531.

H:1937:BRBb

[833] R. G. H. Book review: The Theory of Forward Exchange, by Paul Einzig. Journal of the Royal Statistical Society, 100(3):440–443, ???? 1937. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980532.

M:1937:BRBh

[834] H. W. M. Book review: Industrial Reconstruction and the Control of Competition: The British Experiments, by Arthur Fletcher Lucas. Journal of the Royal Statistical Society, 100(3): 443–445, ???? 1937. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980533.

G:1937:BRBe

[835] C. O. G. Book review: Bank of England Operations, 1890–1914, by R. S. Sayers. Journal of the Royal Statistical Society, 100(3):445, ???? 1937. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980534.

S:1937:BRBf

[836] R. W. S. Book review: The Problem of Investment, by F. I. Shaffner. Journal of the Royal Statistical Society, 100 (3):445–447, ???? 1937. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980535.

G:1937:BRBf

[837] R. F. G. Book review: Trade Revival in a Depressed Area, by D. Caradog Jones. Journal of the Royal Statistical Society, 100(3):447–448, ???? 1937. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980536.

G:1937:BRBg

[838] R. F. G. Book review: Business Statistics and Statistical Method, by H. J. Weldon; Statistics for Professional Students, by R. L. A. Holmes; A Concise Manual of Statistics, by Clement Burton. Journal of the Royal Statistical Society, 100(3):448–450, ???? 1937. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980537.

G:1937:BRBh

[839] H. G. L. Book review: The Railway Worker, by G. M. Rountree. Journal of the Royal Statistical Society, 100 (3):450–451, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980538.

B:1937:BRBe

[840] C. H. L. B. Book review: Pension and Widows’ and Orphans’ Funds, by D. A. Porteous. Journal of the Royal Statistical Society, 100(3):451–453, ???? 1937. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980539.

T:1937:BRBd

[841] C. T. Book review: Musikstatistik, by Leo Wilzin. Journal of the Royal Statistical Society, 100(3):453–454, ???? 1937. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980540.

K:1937:BRB

[842] R. R. K. Book review: Umriss einer Geschichte der Preise und Lohne in Deutschland vom Ausgehenden Mittelalter bis zum Beginn des Neunzehnten Jahrhunderts, by M. J. Elsas. Journal of the Royal Statistical Society, 100(3):455, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980541.

T:1937:BRBe

[843] C. T. Book review: L’Enseignement economique en France et a l’etranger, by Charles Rist; Gaetan Pirou. Journal of the Royal Statistical Society, 100 (3):455–457, ???? 1937. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980542.

Anonymous:1937:ONPc

[844] Anonymous. Other new publications. Journal of the Royal Statistical Society, 100(3):457–460, ???? 1937. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980543.

Anonymous:1937:SNc

[845] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 100 (3):461–475, ???? 1937. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980544.

Anonymous:1937:CNc

[846] Anonymous. Current notes. Journal of the Royal Statistical Society, 100 (3):476–480, ???? 1937. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980545.

Anonymous:1937:SEAc

[847] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 100 (3):481–486, ???? 1937. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980546.

Anonymous:1937:LALc

[848] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 100(3):487–493, ???? 1937. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980547.

Anonymous:1937:VIb

[849] Anonymous. Volume information. Journal of the Royal Statistical Society, 100(4):721–725, ???? 1937. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980388.

Gibbon:1937:PSS

[850] Gwilym Gibbon. The public social services. Journal of the Royal Statistical Society, 100(4):495–557, ???? 1937. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980389.

Leak:1937:SRI

[851] H. Leak. Some results of the Import Duties Act. Journal of the Royal Statistical Society, 100(4):558–606, ???? 1937. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980390.

Thompson:1937:ALB

[852] R. J. Thompson. The agricultural labour bill in England and Wales. Journal of the Royal Statistical Society, 100 (4):607–624, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980391.

Broster:1937:SMD

[853] E. J. Broster. A simple method of deriving demand curves. Journal of the Royal Statistical Society, 100(4): 625–641, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980392.

Anonymous:1937:RCF

[854] Anonymous. Report of the Council for the financial year ended December 31st, 1936, and for the sessional year ending June 15th, 1937, presented at the one hundred and third Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts, John Street, Adelphi, W.C. 2, on June 15th, 1937. Journal of the Royal Statistical Society, 100 (4):642–655, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980393.

Anonymous:1937:POH

[855] Anonymous. Proceedings of the one hundred and third Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts on Tuesday, June 15, 1937. Journal of the Royal Statistical Society, 100(4):656, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980394.

I:1937:BRBb

[856] L. I. Book review: Principles of Medical Statistics, by A. Bradford Hill. Journal of the Royal Statistical Society, 100(4):657–658, ???? 1937. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980395.

S:1937:BRBg

[857] R. F. S. Book review: An Outline of Statistics, by Samuel Hays. Journal of the Royal Statistical Society, 100 (4):658–659, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980396.

H:1937:BRBc

[858] A. B. H. Book review: World Population; Past Growth and Present Trends, by A. M. Carr-Saunders. Journal of the Royal Statistical Society, 100(4): 659–660, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980397.

J:1937:BRBa

[859] J. H. J. Book review: Migration and Economic Opportunity, by Carter Goodrich. Journal of the Royal Statistical Society, 100(4):660–663, ???? 1937. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980398.

J:1937:BRBb

[860] J. H. J. Book review: Regional Shifts in the Bituminous Coal Industry, with special Reference to Pennsylvania, by Wilbert G. Fritz; Theodore A. Veenstra. Journal of the Royal Statistical Society, 100(4):663–665, ???? 1937. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980399.

S:1937:BRBh

[861] E. C. S. Book review: National Income and Outlay, by Colin Clark. Journal of the Royal Statistical Society, 100 (4):665–668, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980400.

S:1937:BRBi

[862] J. C. S. Book review: The Distribution of National Capital, by G. W. Daniels; H. Campion. Journal of the Royal Statistical Society, 100(4):668–669, ???? 1937. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980401.

F:1937:BRBb

[863] A. W. F. Book review: Essays in the Theory of Employment, by Joan Robinson. Journal of the Royal Statistical Society, 100(4):669–671, ???? 1937. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980402.

S:1937:BRBj

[864] R. W. S. Book review: Properity and Depression, by Gottfried von Haberler. Journal of the Royal Statistical Society, 100(4):671–673, ???? 1937. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980403.

W:1937:BRBb

[865] G. R. W. Book review: Some International Aspects of the Business Cycle, by Hans Neisser. Journal of the Royal Statistical Society, 100(4):673–674, ???? 1937. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980404.

D:1937:BRB

[866] I. D. Book review: Trial Census of Distribution in Six Towns, by British National Committee of the International Chamber of Commerce. Journal of the Royal Statistical Society, 100 (4):674–675, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980405.

M:1937:BRBi

[867] H. W. M. Book review: Economic Planning and International Order, by Lionel Robbins; Planning under Capitalism. The Problem of Planning in Great Britain. by A. W. Rather; The Problems and Practice of Economic Planning. by Raymond Burrows. Journal of the Royal Statistical Society, 100 (4):675–678, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980406.

M:1937:BRBj

[868] H. W. M. Book review: Predecessors of Adam Smith: The Growth of British Economic Thought, by E. A. J. Johnson. Journal of the Royal Statistical Society, 100(4):678–680, ???? 1937. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980407.

Anonymous:1937:ONPd

[869] Anonymous. Other new publications. Journal of the Royal Statistical Society, 100(4):680–682, ???? 1937. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980408.

Anonymous:1937:SNd

[870] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 100 (4):683–695, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980409.

Anonymous:1937:SEAd

[871] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 100 (4):696–701, ???? 1937. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980410.

Anonymous:1937:LALd

[872] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 100(4):702–706, ???? 1937. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980411.

Anonymous:1937:PRL

[873] Anonymous. Periodicals received by the Library. Journal of the Royal Statistical Society, 100(4):707–720, ???? 1937. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980412.

Anonymous:1938:FMa

[874] Anonymous. Front matter. Journal of the Royal Statistical Society, 101 (1):i–viii, ???? 1938. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980652.

Jones:1938:PIC

[875] R. Huws Jones. Physical indices and clinical assessments of the nutrition of schoolchildren. Journal of the Royal Statistical Society, 101(1):1–52, ???? 1938. ISSN 0952-8385. URL http:/

                                        /www.jstor.org/stable/2980653.

Isserlis:1938:TSC

[876] L. Isserlis. Tramp shipping cargoes, and freights. Journal of the Royal Statistical Society, 101(1):53–146, ???? 1938. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2980654.

Kendall:1938:RRS

[877] M. G. Kendall and B. Babington Smith. Randomness and random sampling numbers. Journal of the Royal Statistical Society, 101(1):147–

166, ????  1938.  ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980655.

Yule:1938:TTR

[878] G. Udny Yule. A test of Tippett’s random sampling numbers. Journal of the Royal Statistical Society, 101 (1):167–172, ???? 1938. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980656.

Eder:1938:EGP

[879] George Jackson Eder. Effect of gold price changes upon prices for other commodities. Journal of the Royal Statistical Society, 101(1):173–187, ???? 1938. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980657.

Shields:1938:ICD

[880] B. F. Shields. The Irish census of distribution. Journal of the Royal Statistical Society, 101(1):188–201, ???? 1938. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980658.

Ramsbottom:1938:WRU

[881] E. C. Ramsbottom. Wage rates in the United Kingdom, 1934–1937. Journal of the Royal Statistical Society, 101 (1):202–204, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980659.

S:1938:BRBa

[882] F. S. Book review: The Methods of Statistics: an Introduction Mainly for Experimentalists, by L. H. C. Tippett. Journal of the Royal Statistical Society, 101(1):205–206, ???? 1938. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980660.

B:1938:BRBa

[883] M. S. B. Book review: Random Variables and Probability Distributions, by H. Cramér. Journal of the Royal Statistical Society, 101(1):206–208, ???? 1938. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980661.

B:1938:BRBb

[884] M. S. B. Book review: L’Emploi des Observations Statistiques: Methodes d’Estimation, by Georges Darmois. Journal of the Royal Statistical Society, 101(1):208–209, ???? 1938.

ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980662.

K:1938:BRBa

[885] M. G. K. Book review: Interpolation and Allied Tables, by Nautical Almannac. Journal of the Royal Statistical Society, 101(1):209–210, ???? 1938. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980663.

K:1938:BRBb

[886] M. G. K. Book review: Sur la Theorie Mathematique des Jeux de Hasard et de Reflexion, by Rene de Possel. Journal of the Royal Statistical Society, 101 (1):210–211, ???? 1938. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980664.

R:1938:BRB

[887] E. C. R. Book review: The Mathematical Problem of the Price Index, by J. K. Montgomery. Journal of the Royal Statistical Society, 101(1):211, ???? 1938. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980665.

M:1938:BRBa

[888] H. W. M. Book review: A Study of the Capital Market in Post-War Britain, by A. T. K. Grant. Journal of the Royal Statistical Society, 101 (1):211–213, ???? 1938. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980666.

G:1938:BRBa

[889] C. O. G. Book review: British Corporation Finance 1775–1850: A Study of Preference Shares, by George Heberton Evans. Journal of the Royal Statistical Society, 101(1):213–214, ???? 1938.

ISSN 0952-8385. URL http://www.

jstor.org/stable/2980667.

L:1938:BRBa

[890] H. L. Book review: World Prices and the Building Industry, by G. F. Warren; F. A. Pearson. Journal of the Royal Statistical Society, 101(1): 214–215, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980668.

G:1938:BRBb

[891] R. F. G. Book review: Retail Distribution, a Critical Analysis, by Henry Smith. Journal of the Royal Statistical Society, 101(1):215–216, ???? 1938. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980669.

S:1938:BRBb

[892] R. W. S. Book review: Unemployment in the Learned Professions, by Walter M. Kotschnig. Journal of the Royal Statistical Society, 101(1): 217–218, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980670.

M:1938:BRBb

[893] H. W. M. Book review: Economic Planning in Australia, 1929–1936, by William Rupert MacLaurin. Journal of the Royal Statistical Society, 101 (1):218–219, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980671.

G:1938:BRBc

[894] A. T. K. G. Book review: Industrial Assurance: An Historical and Critical Study, by Arnold Wilson; Hermann Levy. Journal of the Royal Statistical Society, 101(1):219–221, ???? 1938. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980672.

B:1938:BRBc

[895] W. A. B. Book review: Report on the British Social Services. Journal of the Royal Statistical Society, 101 (1):222–223, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980673.

Sp:1938:BRB

[896] J. C. Sp. Book review: Problems of Town and Country Planning, by Gwilym Gibbon. Journal of the Royal Statistical Society, 101 (1):223–224, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980674.

T:1938:BRBa

[897] R. J. T. Book review: The World Agricultural Situation in 1935–36. Journal of the Royal Statistical Society, 101 (1):224–225, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980675.

I:1938:BRB

[898] L. I. Book review: An Introduction to the Theory of Statistics, by G. Udny Yule; M. G. Kendall. Journal of the Royal Statistical Society, 101 (1):225–226, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980676.

Anonymous:1938:ONPa

[899] Anonymous. Other new publications. Journal of the Royal Statistical Society, 101(1):226–227, ???? 1938. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980677.

Anonymous:1938:SNa

[900] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 101 (1):228–244, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980678.

Anonymous:1938:CNa

[901] Anonymous. Current notes. Journal of the Royal Statistical Society, 101 (1):245–247, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980679.

H:1938:WSG

[902] H. H. and J. W. William Sealy Gosset, 1876–1937. Journal of the Royal Statistical Society, 101(1):248–251, ???? 1938. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980680.

C:1938:PGW

[903] H. C. Professor George William Daniels. Journal of the Royal Statistical Society, 101(1):251–252, ???? 1938. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980681.

Anonymous:1938:SEAa

[904] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 101 (1):253–258, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980682.

Anonymous:1938:LALa

[905] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 101(1):259–266, ???? 1938.

ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980683.

Anonymous:1938:PRa

[906] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 101 (1):267–274, ???? 1938. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980684.

Anonymous:1938:BMa

[907] Anonymous. Back matter. Journal of the Royal Statistical Society, 101(1):??, ???? 1938. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980685.

Anonymous:1938:FMb

[908] Anonymous. Front matter. Journal of the Royal Statistical Society, 101 (2):i–viii, ???? 1938. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980208.

Houghton:1938:NIN

[909] C. T. Houghton. A new index number of agricultural prices. Journal of the Royal Statistical Society, 101(2): 275–316, ???? 1938. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980209.

Rae:1938:SRI

[910] George Rae. The statistics of the rubber industry. Journal of the Royal Statistical Society, 101(2):317–375, ???? 1938. ISSN 0952-8385. URL http:/

                                        /www.jstor.org/stable/2980210.

Anonymous:1938:WP

[911] Anonymous. Wholesale prices in 1937. Journal of the Royal Statistical Society, 101(2):376–393, ???? 1938. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980211.

Irwin:1938:RAM

[912] J. O. Irwin. Recent advances in mathematical statistics: Bibliography of mathematical statistics (1935, 1936, and first half of 1937). Journal of the Royal Statistical Society, 101(2): 394–433, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980212.

Cochran:1938:RAM

[913] W. G. Cochran. Recent advances in mathematical statistics: Recent work on the analysis of variance. Journal of the Royal Statistical Society, 101 (2):434–449, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980213.

Friedman:1938:MBD

[914] Milton Friedman. Mr. Broster on demand curves. Journal of the Royal Statistical Society, 101(2):450–454, ???? 1938. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980214.

Y:1938:BRB

[915] G. U. Y. Book review: Frequency Curves and Correlation, by W. Palin Elderton. Journal of the Royal Statistical Society, 101(2):455–456, ???? 1938. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980215.

K:1938:BRBc

[916] M. G. K. Book review: The Theory of Statistical Inference, by S. S. Wilks. Journal of the Royal Statistical Society, 101(2):457–458, ???? 1938. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980216.

K:1938:BRBd

[917] M. G. K. Book review: Elements of Statistics, by A. L. Bowley. Journal of the Royal Statistical Society, 101 (2):458–459, ???? 1938. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980217.

M:1938:BRBc

[918] H. W. M. Book review: Seventh Census of Canada, 1931. Volume I, Summary. Journal of the Royal Statistical Society, 101(2):459–461, ???? 1938. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980218.

M:1938:BRBd

[919] H. W. M. Book review: Wages and Income in the United Kingdom since 1860, by A. L. Bowley. Journal of the Royal Statistical Society, 101(2): 461–463, ???? 1938. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980219.

G:1938:BRBd

[920] R. F. G. Book review: A Survey of the Social Structure of England and Wales as Illustrated, by Statistics. by A. M. Carr-Saunders; D. Caradog Jones. Journal of the Royal Statistical Society, 101(2):463, ???? 1938. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980220.

G:1938:BRBe

[921] C. O. G. Book review: Socialism versus Capitalism, by A. C. Pigou. Journal of the Royal Statistical Society, 101(2):464, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980221.

S:1938:BRBc

[922] R. W. S. Book review: The Twilight of American Capitalism. An Economic Interpretation of the New Deal, by A. S. J. Baster. Journal of the Royal Statistical Society, 101(2):464–466, ???? 1938. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980222.

S:1938:BRBd

[923] R. W. S. Book review: N. R. A. Economic Planning, by Charles Frederick Roos. Journal of the Royal Statistical Society, 101(2):467–469, ???? 1938. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980223.

W:1938:BRB

[924] G. R. W. Book review: A Simple Theory of Capital, Wages and Profit or Loss, by E. C. van Dorp. Journal of the Royal Statistical Society, 101 (2):469–470, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980224.

M:1938:BRBe

[925] H. W. M. Book review: Monetary Policy and Economic Stabilisation, by Arthur D. Gayer. Journal of the Royal Statistical Society, 101 (2):470–471, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980225.

M:1938:BRBf

[926] H. W. M. Book review: The National Capital and Other Statistical Studies, by Josiah Stamp. Journal of the Royal Statistical Society, 101 (2):471–472, ???? 1938. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980226.

G:1938:BRBf

[927] C. O. G. Book review: The Balance Sheets of Imperialism, by Grover Clark. Journal of the Royal Statistical Society, 101(2):472–473, ???? 1938. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980227.

E:1938:BRBa

[928] B. E. Book review: Unit Trusts and How they Work, by C. L. Rosenheim; C. O. Merriman; W. Piercy. Journal of the Royal Statistical Society, 101 (2):473–474, ???? 1938. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980228.

Anonymous:1938:ONPb

[929] Anonymous. Other new publications. Journal of the Royal Statistical Society, 101(2):474–476, ???? 1938. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980229.

Anonymous:1938:SNb

[930] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 101 (2):477–484, ???? 1938. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980230.

Anonymous:1938:CNb

[931] Anonymous. Current notes. Journal of the Royal Statistical Society, 101 (2):485–489, ???? 1938. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980231.

Anonymous:1938:SEAb

[932] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 101 (2):490–497, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980232.

Anonymous:1938:LALb

[933] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 101(2):498–502, ???? 1938. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980233.

Anonymous:1938:PRb

[934] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 101 (2):503–510, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980234.

Anonymous:1938:BMb

[935] Anonymous. Back matter. Journal of the Royal Statistical Society, 101(2):??, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980235.

Anonymous:1938:FMc

[936] Anonymous. Front matter. Journal of the Royal Statistical Society, 101 (3):i–viii, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980627.

Glenday:1938:LPE

[937] Roy Glenday. Long-period economic trends. Journal of the Royal Statistical Society, 101(3):511–564, ???? 1938. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980628.

Snow:1938:TCM

[938] E. C. Snow. Is the trade cycle a myth? Journal of the Royal Statistical Society, 101(3):565–591, ???? 1938. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980629.

Kendall:1938:CUW

[939] M. G. Kendall. The conditions under which Sheppard’s corrections are valid. Journal of the Royal Statistical Society, 101(3):592–605, ???? 1938. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980630.

Percival:1938:NCP

[940] D. A. Percival. Notes on the count of a pagan tribe in West Africa. Journal of the Royal Statistical Society, 101 (3):606–616, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980631.

N:1938:BRBa

[941] J. N. Book review: Manuale di Statistica. Introduzione allo Studio Quantitativo dei Fatti Sociali, by Felice Vinci. Journal of the Royal Statistical Society, 101(3):617–619, ???? 1938. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980632.

N:1938:BRBb

[942] J. N. Book review: Wie Statistische Zahlen Entstehen. Die Entscheidenden Methodischen Vorgange, by Franz Zizek. Journal of the Royal Statistical Society, 101(3):619, ???? 1938. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980633.

S:1938:BRBe

[943] F. S. Book review: Mathematical Analysis for Economists, by R. G. D. Allen. Journal of the Royal Statistical Society, 101(3):619–620, ???? 1938. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980634.

S:1938:BRBf

[944] R. W. S. Book review: Some Theoretical Problems Suggested by the Movements of Interest Rates, Bond Yields and Stock Prices in the United States Since 1856, by Frederick R. Macaulay. Journal of the Royal Statistical Society, 101(3):620–623, ???? 1938. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980635.

S:1938:BRBg

[945] R. W. S. Book review: National Income and Capital Formation, 1919– 35. A Preliminary Report, by Simon Kuznets. Journal of the Royal Statistical Society, 101(3):623–624, ???? 1938. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980636.

F:1938:BRBa

[946] A. W. F. Book review: National Income of Sweden, 1861–1930, by Erik Lindahl; Einar Dahlgren; Karin Koch. Journal of the Royal Statistical Society, 101(3):625–627, ???? 1938. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980637.

M:1938:BRBg

[947] H. W. M. Book review: The Economic Policy of France, by George Peel. Journal of the Royal Statistical Society, 101 (3):627–628, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980638.

E:1938:BRBb

[948] B. E. Book review: Practice and Finance of Foreign Trade, by W. W. Syrett. Journal of the Royal Statistical Society, 101(3):628–629, ???? 1938. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980639.

L:1938:BRBb

[949] H. G. L. Book review: The Future of British Railways, by W. J. Stevens; Josiah Stamp. Journal of the Royal Statistical Society, 101(3): 629–630, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980640.

G:1938:BRBg

[950] R. F. G. Book review: Living on a Moderate Income, by E. H. Huntington; M. G. Luck. Journal of the Royal Statistical Society, 101(3):630–631, ???? 1938. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980641.

G:1938:BRBh

[951] D. V. G. Book review: Das Bevolkerungsproblem und seine Auswirkung in der neuen Deutschen Steuerreform, by Karl Barth. Journal of the Royal Statistical Society, 101(3):631–

632, ????  1938.  ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980642.

Anonymous:1938:ONPc

[952] Anonymous. Other new publications. Journal of the Royal Statistical Society, 101(3):632–634, ???? 1938. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980643.

Anonymous:1938:SNc

[953] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 101 (3):635–646, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980644.

Anonymous:1938:CNc

[954] Anonymous. Current notes. Journal of the Royal Statistical Society, 101 (3):647–651, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980645.

Anonymous:1938:AD

[955] Anonymous. Albert Delatour, 1858– 1938. Journal of the Royal Statistical Society, 101(3):652, ???? 1938. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980646.

Anonymous:1938:PG

[956] Anonymous. Paul I. Georgievsky, 1857–1938. Journal of the Royal Statistical Society, 101(3):652–653, ???? 1938. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980647.

Anonymous:1938:GJS

[957] Anonymous. George James Short Broomhall, 1857–1938. Journal of the Royal Statistical Society, 101(3): 653–654, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980648.

Anonymous:1938:SEAc

[958] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 101 (3):655–660, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980649.

Anonymous:1938:LALc

[959] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 101(3):661–668, ???? 1938. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980650.

Anonymous:1938:BMc

[960] Anonymous. Back matter. Journal of the Royal Statistical Society, 101(3):??, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980651.

Anonymous:1938:VI

[961] Anonymous. Volume information. Journal of the Royal Statistical Society, 101(4):821–vii, ???? 1938. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980495.

Anonymous:1938:FMd

[962] Anonymous. Front matter. Journal of the Royal Statistical Society, 101 (4):i–viii, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980496.

Stocks:1938:EOA

[963] Percy Stocks. The effects of occupation and of its accompanying environment on mortality. Journal of the Royal Statistical Society, 101(4):669–708, ???? 1938. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980497.

Kennet:1938:SFH

[964] Lord Kennet. Some footnotes to housing figures. Journal of the Royal Statistical Society, 101(4):709–725, ????

1938. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980498.

Anonymous:1938:PM

[965] Anonymous. Proceedings of the meeting. Journal of the Royal Statistical Society, 101(4):726–729, ???? 1938. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980499.

Nixon:1938:ICS

[966] J. W. Nixon. An international convention on statistics of wages and hours of work. Journal of the Royal Statistical Society, 101(4):730–735, ???? 1938. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980500.

Thompson:1938:SEA

[967] R. J. Thompson. State expenditure on agriculture in Great Britain in 1938– 39. Journal of the Royal Statistical Society, 101(4):736–737, ???? 1938. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980501.

Anonymous:1938:RCF

[968] Anonymous. Report of the Council for the financial year ended December 31st, 1937, and for the sessional year ending June 21st, 1938, presented at the one hundred and fourth Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts, John Street, Adelphi, W.C. 2, on June 21st, 1938. Journal of the Royal Statistical Society, 101 (4):738–751, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980502.

Anonymous:1938:POH

[969] Anonymous. Proceedings of the one hundred and fourth Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts on Tuesday, June 21st, 1938. Journal of the Royal Statistical Society, 101(4):752, ???? 1938. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980503.

K:1938:BRBe

[970] M. G. K. Book review: A First Course in Statistics, by E. F. Lindquist; A Study Manual for A First Course in Statistics. by E. F. Lindquist. Journal of the Royal Statistical Society, 101 (4):753–755, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980504.

N:1938:BRBc

[971] J. N. Book review: Die Statistischen Forschungsmethoden, by Emanuel Czuber; F. Burkhardt. Journal of the Royal Statistical Society, 101(4): 755–757, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980505.

F:1938:BRBb

[972] A. W. F. Book review: An Analysis of Price Behaviour during the Period 1855–1913, by Jorgen Pedersen; O. Strange Petersen. Journal of the Royal Statistical Society, 101 (4):757–758, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980506.

C:1938:BRB

[973] W. G. C. Book review: Lectures and Conferences on Mathematical Statistics, by J. Neyman. Journal of the Royal Statistical Society, 101(4): 758–759, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980507.

K:1938:BRBf

[974] M. G. K. Book review: The Kelley Statistical Tables, by Truman Lee Kelley. Journal of the Royal Statistical Society, 101(4):760–761, ???? 1938. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980508.

K:1938:BRBg

[975] M. G. K. Book review: Tables of the Correlation Coefficient, by F. N. David. Journal of the Royal Statistical Society, 101(4):761–762, ???? 1938. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980509.

D:1938:BRB

[976] J. A. D. Book review: Men Without Work: A Report Made to the Pilgrim Trust, by Archbishop of York; Lord Macmillan. Journal of the Royal Statistical Society, 101(4):762–765, ???? 1938. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980510.

F:1938:BRBc

[977] A. W. F. Book review: Kapitalbildung, by J. Marschak; W. Lederer. Journal of the Royal Statistical Society, 101 (4):765–766, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980511.

S:1938:BRBh

[978] R. W. S. Book review: Les Variations du Mouvement Saisonnier dans l’Industrie de la Construction, by Horst Mendershausen. Journal of the Royal Statistical Society, 101(4): 766–767, ???? 1938. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980512.

S:1938:BRBi

[979] R. W. S. Book review: Common-Stock Indexes 1871–1937, by Alfred Cowles. Journal of the Royal Statistical Society, 101(4):767–769, ???? 1938. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980513.

S:1938:BRBj

[980] R. W. S. Book review: Introduction to the Theory of Employment, by Joan Robinson. Journal of the Royal Statistical Society, 101(4):769–770, ???? 1938. ISSN 0952-8385. URL http:/

                                        /www.jstor.org/stable/2980514.

G:1938:BRBi

[981] D. V. G. Book review: La Population de la France, son Evolution et ses Perspectives, by M. Huber; H. Bunle; F. Boverat. Journal of the Royal Statistical Society, 101(4):770–771, ???? 1938. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980515.

S:1938:BRBk

[982] R. W. S. Book review: Population Pressure and Economic Life in Japan, by Ryoichi Ishii. Journal of the Royal Statistical Society, 101(4): 771–773, ???? 1938. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980516.

T:1938:BRBb

[983] C. T. Book review: Le Juif Delinquant: Etude Comparative sur la Criminalite de la Population Juive et Non-Juive de la Republique Polonaise, by Liebmann Hersch. Journal of the Royal Statistical Society, 101(4):773–774, ???? 1938. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980517.

Anonymous:1938:ONPd

[984] Anonymous. Other new publications. Journal of the Royal Statistical Society, 101(4):774–776, ???? 1938. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980518.

Anonymous:1938:SNd

[985] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 101 (4):777–793, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980519.

Anonymous:1938:CNd

[986] Anonymous. Current notes. Journal of the Royal Statistical Society, 101 (4):794–795, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980520.

Anonymous:1938:SEAd

[987] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 101 (4):796–799, ???? 1938. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980521.

Anonymous:1938:LALd

[988] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 101(4):800–804, ???? 1938. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980522.

Anonymous:1938:PRL

[989] Anonymous. Periodicals received by the Library. Journal of the Royal Statistical Society, 101(4):805–819, ???? 1938. ISSN 0952-8385. URL http:/

                                        /www.jstor.org/stable/2980523.

Anonymous:1939:VIa

[990] Anonymous. Volume information. Journal of the Royal Statistical Society, 102(1):v–vii, ???? 1939. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980134.

Anonymous:1939:FM

[991] Anonymous. Front matter. Journal of the Royal Statistical Society, 102(1):i–iii, ???? 1939. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980135.

Bowley:1939:PE

[992] A. L. Bowley. Production and efficiency. Journal of the Royal Statistical Society, 102(1):1–16, ???? 1939. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980136.

Anonymous:1939:PM

[993] Anonymous. Proceedings of the meetings. Journal of the Royal Statistical Society, 102(1):16–20, ???? 1939. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980137.

Kendall:1939:GDC

[994] M. G. Kendall. The geographical distribution of crop productivity in England. Journal of the Royal Statistical Society, 102(1):21–62, ???? 1939. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980138.

Carruthers:1939:TNP

[995] A. Stanley Carruthers. The trend of net profits of commercial and industrial enterprises, 1928–37. Journal of the Royal Statistical Society, 102(1):63–80, ???? 1939. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980139.

Thompson:1939:NIA

[996] R. J. Thompson. The new index of agricultural prices. Journal of the Royal Statistical Society, 102(1):81–82, ???? 1939. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980140.

Bowley:1939:IIS

[997] A. L. Bowley. The International Institute of Statistics. Journal of the Royal Statistical Society, 102(1):83–85, ???? 1939. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980141.

K:1939:BRBa

[998] M. G. K. Book review: Wahrscheinlichkeit, Statistik and Wahrheit, by R. Von Mises. Journal of the Royal Statistical Society, 102(1):86–89, ???? 1939. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980142.

S:1939:BRBa

[999] B. B. S. Book review: The Assessment of Psychological Qualities by Verbal Methods, by P. E. Vernon. Journal of the Royal Statistical Society, 102 (1):89–92, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980143.

M:1939:BRBa

[1000] H. W. M. Book review: A Century of Bank Rate, by R. G. Hawtrey. Journal of the Royal Statistical Society, 102 (1):92–94, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980144.

H:1939:BRBa

[1001] R. G. H. Book review: Modern Money, by Myra Curtis; Hugh Townshend. Journal of the Royal Statistical Society, 102(1):94–96, ???? 1939. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980145.

E:1939:BRBa

[1002] B. E. Book review: Modern Banking, by R. S. Sayers. Journal of the Royal Statistical Society, 102(1):96–97, ???? 1939. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980146.

G:1939:BRBa

[1003] C. O. G. Book review: The Finance of British Government 1920–1936, by Ursula K. Hicks. Journal of the Royal Statistical Society, 102(1):97–98, ???? 1939. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980147.

M:1939:BRBb

[1004] H. W. M. Book review: Slump and Recovery, 1929–27, by H. V. Hodson. Journal of the Royal Statistical Society, 102(1):99–100, ???? 1939. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980148.

S:1939:BRBb

[1005] R. W. S. Book review: Expectations, Investment and Income, by G. L. S. Shackle. Journal of the Royal Statistical Society, 102(1):100–102, ???? 1939. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980149.

M:1939:BRBc

[1006] H. W. M. Book review: The Lessons of Monetary Experience, Essays in Honor of Irving Fisher. Journal of the Royal Statistical Society, 102(1): 102–103, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980150.

W:1939:BRBa

[1007] G. R. W. Book review: Location Theory and the Shoe and Leather Industries, by Edgar M. Hoover, Jn. Journal of the Royal Statistical Society, 102 (1):103–104, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980151.

L:1939:BRB

[1008] H. L. Book review: Trade Associations in Law and Business, by Benjamin S. Kirsh. Journal of the Royal Statistical Society, 102(1):104–105, ???? 1939. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980152.

G:1939:BRBb

[1009] C. O. G. Book review: De la Nature de la Richesse et de l’Origine de la Valeur, by Auguste Walras. Journal of the Royal Statistical Society, 102 (1):105–106, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980153.

G:1939:BRBc

[1010] R. F. G. Book review: The Worker’s Standard of Living, by International Labour Office. Journal of the Royal Statistical Society, 102(1): 106–107, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980154.

G:1939:BRBd

[1011] R. F. G. Book review: Poverty and Population, by Richard M. Titmus; Lord Horder. Journal of the Royal Statistical Society, 102(1):107–108, ???? 1939. ISSN 0952-8385. URL http:/

                                        /www.jstor.org/stable/2980155.

T:1939:BRBa

[1012] R. J. T. Book review: Wheat Commission. Report upon the Administration of the Wheat Act, 1932, from July 1, 1932 to July 31, 1937. Journal of the Royal Statistical Society, 102(1):108, ???? 1939. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980156.

T:1939:BRBb

[1013] R. J. T. Book review: An Economic Survey of Buckinghamshire. Part I. Farms and Estates. Journal of the Royal Statistical Society, 102(1):108–

109, ????  1939.  ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980157.

Anonymous:1939:ONPa

[1014] Anonymous. Other new publications. Journal of the Royal Statistical Society, 102(1):109–110, ???? 1939. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980158.

Anonymous:1939:SNa

[1015] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 102 (1):111–132, ???? 1939. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980159.

Anonymous:1939:CNa

[1016] Anonymous. Current notes. Journal of the Royal Statistical Society, 102 (1):133–136, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980160.

M:1939:AH

[1017] H. W. M. Alfred Hoare. Journal of the Royal Statistical Society, 102 (1):137–138, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980161.

A:1939:WWK

[1018] P. A. A. W. Waterlow King. Journal of the Royal Statistical Society, 102(1):138, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980162.

Anonymous:1939:SEAa

[1019] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 102 (1):139–146, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980163.

Anonymous:1939:LALa

[1020] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 102(1):147–158, ???? 1939. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980164.

Anonymous:1939:PRa

[1021] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 102 (1):159–166, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980165.

Crump:1939:ETR

[1022] Norman Crump. The economics of the third Reich. Journal of the Royal Statistical Society, 102(2):167–212, ???? 1939. ISSN 0952-8385. URL http:/

                                        /www.jstor.org/stable/2980003.

Leak:1939:CTB

[1023] H. Leak. The carrying trade of British shipping. Journal of the Royal Statistical Society, 102(2):213–266, ???? 1939. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980004.

Anonymous:1939:WP

[1024] Anonymous. Wholesale prices in 1938. Journal of the Royal Statistical Society, 102(2):267–284, ???? 1939. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980005.

Yates:1939:AWC

[1025] F. Yates. The adjustment of the weights of compound index numbers based on inaccurate data. Journal of the Royal Statistical Society, 102 (2):285–288, ???? 1939. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980006.

Ramsbottom:1939:WRU

[1026] E. C. Ramsbottom. Wage rates in the United Kingdom in 1938. Journal of the Royal Statistical Society, 102 (2):289–291, ???? 1939. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980007.

Anonymous:1939:SDC

[1027] Anonymous. The statistical dinner club 1839–1939. Journal of the Royal Statistical Society, 102(2):292–294, ????

1939. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980008.

N:1939:BRB

[1028] J. N. Book review: A Study in Analysis of Stationary Time Series, by Herman Wold. Journal of the Royal Statistical Society, 102(2):295–298, ???? 1939. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980009.

Y:1939:BRBa

[1029] F. Y. Book review: Statistical Methods for Research Workers, by R. A. Fisher. Journal of the Royal Statistical Society, 102(2):298, ???? 1939. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980010.

K:1939:BRBb

[1030] M. G. K. Book review: Statistical Tables for Biological, Agricultural and Medical Research, by R. A. Fisher; F. Yates. Journal of the Royal Statistical Society, 102(2):298, ???? 1939. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980011.

K:1939:BRBc

[1031] M. G. K. Book review: Tables for Calculating by Machine Logarithms to 13 Places of Decimals, by Frederic Deprez. Journal of the Royal Statistical Society, 102(2):299, ???? 1939. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980012.

K:1939:BRBd

[1032] M. G. K. Book review: Duodecimal Arithmetic, by George S. Terry. Journal of the Royal Statistical Society, 102 (2):299–300, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980013.

Y:1939:BRBb

[1033] F. Y. Book review: The Synthetic Optimum of Population, by Imre Ferenczi. Journal of the Royal Statistical Society, 102(2):300–301, ???? 1939. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980014.

G:1939:BRBe

[1034] M. G. Book review: Medical and Sanitary Reports from British Colonies, Protectorates and Dependencies for the Year 1936, by P. Granville Edge. Journal of the Royal Statistical Society, 102 (2):301–302, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980015.

R:1939:BRBa

[1035] E. C. R. Book review: Introduction to Economic Statistics, by W. L. Crum; A. C. Patton; A. R. Tebbutt. Journal of the Royal Statistical Society, 102 (2):302–304, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980016.

H:1939:BRBb

[1036] R. G. H. Book review: Oxford Economic Papers. Journal of the Royal Statistical Society, 102(2):304–

307, ????  1939.  ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980017.

H:1939:BRBc

[1037] R. G. H. Book review: Value and Capital: An Inquiry into Some Fundamental Principles of Economic Theory, by J. R. Hicks. Journal of the Royal Statistical Society, 102(2):307–312, ????

1939. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980018.

S:1939:BRBc

[1038] R. W. S. Book review: Reinvestment Cycles and their Manifestation in the Norwegian Shipping Industry, by Johan Einarsen. Journal of the Royal Statistical Society, 102(2):312–313, ???? 1939. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980019.

M:1939:BRBd

[1039] H. W. M. Book review: International Combines in Modern Industry, by Alfred Plummer. Journal of the Royal Statistical Society, 102(2): 313–314, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980020.

M:1939:BRBe

[1040] H. W. M. Book review: Argentine Meat and the British Market, by Simon G. Hanson. Journal of the Royal Statistical Society, 102(2):315–316, ???? 1939. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980021.

M:1939:BRBf

[1041] H. W. M. Book review: The Growth of Collective Economy, by F. E. Lawley. Journal of the Royal Statistical Society, 102(2):316–317, ???? 1939. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980022.

G:1939:BRBf

[1042] C. O. G. Book review: A History of Economic Thought, by Erich Roll. Journal of the Royal Statistical Society, 102(2):317–318, ???? 1939. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980023.

Anonymous:1939:ONPb

[1043] Anonymous. Other new publications. Journal of the Royal Statistical Society, 102(2):318–320, ???? 1939. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980024.

Anonymous:1939:SNb

[1044] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 102 (2):321–330, ???? 1939. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980025.

Anonymous:1939:CNb

[1045] Anonymous. Current notes. Journal of the Royal Statistical Society, 102 (2):331–332, ???? 1939. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980026.

B:1939:GHL

[1046] A. L. B. G. I. H. Lloyd. Journal of the Royal Statistical Society, 102(2):333, ???? 1939. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980027.

Anonymous:1939:SEAb

[1047] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 102 (2):334–339, ???? 1939. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980028.

Anonymous:1939:LALb

[1048] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 102(2):340–345, ???? 1939.

ISSN 0952-8385. URL http://www.

jstor.org/stable/2980029.

Anonymous:1939:PRb

[1049] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 102 (2):346–353, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980030.

Greenwood:1939:SDU

[1050] Major Greenwood. The social distribution of university education. Journal of the Royal Statistical Society, 102 (3):355–383, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980064.

Rhodes:1939:JD

[1051] E. C. Rhodes. Juvenile delinquency.

Journal of the Royal Statistical Society, 102(3):384–405, ???? 1939. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980065.

Hartley:1939:RAM

[1052] H. O. Hartley. Recent advances in mathematical statistics. Journal of the Royal Statistical Society, 102(3): 406–444, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980066.

Muhsam:1939:NMV

[1053] Helmut Muhsam. A note on migration and Verhulst’s logistic curve. Journal of the Royal Statistical Society, 102 (3):445–448, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980067.

E:1939:BRBb

[1054] W. P. E. Book review: Karl Pearson, An Appreciation of Some Aspects of his Life and Work, by E. S. Pearson. Journal of the Royal Statistical Society, 102 (3):449–451, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980068.

K:1939:BRBe

[1055] M. G. K. Book review: Principles of the Mathematical Theory of Correlation, by A. A. Tschuprov; M. Kantorowitsch. Journal of the Royal Statistical Society, 102(3):451–452, ???? 1939. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980069.

K:1939:BRBf

[1056] M. G. K. Book review: An Introduction to Modern Statistical Methods, by Paul R. Rider. Journal of the Royal Statistical Society, 102(3):452, ???? 1939. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980070.

R:1939:BRBb

[1057] E. C. R. Book review: The Natural History of Population, by Raymond Pearl. Journal of the Royal Statistical Society, 102(3):452–454, ???? 1939. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980071.

G:1939:BRBg

[1058] D. V. G. Book review: Apercu de la Demographie des Divers Pays du Monde 1929–1936. Journal of the Royal Statistical Society, 102(3):454–455, ???? 1939. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980072.

G:1939:BRBh

[1059] D. V. G. Book review: Das Bevolkerungsproblem in Frankreich, by H. von Posadowsky-Wehner. Journal of the Royal Statistical Society, 102(3):455, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980073.

B:1939:BRB

[1060] H. S. B. Book review: Studies in Income and Wealth. Journal of the Royal Statistical Society, 102(3): 456–458, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980074.

F:1939:BRB

[1061] A. W. F. Book review: The Significance and Basic Postulates of Economic Theory, by T. W. Hutchison. Journal of the Royal Statistical Society, 102(3):458–459, ???? 1939. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980075.

G:1939:BRBi

[1062] C. O. G. Book review: Economics: A General Textbook for Students, by Frederic Benham. Journal of the Royal Statistical Society, 102(3):459–460, ???? 1939. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980076.

G:1939:BRBj

[1063] R. F. G. Book review: Accounting for Economic Management, by D. S. Blacklock. Journal of the Royal Statistical Society, 102(3):460–461, ???? 1939. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980077.

M:1939:BRBg

[1064] H. W. M. Book review: Consumers’ Co-operation in Great Britain: An Examination of the British Co-operative Movement, by A. M. Carr-Saunders; P. Sargant Florence; Robert Peers; Cooperation at Home and Abroad: A Description and Analysis. Volume II, 1908–1938, by C. R. Fay. Journal of the Royal Statistical Society, 102 (3):461–462, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980078.

G:1939:BRBk

[1065] R. F. G. Book review: The People’s Food, by William Crawford; H. Broadley. Journal of the Royal Statistical Society, 102(3):463–464, ???? 1939. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980079.

G:1939:BRBl

[1066] R. F. G. Book review: The Home Market, by G. Harrison; F. C. Mitchell. Journal of the Royal Statistical Society, 102(3):464, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980080.

T:1939:BRBc

[1067] R. J. T. Book review: Economics of Peasant Farming, by Doreen Warriner. Journal of the Royal Statistical Society, 102(3):464–465, ???? 1939. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980081.

Anonymous:1939:ONPc

[1068] Anonymous. Other new publications. Journal of the Royal Statistical Society, 102(3):466–467, ???? 1939. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980082.

Anonymous:1939:SNc

[1069] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 102 (3):468–481, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980083.

Anonymous:1939:CNc

[1070] Anonymous. Current notes. Journal of the Royal Statistical Society, 102 (3):482–484, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980084.

Anonymous:1939:SEAc

[1071] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 102 (3):485–491, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980085.

Anonymous:1939:LALc

[1072] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 102(3):492–499, ???? 1939. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980086.

Anonymous:1939:VIb

[1073] Anonymous. Volume information.

Journal of the Royal Statistical Society, 102(4):652–657, ???? 1939. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980166.

Gordon:1939:CMD

[1074] A. P. L. Gordon. The capital market of to-day. Journal of the Royal Statistical Society, 102(4):501–531, ???? 1939. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980167.

Wishart:1939:SAT

[1075] John Wishart. Some aspects of the teaching of statistics. Journal of the Royal Statistical Society, 102(4): 532–564, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980168.

Lewis-Faning:1939:SRD

[1076] E. Lewis-Faning. Statistics relating to the deterrent element in flogging. Journal of the Royal Statistical Society, 102 (4):565–578, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980169.

Anonymous:1939:RCF

[1077] Anonymous. Report of the Council for the financial year ended December 31st, 1938, and for the sessional year ending June 20th, 1939, presented at the one hundred and fifth Annual General Meeting of the Royal Statistical Society, held in the Hall of the Royal Society of Arts, John Street, Adelphi, W.C. 2 on June 20th, 1939. Journal of the Royal Statistical Society, 102 (4):579–591, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980170.

Anonymous:1939:POH

[1078] Anonymous. Proceedings of the one hundred and fifth Annual General Meeting of the Royal Statistical Society held in the Hall of the Royal Society of Arts on Tuesday, June 20th, 1939. Journal of the Royal Statistical Society, 102(4):592–594, ???? 1939. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980171.

M:1939:BRBh

[1079] H. W. M. Book review: The Gold Standard in Theory and Practice, by R. G. Hawtrey. Journal of the Royal Statistical Society, 102(4):595–596, ???? 1939.

ISSN 0952-8385. URL http://www.

jstor.org/stable/2980172.

W:1939:BRBb

[1080] J. E. W. Book review: Essays in the Theory of Economic Fluctuations, by M. Kalecki. Journal of the Royal Statistical Society, 102(4):596–599, ???? 1939. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980173.

H:1939:BRBd

[1081] R. G. H. Book review: Central Banking, by M. H. de Kock. Journal of the Royal Statistical Society, 102(4): 599–600, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980174.

G:1939:BRBm

[1082] C. O. G. Book review: The History and Mechanism of the Exchange Equalisation Account, by Leonard Waight. Journal of the Royal Statistical Society, 102(4):600–601, ???? 1939. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980175.

G:1939:BRBn

[1083] C. O. G. Book review: Public and Private Property in Great Britain, by H. Campion. Journal of the Royal Statistical Society, 102(4):601–602, ???? 1939. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980176.

S:1939:BRBd

[1084] R. W. S. Book review: A Study in English Local Authority Finance, by Joseph Sykes. Journal of the Royal Statistical Society, 102(4):602–603, ???? 1939. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2980177.

G:1939:BRBo

[1085] R. F. G. Book review: Incomes, Means Tests and Personal Responsibility, by P. Ford. Journal of the Royal Statistical Society, 102(4):603–604, ???? 1939. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980178.

A:1939:BRB

[1086] J. E. A. Book review: Railways, Roads and the Public, by Charles Stuart-Williams; Ernest Short. Journal of the Royal Statistical Society, 102(4): 604–605, ???? 1939. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980179.

M:1939:BRBi

[1087] H. W. M. Book review: State Interference in South Africa, by F. J. van Biljoen. Journal of the Royal Statistical Society, 102(4):605–606, ???? 1939. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980180.

G:1939:BRBp

[1088] M. G. Book review: Walter Bagehot, by William Irvine. Journal of the Royal Statistical Society, 102(4): 607–608, ???? 1939. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980181.

M:1939:BRBj

[1089] H. W. M. Book review: An Eighteenth Century Industrialist: Peter Stubs of Warrington, 1756–1806, by T. S. Ashton. Journal of the Royal Statistical Society, 102(4):608–609, ???? 1939. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2980182.

Anonymous:1939:ONPd

[1090] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 102(4):609–610, ???? 1939. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980183.

Anonymous:1939:SNd

[1091] Anonymous. Statistical notes. Journal of the Royal Statistical Society, 102 (4):611–620, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980184.

Anonymous:1939:CNd

[1092] Anonymous. Current notes. Journal of the Royal Statistical Society, 102 (4):621–623, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980185.

F:1939:CLC

[1093] A. W. F. Clement-Leon Colson b. 1853; d. 1939. Journal of the Royal Statistical Society, 102(4):624, ???? 1939. ISSN

0952-8385. URL http://www.jstor.
org/stable/2980186.

Anonymous:1939:CHW

[1094] Anonymous. Charles Henry Wickens b.

1872; d 1939. Journal of the Royal Statistical Society, 102(4):625, ???? 1939. ISSN 0952-8385. URL http://www.

jstor.org/stable/2980187.

Anonymous:1939:SEAd

[1095] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 102 (4):626–630, ???? 1939. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2980188.

Anonymous:1939:LALd

[1096] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 102(4):631–636, ???? 1939. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2980189.

Anonymous:1939:PRL

[1097] Anonymous. Periodicals received by the Library. Journal of the Royal Statistical Society, 102(4):637–651, ???? 1939. ISSN 0952-8385. URL http:/

                                        /www.jstor.org/stable/2980190.

Anonymous:1939:BM

[1098] Anonymous. Back matter. Journal of the Royal Statistical Society, 102(4):??, ???? 1939. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2980191.